A Complete Bibliography of Publications in IEEE Spectrum (1990–1999)

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

26 December 2021 Version 1.01

Title word cross-reference

$12 [Swe98c]. $19.95 [Swe98c]. $25 [Rob99]. 3 [Ell99b]. $30 [Ano97b, Wis99]. $32.50 [Mor98a]. $35.95 [Fie97]. _{DDQ} [HS96a].

-D [Ell99b].

// [GS97].

0-8027-1346-7 [Rob99].

100-million [Ano99a]. 12 [McA98]. 1940’s [Ede99]. 1958 [Lik92]. 1992-data [Low92]. 1996 [Ano96a]. 1997 [Ano97-54, Ano97a, Hei97]. 1999 [Ano99b, Ste99].

2 [Bel90g, Noa97]. 200-Year [Rob99]. 2000 [CP98, LP98]. 21st [Jur90c]. 25 [Doo94b].

3 [PS92, Wil93]. 3-V [PS92, Wil93].

4 [Koe99].

5 [Wil93]. 5-V [Wil93].

64 [Wat91a].

’90 [AE90, Bel90g, Fit90g, MV90, Mur90d, SC90, Sti90b, Zor90g, Zor90h, Bel90b, Che90b, Che90a, Fit90a, Jur90a, Kap90a, Mur90a, Voe90b, Voe90a]. ’90s [Kor94, Bel90a, Cam90]. ’92 [Kap90b].

= [Ada92a, Cen90]. abstract [Fos93]. abstraction [LHD^{+}90]. AC [Hin93, Rad96]. academe [Gup94]. Accelerator [Doo96b]. access [DR99c]. accessible [PA91]. accident [Obe99d]. according [Luc98f]. acquisition [Cer92, Fow94, Hou95, Rie93, Tra93]. across [Kur99]. Act [Bra96a, RTHM92, Tro99]. action [Kor93]. actions [Ben93]. Active [Kor94, Ell99a]. acts [Bel90e]. actuators [HMGT90]. Adaptable [Ell99b, Mas99a]. Adapting [Tra93]. Adaptive [Cox93]. Add [Com92a, Ano91a]. Add-ons [Com92a, Ano91a]. addiction [Luc99c]. adding [Eve98]. Addressing [BS99]. adherents [Fit90b]. Adopting [Dav94]. adorns [Ell99b]. ADSL [DR99e]. advance [DT92]. Advanced [Bel96a, Mal93, Swe97a, Kor94, SB99b]. advancement [YL90]. advances [Che90b, Swe94b]. Advancing [Kob94]. adversaries [Jur90b]. advertisements [Hos95a]. Aerodynamics [Ano97b]. Aerospace [Bra99, Doo93a, Doo94a, Doo95a, Doo96a, Doo97a, Doo98, Esc90, Kas94, AE90, AE91]. aerospace/military [AE91]. affair [Run99]. Affordable [Hin90, Mir96a]. African [Mak94]. after [Doo94b, Rie91a, Sha93, Swe96b, Sym98]. after-effects [Swe96b]. against [Mar90]. Age [Mar96, Ada91b, PF96, AAH98]. agency [Zor91b]. Agents [Mor98b, Sin96]. aging [TPB95]. ago [Jon97, Jon98f, Jon98g, Jon98h, Jon98i, Jon98j, Jon98k, Jon98b, Jon98c, Jon98d, Jon98e, Jon98l, Jon99b, Jon99c, Jon99d, Jon99e, Jon99f, Jon99g, Jon99h, Jon99i, Jon99j, Jon99k, Jon99l, Jon99m, Swe97f]. agreement [Ove97]. Ahead [Fuq96, Par96, RW98, Ano98-103, Ros91c]. aided [CK96, Low97, RGH90, VG96]. Aids [Han98a]. aim [WK90]. Air [Ada91a, Per91, PA91, LRC99, Per97c, WH99, Bra96a]. airborne [For97]. airport [Mec95]. Alfred [Bel94a]. algorithm [Kar96a]. algorithms [Eve98]. Ali [Ano96-42]. all-metal [Joh94]. allocation [HT95]. Alpha [Com92b]. Alps [Sto95]. always [Fit90h]. am [Luc96e]. America [Rud96]. American [Fos97f, Mor98a]. Americans [Sti90a]. among [Wis96]. amplifiers [Bur93]. analog [Cha97b, Gep99c, Hin90, Rho96]. analyses [HW94]. Analysis [Bel94b, GS99, BR97, Bet99, Bra99, Bre99, BDG96, Cha97a, Com97b, Com98a, Com99b, Doe97, Doo98, DR99f, Fer98, Fos98c, Gep98g, Gep99g, Hel98, Hin98, Jam98, JJ91, Kap97b, Kap98a, Kap99b, Kor94, Kor99, Lan98, Lan99, LK99b, Mar99a, Mas98, Mas99c, Mir98, Per98b, Phi98, RS99a, Rie98e, Ros98, SS99, Spa97, Sta93, Swe98e, Swe99f, Trl97]. analyzed [Win96]. anarchy [Luc96d]. Anatomy [SCS91]. Andrew [Hil99]. annual [Com96d]. anticipates [LHD^{+}90]. Any [Luc97a, Lan94, Obe99c]. anywhere [Luc97a]. Apollo [Doo94b]. apostles [Mem97]. Apple [Hin99]. appliances [BC98]. Application [Bra95a, Gab94a]. applications [Ada93a, Aka97, Bar98, Eub98, JPN^{+}92, Trl96, Bel90b, Che90b, Che90a, Fit90a, Jur90a, Kap90a, Mur90a, Voe90b, Voe90a]. Applied [Kar96b]. applying [Ano90d]. Approach [Mos96, Fre98, Noa97, Wat92a]. Approaching [HI92]. Appropriate [Kap94]. architecting [Rec92, Rie92c]. Architects [Ada96a]. architecture [Com95a, PA97]. architectures [BH95, Gep98b]. area [GB91, MC90, Sto97, Wic96b]. area-wise [Sto97]. arsenals [Zak99]. arsenide [Cat90, Dey95]. Art [Mil98b, Cha99, Ell99b, Mil98d, Mur90c, Rec92]. artificial [Ada91a, Cla98, DM96]. artists [Mil98c]. Asia [Bal91]. Asian [LRC99]. ASIC [Lev92]. ASICs [DR91, Smi92a]. aspects [Fra96, Mag97]. assembly [BK93]. assessment [KW92]. astronomy [But97, Zor95b]. athletes [Per90a]. ATM [Lan94, Mil94a, Rui94]. audio [Jur96]. Authors [Ano96-72]. Auto [Kap97a]. Automate [GOI98]. automated [Kuk93]. Automating [Top91]. automation [Cai97, Mar99a, Mur90b, Swe95]. Automotive [KWMH96]. autonomous [Pas95]. Autopilots [Zor94a]. Award [Ano96o, Ano96-27, Ano96z, Ano96-40, Ano96-42, Ano96-41, Ano96-52, Ano96-69, Ano96-70, Ano96-71, Ano96-73, Ano96-74, Ano96-75, Ano96-76, Ano96-77, Ano96-78, Ano96-91, Ano96-103, Ano96-104, Ano96-105, Ano96-124, Ano96b]. Awards [Ano96-122, Ano96-123, Ano97-54, Ano96a, Ano98-63, Ano97a, Ano98-62, Ano99-54]. ax [Fow94].

B [Ano96-77, Ano96-103, DP91]. B-ISDN [DP91]. B.C [Ede99]. Back [Ano97c, Asp90]. Backing [Nor99, Spi94]. backlot [Bra95b]. Bad [Jon98a, Bel90a, BP92]. baggage [Edw97]. Baker [Ano96-124]. Balancing [Bra96a, Log99, Cat92]. Ballistic [Ano97c, LP97a, Mos97, SLLP97]. ban [Swe96a]. bandwidth [Wil97]. Banking [McC97]. barrier [Rie92d]. Bart [Ada96h]. based [Apf97, GF98, UT99]. Basic [Hel98, Hel96, Hel97, Mor99a, Saw95]. Basics [Asp90]. basis [Bel93a]. bat [SSD92]. batteries [MH98, Rie95a]. battery [Rie92d, lH98]. battle [Com97a, Com98b]. battlefield [Zor91c]. Battling [Swe99a]. Baum [Ano96o]. bay [HJ93]. Be [Gad98, Ada90b, Kel99, von94]. Beam [Doo96b, Har99b]. Beckman [Bel96a]. becomes [Run99]. becoming [Dil94]. beer [Gep99b]. beget [Eub98]. Beginnings [Fos97f]. behaviour [Tay99]. Behind [Mor97a, Nah91]. Below [Tiw94]. Berlin [BP98]. best [Bel91d, Kra96]. bets [Gut91]. Better [Swe96a, Mar98a, Rie95a, SSH99, Swe99c, Zor90c]. Bettyann [Fie97].

Between [Sto95]. Beyond [BC93, Bra96b, Bel92a, Doo95b]. Bicycles [Bel93a]. big [ARA91, Che90c, GOI98, Run99, Ste99]. Bigger [MM98, Swe99c]. Billy [Mil98c]. biography [Ada96h, Bel94a, Kap99a, Per95d, Per95e, Per98a, Zac95]. Biological [NM90]. Biomechanically [Per90a]. biometrics [Mil94b]. Biotechnology [But96]. bipolar [Cre95]. birth [Swe97b]. bit [Bel90e, Luc94a]. Blind [Ano97-82, NMGW96, WR96]. Blinking [Luc97b]. Block [Han96, Han97, Han98b]. Blocking [HR96]. blowing [Fit90h]. Blue [GON93, Ano97-38, Flo90, Mar98d]. Blueprinting [Kra98]. Board [Ano90a]. boardroom [Per94b]. boards [Ber99, BC99, Cha97a]. Boeing [Nor95]. bolts [Mar98d]. bomb [Kra98, LY96, MM98, Zor92d]. bonds [Zor91a]. Bone [Fie97]. Book [Bai96, BDF99, Bro99, Cla98, Day96, Ede97, Ede99, Ell96, Eng97a, Fay99, Fie97, Fos96a, Fos99a, Gad98, Kam99, Mac98, Mar96, Mar97a, Mor99a, Mor99b, Pfl96, Rob97, Rob99, Rod99, Ste98a, Swe98c, Wis99, Zor97, Bel96b, Bry96, But96, Wis96]. Books [Ano98a, Bel98, Bra96b, Kra98, Lat97, Lin96, Mem98, Mes98, MM98, Rob97, Ste96, Swa96, Swe96e, Swe98d, Swe98g, Swe99e, Swe99d, Zim97, Alt96, Ano96-95, Ano96-96, Ano96-97, Ano96-98, Ano96-99, Ano96-100, Ano96-101, Ano96-102, Ano97-71, Ano97-72, Ano97-73, Ano97-74, Ano97-75, Ano97-76, Ano97-77, Ano97-78, Ano97-79, Ano97-80, Ano97-81, Ano98-77, Ano98-78, Ano98-79, Ano98-80, Ano98-81, Ano98-82, Ano98-83, Ano98-84, Ano98-85, Ano98-86, Ano98-87, Ano98-88, Ano99-68, Ano99-69, Ano99-70, Ano99-71, Ano99-72, Ano99-73, Ano99-74, Ano99-75, Ano99-76, Ano99-77, Ano99-78, Ano99-79, Kar96b, Log99, Mun96]. boosts [FASN96]. born [Lik92]. both [Cla98, Fos97f]. bottleneck [GOI98]. bound [Fin90]. boundaries [Ada90d]. box [Cic95, Flo90]. Bozos [Luc96a]. Bracing [KA90]. brains [KS92]. Brazil [DR96, Fil96, Ros96]. Breakthroughs [GS98]. Breezy [Pac99]. Bring [DR99a]. Bringing [LK99a, Mir99a]. brings [SZ96]. brink [Com92c]. Broad [Bis95, Buc95, Dua95, PM98]. Broadcasting [Jur96, Rei92a]. broadly [Fox99]. Brooding [CP98]. Browder [Ano96o]. Brunetti [Ano96-27]. Brunswick [Fie97]. budget [Gep99b]. bug [CP98]. Building [Str96]. Building-integrated [Str96]. bus [KHSK95, Luc96a, Wic97]. buses [Box95]. Bush [Mor98a, Zac95]. busier [SW97]. Business [MF97, Jur90e, Tan98]. bust [Bel91d]. butting [Sto98]. bye [Jur92b].

C [Mir96c, Mir97]. Cable [DR99b, RN97]. cabling [Tho96]. cache [McM96, Pix96]. CAD [Fox93]. CAE [Ano90e, Cen90]. calculus [Swa96]. Calendar [Ano96p, Ano96q, Ano96r, Ano96s, Ano96t, Ano96u, Ano96v, Ano96w, Ano96x, Ano96y, Ano97p, Ano97q, Ano97r, Ano97s, Ano97t, Ano97u, Ano97v, Ano97w, Ano97x, Ano97y, Ano97z, Ano97-27, Ano98n, Ano98o, Ano98p, Ano98q, Ano98r, Ano98s, Ano98t, Ano98u, Ano98v, Ano98w, Ano98x, Ano99o, Ano99p, Ano99q, Ano99r, Ano99s, Ano99t, Ano99u, Ano99v, Ano99w, Ano99x, Ano99y, Ano99z]. Calibrating [Sha99]. California [BH96]. calls [Bel96c, Tur91]. calorie [Bla97]. Calstart [Lyn93]. campus [Fal96, Hil99]. Can [SH99, Ada90b, Ada92d]. Cane [Ano97-82]. cannot [Box95]. Canon [Chr90]. capacity [FASN96, RN97]. capitalism [Ste99]. Capitol [Rod99, Swe94c]. car [DU98, Jur90d, Tho96]. carbon [SB99c]. cards [Ste92]. Care [Fos97f]. career [Hos95a, YL90]. Careers [Bel96c, Gep99a]. Carl [Ano96o]. cars [CW94, Jur91b, von94]. case [Ano90b, Atw91, Gum99, Ano90b, Jon92]. cash [CB97, Gem97, ter97]. Cassini [Mar98c]. catalyst [Mil98c, Whi92]. catch [RLH93]. Catching [Ros96]. CCDing [McC92]. cell [FP93]. cells [Gil98, Ito90, Wol99]. cellular [Fis93, Gut91]. Center [Com95b]. Century [Rob97, Swe96e, Doo94b, Jur90c, BDF99, Fie97, Mor98a]. challenge [ZS90]. challenged [Zor90a]. Challenger [Bel96b]. challenges [Bel90f, IK99, Jur91a, LP97a, Per90d, SLLP97]. Champagne [Gep99b]. Chandy [Ano96-75]. change [Mor97a, Per92b, Zor93a]. Changed [Ste96, GS97]. Channeling [Roe97]. chaos [HJ93]. chapter [Kam99]. Characterizing [RDK90]. Charge [Toe98, Hos94b]. Charles [Ano96z, Kap99a]. Charms [Swe98g]. Chart [Mir95]. Charting [BH96]. charts [Doe97]. cheaper [Mar98a]. Check [Mos96]. Checking [Per99a, CMR96, McT96]. checks [CCS99]. Chemistry [Saw95]. Chernobyl [Swe96b]. Chiariglione [Swe97b]. Chicago [Swe98c]. China [Gep95a, HS99, SH99, WH99]. Chip [Tho97, BT93, CP96, Fuj96, HMGT90, HR96, You93, Zor99]. Chip-scale [Tho97]. Chipmaking [Gep95a]. chips [KM96, Mas90, Run99]. Cho [Bel94a]. choices [HK99]. choose [MP97]. Choosing [Hou95]. Christianson [Rob99]. churn [Luc98c]. CIM [Jai93, SJ93]. cinema [Bra95b]. circa [KWMH96]. circuit [Ber99, Bla96, Cam97, Lau97, Rie91b]. Circuits [Alt96, Ano96-105, Cha97b, Rho96]. class [Ben98, Shi91]. classic [Mor99b]. classroom [GF98]. Claude [Hor92]. Clean [HD96]. Cleaning [Per93a]. cleans [Gel96]. Cledo [Ano96-27]. Client [Lar94]. climate [Per93b, Zor93a]. climb [YL90]. clip [Ell99b]. clip-art [Ell99b]. Clock [Luc98a]. close [Doo94c]. clues [Jam98]. CMOS [Rie91b, SHM96, YS91]. Co [Rob99]. coal [BS99, SH99]. cockpit [Swe95]. code [BP92, Mir96c]. coexistence [Zor90c]. Coherence [Ell96, McM96]. Collaring [Ico97]. combats [Ada94b]. combined [BP98]. come [GC94]. comes [Fox98, PF96]. Coming [Ano96-29, Ano96-30, Ano96-31, Ano96-32, Ano96-33, Ano96-34, Ano96-28, Ano96-35, Ano96-36, Ano96-37, Ano96-38, Ano96-39, Ano97-28, Ano97-29, Ano97-30, Ano97-31, Ano97-32, Ano97-33, Ano97-34, Ano97-35, Ano97-36, Ano97-37, Ano98y, Ano98z, Ano98-27, Ano98-28, Ano98-29, Ano98-30, Ano98-31, Ano98-32, Ano98-33, Ano98-34, Ano98-35, Ano98-36, Ano99-27, Ano99-28]. comments [SLLP97]. commercial [SPM90]. commitment [Bar92]. commodity [Pet95]. Communication [Ano96b, Wic97]. Communications [Ano96-75, Ano98-89, BAL96, BR97, Cam90, Cru99, RS99a, Rie98e, Bor94, DP91, Fuj96, GB91, Hen97, Low92, Low95, Rei92b, Ros91a, Sar93, Sha91, Ste98b, Sti90b, Tay92, WBR91]. Community [Sti90a]. Compact [Fra95b, Lub95]. Companies [KWCR92, McA98, Jur90b, Shi91]. company [Hos95a]. compatible [NP90]. compelling [Gum99]. Compendium [Mem98]. Competing [Ada90a]. competition [Bel90c, Tan98]. Competitive [Kok93, Whi93a, KW92]. competitiveness [Che90d, RPD^{+}91, Whi92]. Complete [Bli96]. complex [Kal94, Rec97]. compliant [LP98]. Complying [Gre94]. compression [ARA91, Mem97]. computational [WJT96]. computationalism [Swe99d]. Computer [CK96, Low97, RGH90, VG96, Ano97-38, CCS99, Com93g, CMS97, Hin99, Mar90, Ng92, NP90, Whi93b]. Computer-aided [CK96, Low97, RGH90, VG96]. Computers [Ano96-75, Com96b, Com98a, DR99f, Jul97, KWC93, Mar97a, Bel91a, CZ90, Com97a, Com98b, Gep96b, Jul95, Khe94, Khe95, San96, Sei99, Zor92f, Zor93b]. computing [Fal96, Har92b, Hil99, SK92a]. Concept [Asp90]. concern [Chr90]. Concordia [Kap99a]. Concurrent [RWC91, SSZ93, Shi91, Tur91, WBR91, MC94]. Conducting [SS92]. conferencing [DR98]. conflicting [Mur90b]. Conformal [Mir98]. Congressional [Swe94c]. conjunction [But97]. Conn [Ano97b]. conservation [Per95b]. conserves [KHSK95]. Consolidating [PR92]. consortia [Zor90b]. Consorting [Mar98b]. consortium [Lyn93]. constitutional [DSD96]. consume [SH99]. Consumer [Ano96-76, Bra96c, Jur92c, Mil99a, Per90b, Per94a, Per95a, Per97a, Per98b, EB99, Jur90a, Jur91c, Jur93]. contenders [Com93c, Gep93]. Contents [Slo96, Ano96-106, Ano96-107, Ano96-108, Ano96-109, Ano96-110, Ano96-111, Ano96-112, Ano96-113, Ano96-114, Ano96-115, Ano96-116, Ano96-117, Ano97-84, Ano97-85, Ano97-86, Ano97-87, Ano97-88, Ano97-89, Ano97-90, Ano97-91, Ano97-92, Ano97-93, Ano97-94, Ano97-95, Ano98-90, Ano98-91, Ano98-92, Ano98-93, Ano98-94, Ano98-95, Ano98-96, Ano98-97, Ano98-98, Ano98-99, Ano98-100, Ano98-101, Ano99-82, Ano99-83, Ano99-84, Ano99-85, Ano99-86, Ano99-87, Ano99-88, Ano99-89, Ano99-90, Ano99-91, Ano99-92, Ano99-93]. Contest [Han96, Han97, Han98b]. continent [Fin90]. Continuous [Doo96b]. Contract [Fit90b]. Contrasting [Che90c]. control [Ada91a, Bel95a, Doo96d, Ell99a, Gel96, LD96, Pas95, PC98, Per91, PA91, Per97c, Raj94, SZ96, Smi93, Whi93b, Ano96-40]. controllers [PA97]. controls [Mas96, SW97]. conundrum [BS99]. Conversion [GMKJ92]. convert [RGG^{+}92]. cool [Bar98]. Cooperating [RKGB92]. Coping [Com95a]. coprocessors [Fer91]. Cornell [Com95b]. Corporate [SB98a, BC99]. Cortical [NMGW96]. Cost [Yea96, DMP96, Obe99c, SS91]. Count [Fos97f, Yea96]. countermeasures [Ada92b]. country [Doo96c]. courage [RGG^{+}92]. course [BH96, Obe99d]. courses [McA98]. court [EB99]. Courts [Mes98]. cover [Ano96c, Ano96d, Ano96e, Ano96k, Ano96l, Ano96m, Ano96f, Ano96n, Ano96g, Ano96h, Ano96i, Ano96j, Ano97d, Ano97e, Ano97f, Ano97g, Ano97h, Ano97i, Ano97j, Ano97k, Ano97l, Ano97m, Ano97n, Ano97o, Ano98b, Ano98c, Ano98d, Ano98e, Ano98f, Ano98g, Ano98h, Ano98i, Ano98j, Ano98k, Ano98l, Ano98m, Ano99c, Ano99d, Ano99e, Ano99f, Ano99g, Ano99h, Ano99i, Ano99j, Ano99k, Ano99l, Ano99m, Ano99n]. CPU [Gep93]. Craford [Per95e]. Cray [McT96]. Creating [Lar95]. creativity [Fit90d]. Credits [Sir97]. crescent [Bal91]. Crime [Mor97c]. criminals [HSW95]. critical [DL93]. Crossing [Gro98]. crossroads [Sai90]. Crucial [Lat97]. crunch [Luc94b]. cryptography [Ada92a]. Culture [Bel96b, Fit91a]. current [Bel90d]. currents [PC98]. curricula [Wat92b]. curve [Ast95]. custom [Hin95, Pet95]. customers [Swe99a]. cuts [Per90c]. Cutting [SS91]. cyber [Swe98f]. cybercrook [Ico97]. cyberspace [McC97]. cycle [BP98].

D [Ano97-38, Fos97f, Ell99b]. DACs [Hen97]. dad [Ano97-38]. dangers [Swe97c]. Daniel [Ano96-124]. dark [McC92]. DARPA [RWC91]. Data [Ada92a, Ada92b, Bor94, BP92, CDEJ99, Low95, RT98, Sar93, Sch91, Ano90c, Cer92, Che99, DP91, Doe97, Fos98d, Fox99, GB91, GM93, Hou95, Jam98, JS99, Kor93, Kor94, Lan94, Low92, LLKS95, Mas99b, Pen99, Rad96, Rie93, Roe97, Ros91a, SS99, Sha91, Sti90b, Tra93, WS93, Yal97]. databases [Atw91, Sta93]. datanets [GAL98]. dates [Gep96b]. Dave [BDF99]. David [Ano96-122, Ano96-41]. dawn [Fox95]. days [Jon98a]. DC [Hin96h]. Dealing [HS96a, GM93]. debits [Sir97]. Debugging [Fuj96, Pix96]. DEC [Com92b]. decade [Per91]. decentralized [Bel90f]. Decision [Bel96b, IO96]. deck [Swe95]. Deep [Ano97-38, Waf93, LY96, SHM96, VS97, Zor94a]. deep-submicron [LY96, SHM96, VS97]. defects [SHM96]. Defense [Ada90b, Ano97c, Fow94, Zor90a, Ada92d, Esc90, GMKJ92, KA94, LP97a, Mos97, Pan97, SLLP97]. defensive [Ber99]. define [PA91]. defying [Ben98]. Degrees [DFC94, Pan97]. delays [Gra99]. delivery [CK99]. deluge [GM93]. Demarche [Lat97]. density [Lub95]. dependability [CCS99]. deployable [Ada90b]. depth [Kap97a]. desert [Zor91d]. Design [BK93, Dom94, Gep99c, GW98, Han98a, Luc99a, PP93, Yof96, Ano98-103, Ber99, Bil96, Bla96, Cha99, Doo96d, Fre98, GOI98, Gep98c, Goy94, Hin90, HW94, Kar96a, Kur92, Low97, Mar99a, Mur90a, Mur90b, Rho96, Smi93, Sou98, SB99b, TC97, TT95]. design-automation [Mur90b]. Designing [HHSM93, Sel90, WL98, Har92a]. designs [Pac99, Sha99, Sto99]. desktop [Com97a, DR98]. destructive [Bel91d]. detection [FASN96]. developed [Com92b]. Development [Gep97f, Pfl96, Doo97a, Faw94, Hos95a, Per90c, Swe97a, Top91, YL94]. developments [Com96b, Gep96d]. Deviance [Bel96b]. devices [Cha97b]. Diagramming [Fin97]. diagrams [Lau97]. dial [Dix95]. dial-tone [Dix95]. diaries [Ano97-83]. diet [PS92, WS93, Yal97]. different [GLS94]. Digital [BBH97, Bli96, Boo99, DT92, Ede97, Fox98, Han98a, Jur92a, Mar96, Mem97, PF96, Bel99, Bra95b, Dal90, Fox95, Gep99c, Ham95, Jur91a, Jur96, Nin95, Per95c]. digitizes [Kap97a]. dimension [AC97, Bra96b]. diminishing [Luc98b]. directions [Bel93c]. dirty [Whi93b]. disabilities [Ada94b]. disabled [Laz99]. disappearing [Luc94b]. disc [Ast94]. Discussion [AHP^{+}95]. disk [Bel99, FASN96, Fri96]. disks [Lub95]. Dismantling [MW96]. dispatch [VG96]. Display [Luc93, Wer95]. Displaying [Ano90c]. displays [Wer97]. disposal [Bea97]. distance [Dom94]. distributed [Har92b]. Distribution [Ano96-71]. diversify [Zor90a]. Diversifying [CK99]. Diversity [Bar92, Bat92, DF92, Ric92, Smi92b, DFC94]. division [Jur90e]. Do [Gal96, PG96, Mar98d]. document [Jur91d]. Documentation [Mau94]. documents [Ell99b]. DOE [Col92, Zor91b]. does [Com99h]. Doing [SSH99, Ung96]. Donald [Ano96-42, Per98a]. DON’Ts [Fow93]. Doubling [SLHM94]. Down [Ell99a, BH92, Fox98]. downsize [Zor90a]. downsizing [Ada92d]. DRAM [IK99]. DRAMs [FPR^{+}92, Sal92]. dream [Ano97-38]. drive [Com94c, FASN96]. DSP [Gep98b, Gum99, IO96, Kar96a, LB96, Mat90, Mat91, Pac99, Rob92]. DSPs [EB99, Ste98b]. dumb [Luc97f]. dummies [Obe99a]. Durham [Ano96-104]. Dutton [Ano96-73]. dynamic [Bel99, JPN^{+}92, Mas90].

e-beam [Har99b]. e-cash [Gem97, ter97]. E-mail [PA92, Per92a]. e-safe [BC97a]. ear [SSD92]. Earl [War96]. Early [Mos96]. Earth [Fox98]. earthing [St.92]. East [Bal91, Per90e]. Eavesdropping [Kra96]. EC [Kap90b]. Ecology [All96]. economic [Che90c, Rod99]. economics [ter97]. Economy [Ben98, Ada90a, Fel97, UT99, Ste99]. Economy-class [Ben98]. edge [LD96, Wer95]. EDIF [ER90]. Edison [Wis99]. edited [BDF99]. Educating [Gep95b, Swe98a]. Education [Bis95, Buc95, Dua95, AW98, Bel93e, Cur93, EJ93, GF98, Gup94, Joh93, Mas95, McA98, Sur90, Tad98]. Edward [Ano96-52]. Edwin [Fef99]. Ee [Ano96-50, Rie96c, Rie97a, Rie97g, Ano97-39, Ano98-37, Ano98-38, Ano98-39, Ano98-40, Ano98-41, Ano98-42, Ano98-43, Ano98-44, Ano98-45, Ano98-46, Ano99-29, Mor98a]. EEs [Ano96-43, Ano96-44, Ano96-45, Ano96-46, Ano96-47, Ano96-48, Ano96-49, Ano96-51, Ano97-40, Ano97-41, Ano98-47, Ano99-38, Ano99-39, Ano99-40, Ano99-30, Ano99-31, Ano99-32, Ano99-33, Ano99-34, Ano99-35, Ano99-36, Ano99-37, Fow93, Per94b, Rie96a, Rie96b, Rie97b, Rie97c, Rie97d, Rie97e, Rie97f, Rie97h, Rie98a]. effectiveness [Pan97]. effects [Bra95b, LRC99, NM90, RTD97, Swe96b]. efficiency [Ada91a, Div96, Rie92b, Zor91g]. efficient [Rie93, SSH99]. efficiently [RMAA91]. effort [Cur93]. Electric [Ano96-71, Ram94, Rie92a, Rie92c, Rie92d, Rie92b, Gil98, GR97, HS98, KHSK95, Kur99, Lyn93, MT96, lH98]. Electrical [Luc98b, Che90b, Ede99, KWMH96, Per94b]. electricity [BH96, HS96b, Rud96]. Electromagnetic [Fit90c, MN90, NM90, MP97]. Electromagnetics [Fos95, Mir96a]. electromagnetism [Mal93]. Electromigration [LY96]. Electron [Doo96b, Har99b]. Electronic [Ada92c, Mar99a, Par97, Per92b, PA92, Per92a, Per97b, BLK98, CB97, HSW95, Jur95, Kel97, Kha96, Kra96, Kur92, NGOS98, Per98a, RTD97, Sto99]. Electronics [Ano96-76, Bel95a, Fra96, KA94, Mag97, Zor90b, Ada94a, Ada95a, Ada96i, Bra96c, Fit90g, Fit91c, Hin96h, Jur90a, Jur90d, Jur91c, Jur92c, Jur92d, Jur93, Kap90a, Kap91, Kap92b, Kap93, Kap95, Kap96, Kap97b, Kap98a, Kap99b, LK99b, Mil99a, Mor97d, Per90b, Per93a, Per94a, Per95a, PG96, Per97a, Per98b, RPD^{+}91, Ros98, Ste93]. Electrotechnology [Gel96]. elements [Jon92]. Elephants [Par96]. Elevators [Ish94]. Elizabeth [Zor92a]. Elusive [Len96]. Emanuel [Ano96-52]. Embedded [IK99, GOI98, LK99a, LP98]. Embedding [JJ91, Mat91, Rob92]. emergencies [GR97]. emerges [All97, Mat90]. Emerging [Bry96, UT99]. emissions [Zor94d]. Employed [Ste98a]. Employment [Bel96c, Bel90a]. emulating [Zor91c]. Encouraging [Fit90d, RWC91]. end [Jul94, RT98]. endanger [PG96]. Endless [Mor98a]. Enemy [Zim97, Mar98b]. Energy [Fit91b, Fit91a, HS99, Zor91b, Zor91a, Bla97, Log99, Per95b, SSH99, Swe96c, Swe97e, Swe98e, Swe99f, SB99c, Zor90f, Zor90h, Zor91f, Zor92g, Zor93c, Zor94c, Zor95a]. energy/resource [Per95b]. Engineer [Mor98a, Ada96h, Ano96-92, Ano96-93, Ano96-94, Ano97-67, Ano97-68, Ano97-69, Ano97-70, Ano98-76, Che90f, Fre98, Gep95b, GMKJ92, Gra99, Kap99a, Kel99, LP98, Mil98c, Sou98, SLHM94]. engineered [Per90a, Zor92d]. Engineering [Bel94c, Cer92, Com92a, Com94a, Dai92, Fos92, Fos93, Fos98a, Gal96, Har92a, HPV92, JB96, Kap92a, Kap92c, KS92, Kor93, Rie93, Rie98b, Rob92, Ros92a, SK92a, San90, Sky94, Smi92a, Swe98a, Ano91b, Ano91a, Che90b, Com95c, Com97b, Com99b, Cre95, Gab94b, Gep95d, GS97, Gup94, HK99, Hel99, Hin96g, Hin98, HFYT91, Hos94a, Kap99c, Luc98b, MC94, Mas95, Per98a, RMAA91, Rau93, RWC91, Rod99, SSZ93, Shi91, Sur90, Tur91, WBR91, YL90, Zor90c, BDF99]. Engineers [Com93a, Fos96a, Gep97g, KP98, Per92c, Fit90h, Gep96a, Gep99a, Per94b, Ric92, RGG^{+}92, Wol93]. enhances [SS96]. Enough [HD96]. ensure [PA91]. entails [Bel90d]. Entertainment [Zor97]. Entrepreneurial [Gep96a]. entrepreneurs [KP98]. Environment [GM93, Per93b, Zor93a, Bet99, Fer98, Fra96, Mag97, Per95b]. environments [WB97]. Envision [BDF99]. epidemiology [KWC93]. equals [Wis96].

Equation [Mar97a]. Equations [Ste96, Fos96c, Mir99a]. equipment [JJ91, Kur92, Per91]. Equivalence [McT96]. era [JPN^{+}92, Kha96]. Ericsson [Gut91]. error [Luc97c]. especially [GS97]. ESPRIT [Gut90b]. essay [BC93]. Essential [Len96]. estimate [LB96]. Ethernet [Dav97]. Ethical [Swe98a]. ethics [HL99]. Euro [Gro98]. Europe [Fox95, Gep97g, Gep98a, Joh98, Kur99, Swe94a, SB98b, Tan98]. European [PR92, PS98, Sti90a, SB98a]. EV [Ano96-53, Ano96-54, Ano96-55, Ano98-48, Ano98-49, Ano99-41, Rie97i, Rie97j, Rie97k, Rie97l, Rie98b, Rie98c, Rie98d, Rie99a, Rie99b, Rie99c]. Evaluating [Byr97, Ven90]. event [Bel90g, Bel91c, Per94e, Wat91a]. every [Pra94]. everyone [Tur91]. everything [GS97]. Everywhere [Wal96]. evidence [PG96]. evolution [CK92]. EVs [RW98, Toe98]. excerpt [BDF99]. exegesis [Mem97]. expanding [KR90]. expediency [Com93g]. experience [Swe94c]. expert [PA91]. Experts [BDF99]. Exploiting [Zim97]. exploration [Mar98a, WK90]. exploratory [Jam98]. Exploring [Pic95]. express [CDEJ99]. Extension [Hac93]. eye [Ano97-82, DM96, Lau97, WJT96].

fabrication [CB99]. face [Fos98a]. Faceted [Swe98c]. Facility [Doo96b]. Facsimile [Coo93]. factor [Luc96b]. factory [Ath98, Zor91d]. facts [Bel94c, Wic97]. failures [Ano96-56, Bra98, Eng96, Eng98, Fit90d, Gep96b, Gep97a, Gep97b, Gep99d, Gep99e, Obe97, Swe98b, Swe99b]. false [Coo93]. fantastic [Foc99]. Farouk [Ano96-71]. fashion [Gra98]. fast [FM92, FPR^{+}92, JPN^{+}92, Ng92, Pri94, Sal92]. Faster [Mar98a]. fat [CWG98]. Fatal [Luc97c]. father [Gep93, JNM^{+}98]. Fault [CCS99, VS97]. Faults [Ano96-56, Bra98, Eng96, Eng98, Gep97a, Gep97b, Gep99d, Gep99e, Obe97, Swe98b, Swe99b, Gep96b].

FCC [Col94]. FED [CWG98]. Federal [Ada90c]. feel [Ric92]. Fellow [Swe94c]. fend [LS98]. FETs [Ven90]. Few [AW98]. fibre [DR95]. fibre-optic [DR95]. Field [Ano97a, Ano98-62, Ano99-54]. Fieldbus [SZ96]. fields [Fit90c, Kir96, MN90, NM90, Per94f, SK92a, Swa91]. fighter [RGH90]. Filing [Eld93]. film [All97]. Finding [VS97, WL98]. finds [Cat90]. Fine [AAH98, Cha99, Gra99]. Fine-tuning [AAH98, Gra99]. Fink [Ano96-42]. fire [Zor91b]. Firewalls [LS98]. FireWire [Wic97]. firm [Gut92]. firms [Hac93]. First [Wis96]. Fish [Zor90c]. fitness [Bel96c]. fittest [YL94]. Fitting [Cla92]. Five [Ste96]. fix [Han93]. Fixed [DR99c, IO96]. FLAG [DR95]. Flash [DH93, Eld93]. flat [Wer93, Wer97]. Flatland [Win96]. flexibility [Fit90b]. Flexible [Hin93, Com95a]. flickers [Gut92]. flight [PG96, Swe95]. floating [IO96]. flood [Roe97]. floor [Ber97, Zor91d]. flops [Rod98]. flowering [Wer97]. flowers [SK92b]. flush [Pix96]. flying [Gep98b]. Flywheels [Hul97]. FMS [Bel93a, Luc93]. focus [Kas94]. focuses [Gom96]. Forbidden [Ano97-96]. Forcast [GS99]. Force [Swe98f, Bra96a]. forces [Ada90b, Per92b]. Ford [Bat95]. fore [Gil98]. forecast [Bel94b, BR97, Bet99, Bra99, Bre99, Com97b, Com98a, Com99b, Doo98, DR99f, Fer98, Fos98c, Gep95c, Gep97f, Gep98g, Gep99g, Hel98, Hin98, Kap97b, Kap98a, Kap99b, Kor99, Lan98, Lan99, LK99b, Mar99a, Mas98, Mas99c, Per98b, Phi98, RS99a, Rie98e, Ros98, Spa97, Swe98e, Swe99f, Trl97]. Forecasting [Rob97]. Foreign [SD94]. forfeit [Tro99]. format [Pen99]. Formula [RPD^{+}91]. Forum [Ano96-57, Ano96-58, Ano96-59, Ano96-60, Ano96-61, Ano96-62, Ano96-63, Ano96-64, Ano96-65, Ano96-66, Ano96-67, Ano96-68, Ano97-42, Ano97-43, Ano97-44, Ano97-45, Ano97-46, Ano97-47, Ano97-48, Ano97-49, Ano97-50, Ano97-51, Ano97-52, Ano97-53, Ano98-50, Ano98-51, Ano98-52, Ano98-53, Ano98-54, Ano98-55, Ano98-56, Ano98-57, Ano98-58, Ano98-59, Ano98-60, Ano98-61, Ano99-42, Ano99-43, Ano99-44, Ano99-45, Ano99-46, Ano99-47, Ano99-48, Ano99-49, Ano99-50, Ano99-51, Ano99-52, Ano99-53]. forums [Bat92]. fossil [Fra95b]. fossil-fueled [Fra95b]. Fouke [BDF99]. Four [Eve98]. FPGA [Faw94]. framework [Col93]. France [LLKS95]. frauds [Flo90]. Frederik [Ano96-69]. Free [Luc99b, Mor98a, Mil98a, SB99c, Tru98]. freeform [CB99]. frequency [Bar90, HT95, SS99]. Fresh [Gep96b, War96]. Friendly [Jur90b, Ano90e, Bil96]. Front [Ano96c, Ano96d, Ano96e, Ano96k, Ano96l, Ano96m, Ano96f, Ano96n, Ano96g, Ano96h, Ano96i, Ano96j, Ano97d, Ano97e, Ano97f, Ano97g, Ano97h, Ano97i, Ano97j, Ano97k, Ano97l, Ano97m, Ano97n, Ano97o, Ano98b, Ano98c, Ano98d, Ano98e, Ano98f, Ano98g, Ano98h, Ano98i, Ano98j, Ano98k, Ano98l, Ano98m, Ano99c, Ano99d, Ano99e, Ano99f, Ano99g, Ano99h, Ano99i, Ano99j, Ano99k, Ano99l, Ano99m, Ano99n]. Frontier [Mor98a]. frontiers [Bra95c, Kur99]. Fuel [Gil98, Wol99, FP93, Ito90, KHSK95, SSH99]. fuel-cell [FP93]. fuel-efficient [SSH99]. fueled [Fra95b]. Functional [LHD^{+}90]. fundamentals [Cox92]. further [Ano96-120, Ano96-121, Ano97-97, Ano97-98, Ano97-99, Ano97-100, Ano98-102, Ano99-95]. fusion [Swe94b]. Future [LP97a, FP93, Joh98, Kel97, Per97c, RN97, Rie98b, Sch97, Ste92, SB98a, Wer93, SLLP97]. Fuzziness [Kar96b]. Fuzzy [Cox92, SK92b, Ada96h, Cox93, Per95d, Sel90].

G [Ano96-42, Ano96-74, Ano96-124,

Ano97-38, Mor98a]. GaAs [Ven90]. gaining [BK99]. gains [ARA91, Fit90b]. Gale [Rob99]. Gallium [Cat90, Dey95]. game [Fit90h]. garden [Gep95d]. gateway [Hol97]. gathering [Lub95]. Ge [Cre95]. Gem [Swe98c]. generation [CDEJ99, Joh90, Kap94, Mas90, SSH99]. generations [Zor90b]. generator [Wol90]. generators [Bar90, Bla97]. geomagnetic [KA90]. George [Per95e, Hei97, Sha94]. geothermal [Zor92b]. Get [Tru98, BC98, Com96k, Zor91g]. gets [Mal93, YL90]. Getting [Bat95, Fuq96, Lau97]. ghosts [Jur92b]. giants [Dey95]. gigabit [CDEJ99, Dav97]. Give [Sou98]. giving [Gep96b]. glass [DR95, Swe95]. glitches [LHD^{+}90]. glitz [Luc96b]. Global [Fit90e, Mor97a, Ada90a, Dua96, Jur91b, LRC99, Obe99b, Rob99, Ste99, ZS90, Get93]. GlobalR&D [Jur90c]. globe [Soo94]. go [BH92, Box95, Com93c, Jur91b, SB98b, Wic96b]. goes [DH93, Ham95, SPM90, Zor94b]. Going [DRS96, Fri96, PS92]. gold [AW98, Bla96]. Golem [Fos99a]. Good [Doe97, Hin96f, Jur92b, Bel90a, Bla96, DR97, Hos94a]. Good-bye [Jur92b]. Good-looking [Doe97]. government [Ada92d, Ram94]. Graaff [Wol90]. grab [Fow94]. gracefully [Zor90a]. Graduate [Ano96-70]. Grand [Mar98c, Mos97, Nah92]. Graph [Mir95]. graphics [Fos92]. grass [LLKS95]. gravitational [Lew95, Sel96c]. Gravity [Fay99]. great [Flo90, Hin99, Rie92d, lH98]. Greatbatch [Ada95c]. greater [BK99]. Green [Per94c]. Greenhouse [Rob99]. grid [HS96b, JS99, Mas96, SB99b, Tay99]. grids [GR97, Leu97, RN97]. Grounds [St.92]. groves [Gup94]. Growing [Kuk93, Com96d]. grows [ER90]. grungy [Gup94]. Gruntwork [Fin97]. Guide [Ano96-72, Ano91b, Ano91a, DP91, GB91, HFYT91, RMAA91, Sha91]. Gulf [Ada91b, Rie91a, Zor91c, Zor91d]. guys [Sha93].

H [Ano96-42, Hei97, Sha94]. Hal [Ano97-38]. half [Tiw94]. half-micron [Tiw94]. Halperin [Ano96-71]. halt [Ada92e]. hand [Hos95b]. Handbook [Fos96a]. handling [Sch91]. happens [Hos94a]. hard [FASN96, Sto95]. hard-disk [FASN96]. hardcover [Swe98c]. harder [YL90]. Harnessing [Che90d, Jay95]. Harold [Ano96-104]. haul [Toe98]. haute [Com93b]. Haven [Ano97b, Kap98b]. Hawaii [Zor92b]. haze [LRC99]. HDTV [BCD^{+}95, Bel95b, Fox95, Jur91a, Lub97, Nin95, Per95c]. head [Com93c]. head-to-head [Com93c]. Heading [GR97]. Heads [Sto98]. Health [Fos97f]. heard [CMS97]. heat [Sou98]. heel [Mir99a]. Heilmeier [Hei97, Sha94]. hell [Kur92]. help [Ada92d, Hac93, Jur90b, Laz99]. Helping [Laz99, Fit90d]. helps [Eub98]. Henry [Per99b]. here [Luc96e, Mil94a]. Herman [Ano96-71]. heterojunction [Cre95]. HEV [Bat95]. HF [Hea90]. hi [Ros91c]. hi-tech [Ros91c]. hierarchical [Pen99]. High [Gep98b, Hin96h, RN97, SS99, Bar92, Bat92, Bla97, Cha97a, DF92, GB91, GGJ^{+}92, Kap98b, Kur92, Lub95, Ric92, RDK90, Smi92b, SS91, Tad98]. high-density [Lub95]. high-energy [Bla97]. High-flying [Gep98b]. High-resolution [SS99]. high-speed [Cha97a, GB91, RDK90]. high-tech [Bar92, Bat92, DF92, Kap98b, Ric92, Smi92b]. High-voltage [Hin96h]. Higher [Wer94, Fox93, Tad98]. highly [Mas99a]. highways [CW94, Jur91b]. HiQ [Fos97a]. Hiroyuki [Ano96-41]. hit [Zor90e]. hits [Rod98]. Hitting [Joh93]. Hoff [Per94d]. Holtzman [Fie97]. Home [BC98, DR99a, Ric92, Zor94b]. Homemade [Bha94]. homes [DR99d, Wol99]. hones [DU98]. Hong [Mec95]. Honk [Luc96d]. Honor [Hei97]. horizon [GON93, Hea90]. horse [Nor99]. hospitals [Wol99]. hot [San96]. hotter [Mal93]. house [Hos95a]. http [GS97]. Hubble [Bel91c, Han93]. human [Fos98a, NP90]. human-computer [NP90]. Humor [Bai96]. Hybrid [HS98, KHSK95]. Hybridizing [Bru94]. Hybrids [Wou95]. hydride [SOGC98]. hydroelectric [Bar90]. hydroelectricity [Zor90c]. hype [Car93, Edw97].

Ibuka [Ano96-76]. IC [Ano99a, Cha99, Fre98, Gep98c, Har99a, TC97]. ICs [Bal95, Ben98, Gep99c, Har92a, HFYT91, KS92, LY96, May94, PS92, SHM96, VS97, Wil93]. ideal [GAL98]. Identifying [SHM96]. identity [Mil94b]. IDL [Fox99]. IEEE [Ano99b, BDF99, Hei97, Ano96-72, Ano96a, Ano97-54, Ano98-63, Ano98-62, Ano99-54, Kuh91, Swe94c, Ano96c, Ano96d, Ano96e, Ano96k, Ano96l, Ano96m, Ano96f, Ano96n, Ano96g, Ano96h, Ano96i, Ano96j, Ano97d, Ano97e, Ano97f, Ano97g, Ano97h, Ano97i, Ano97j, Ano97k, Ano97l, Ano97m, Ano97n, Ano97o, Ano98b, Ano98c, Ano98d, Ano98e, Ano98f, Ano98g, Ano98h, Ano98i, Ano98j, Ano98k, Ano98l, Ano98m, Ano99c, Ano99d, Ano99e, Ano99f, Ano99g, Ano99h, Ano99i, Ano99j, Ano99k, Ano99l, Ano99m, Ano99n, Han96]. if [Luc96d]. ignite [Com98b]. II [Ada92e, Bel95e, Bel96e]. illusion [Mar98c]. illustrate [SB99b]. Image [Ede97]. Imaging [Fie97, Kir96, McC92]. Imamura [Ano96-103]. immediate [Gom96]. Immink [Ano96-76]. impact [HFYT91, Zor90b]. Imperative [Fos97f]. implants [NMGW96, WR96]. implications [Ada93a]. important [Gut90a]. impossible [Fef99]. Improved [Fos97b, Ada91a]. Improving [Fis92, Mos96, Sto99, Tay99]. improvisation [Bel90d]. in-house [Hos95a]. increase [Kin93]. Incredible [Bel91a, Tau99]. India [Bha94, SB99a, Zor94e]. indigestion [Gep96b]. indoors [WL98]. Industrial [All96, Gep94a, Kap95, Kap96, Kap97b, Kap98a, Kap99b, Kap90a, Kap91, Kap92b, Kap93, RWC91]. Industries [Ada90d, Sha93]. industry [Che90d, Fal96, Fra96, HS96b, Lyn93, Mag97, Per93a, Wer95]. Infinite [Hin99, Pic95]. Information [Ano96-103, But96, Luc99c, Ada91b, GE95, Gro98, HL99, Hil91]. Infrastructure [MW96, Roc97, Ros96]. ingenuity [Bel90d]. initiative [RWC91]. initiatives [Com93b]. injection [CCS99]. Innovation [Mor99a]. Innovations [Ano99-55, Gep97c, Gep97d, Gep97e, Gep98d, Gep98e, Gep98f, Gep99f, Jon99a]. innovative [Mac91]. input [Tur91]. insane [Hin99]. insanely [Hin99]. Insisting [Fef99]. Institute [Bel96a]. instrument [San90]. instrumentation [Fit90a, WBR91]. instruments [Fos96d]. integrated [Jur91d, PM98, Str96]. Integrating [Cai97]. integrity [Goy94]. intelligectual [HL99]. intelligence [Ada91a, Kok93]. Intelligent [Hay93, Pas95, Cla98]. interact [HW94]. Interactive [Com96c, Per96, Rid98]. interactively [Win96]. interactivity [Ano96-133, Lar95]. interchange [Jur91d]. interconnecting [Sha91]. interconnections [Wat92c]. Interesting [Ste98a]. interfaces [FPR^{+}92, NP90]. Interim [Swe97c]. internal [Gra99, Run99]. International [Ano96b, Bel96d, Bel96e, Ben93, Doo95b, Obe98]. Internet [Ada95b, And97, Ano98-89, Asp97, BCW97, CK99, CDEJ99, Com96c, Com96d, Com97c, Com99c, DRS96, DR99a, DR99c, Fos99b, GF98, JNM^{+}98, Lan98, Lan99, Oli99, Sir97]. Internet-based [GF98]. interrupt [Luc98d]. Introducing [Hin95]. invasions [LS98]. Inventing [Asp97]. Invention [Wis99]. inventor [Per95d]. investigation [Obe99d]. investigator [Ico97]. investment [McC97]. ion [Ben98]. ion-defying [Ben98]. Iraq [Zor92d]. ISBN [Rob99]. ISDN [DP91, Fra95a, Pet95]. isn’t [San90]. Israel [Wis99, Kap98b]. issue [Bel94b]. itself [McC97].

J [Ano96o, Ano96-52, Ano96-105]. Jack [Ano96-73]. Jan. [Lik92]. Janie [BDF99]. Japan [Cro91, KW92, MK90, Sai90, SK92b, Sha93, WK90]. Japanese [Nin95]. Java [Ath98, Che99, Hin96f]. jive [Hin96f]. Job [Luc98c, Hos95a]. Jobs [Bel93b]. jockey [Hin97c]. John [Ano96-91, Wis99]. Johnson [Ano96-103]. joins [Dey95]. Jr. [Per94d]. Judith [Ano96-74]. jukebox [Byr97]. Jumping [Ast95]. junction [CDEJ99, RH97]. junk [Luc97d]. jury [Fit90c, MN90, NM90]. Just [Bel90d, Yea96, Chr90, DR99b].

Kailath [Ano96-42]. kana [MK90]. Kang [Ano96-70]. kanji [MK90]. Karan [Ano96-123]. Kaufmann [Ano96-104]. Keeping [HSW95, HJ93, MA99, Bel90d]. keeps [Fri96]. Kees [Ano96-76]. Keithley [Ano90d]. Ketchen [Ano96-77]. Kevles [Fie97]. Key [Hin96f, Yof96, GS98]. Kilowatts [Bur93]. kitchen [Bel95b]. klatch [Hin96f]. Kluver [Mil98c]. knits [Lan94]. Know [Lin96, Fos99a, Joh93]. know-how [Joh93]. knowledge [Ano97-96, UT99]. knowledge-based [UT99]. Kobayashi [Ano96-75]. Koji [Ano96-75]. Kong [Mec95]. Kosko [Ada96h]. Koyanagi [Ano96-27]. Kuala [Rob95]. Kunishi [Ano96b].

L [Ano96-123, Doo94b]. laboratories [Ada90c]. Laboratory [San90]. labs [Col92]. Labview [Fos97d]. ladder [YL90]. land [CMS97, Fef99]. landing [Doo94b]. lane [Pri94]. language [Rho96]. LANs [Dav97, Sha91]. Laplace [Nah91]. laptop [Luc96c]. laptops [Tra93]. Large [Khe94, Khe95, Fos96c, Fos99a, GR97, Whi93b, Zor92f, Zor93b]. laser [For97, Lik92]. lasers [GON93]. last [GC94, RTHM92]. Launch [Bel96b, Doo96c]. Launcher [Doo96c]. Laverick [Zor92a]. Law [Tro99, Dil94, Rau93, Luc98e, Sch97].

Layoffs [GC94, Bel94c]. lead [Tru98]. leaders [SB98a]. leadership [JWA^{+}91]. leading [LD96]. lean [BC93]. leap [Rad96]. Learning [PB96, Wol92]. least [Pra94]. leave [GS98, Pel97]. LEDs [Wer94]. Leeds [Ano96-77]. legacy [Ada91b, Ano97-38, Rie91a, Zor91c, Zor91d]. Leland [Ano96-124]. length [Fay99]. less [DdMMNM96, KHSK95, SSH99, SH99]. Lessons [Bel98, Tab96, Obe98]. Let [Gal96]. lethal [Ada90b]. level [CMR96, Fox93, Fre98]. Liebman [Ano96-78]. Life [Luc98d, Wis99, Fef99]. Light [Bel90e, GAL98, Hau93, Mor91]. lights [MA99]. Like [Mar97a, Bel90e, Luc93, Mir95]. Lilliput [BLK98]. limit [HI92]. limits [YS91]. Linda [CS93]. line [DRS96, Kuk93, TC97]. lines [BH92]. link [DR95, GGJ^{+}92]. listens [CMS97]. lithography [FK91, Gep96c, Har99a, Zor92c]. little [Sha93, LP97b]. Lives [Sei99]. loads [DMP96, DL93]. Local [Wic96b, Bru94, GB91, Swe99a]. Local-area [Wic96b, GB91]. locally [Sto97]. Locking [BC97a]. Logic [DR91, Cla92, Per95d, SK92b, Sel90, Smi92a]. long [Ast94, Bel90c, Toe98]. long-term [Bel90c]. look [Com93d, Dix95, Gal94, Rod98]. Looking [Hea90, Hos94a, Doe97]. looks [Sha93]. loop [Bru94, Hin99]. loosening [Zor91a]. losers [Ste99]. losing [Fit90h, Ste99]. Losses [HS96a]. Lotfi [Per95d]. lots [SJ93]. Love [Pel97]. Low [Bla97, Div96, Fit90b, Fre98, Jul94]. Low-calorie [Bla97]. low-end [Jul94]. low-power [Fre98]. Lumpur [Rob95]. Luttwak [Ste99].

M [Ano96-71]. M. [Per95e]. machines [Hin97c, MWP97]. macros [Gra99]. made [Cer92, Fos93]. magazine [Hei97]. Magnetic [Woo90, Bel93e, Ede99, Fri96, Kir96, MWP97, Per94f]. Magnetism [Ryd91]. magnets [Ryd91]. mail [Ada92c, Per92b, PA92, Per92a]. main [Bel90g, Wat91a]. mainframes [Zor90g, Zor91e]. mainstream [DH93]. Maintain [Eng97a]. Major [Fos96b]. make [Apf97]. makes [ARA91, Bar92]. Making [DSD96, HK99, BBH97, GE95, Gut90a, Kuh91, Ric92, Tur91, LP98]. management [Bel93f, Fit91b, Fit91a, Hei97, Per90e, Swa93, Zor91b, Zor91a]. Managing [Goy94, Pfl96, HPV92, MN90]. Mani [Ano96-75]. manipulation [Mas99b]. Manna [Day97]. manufacturability [Sto99]. manufacture [DU98, Har99a, Kap97a, PP93]. Manufacturing [JT93, May94, Ben93, EJ93, Gut92, Hay93, Joh93, Jur90c, Sha93, Whi93b]. Many [Swe98c, GS98]. Many-Faceted [Swe98c]. march [BPM94]. Marcian [Per94d]. Marco [Ano96z]. Marcus [Ano96-104]. marine [Doo96c]. Mark [Ano96-77, GS98, Tiw94]. market [BH96, Che90f, Fra95a, Spi96]. marketing [Gut90a]. marketplace [Ada90c]. Mars [Obe99d]. Masaru [Ano96-76]. mass [BC93]. Master [EJ93]. Masterson [War96]. Math [Bra95c, Fos97d, Fos99b, Kor99, Swe98g, Fer91, Fos91, Fos92, Fos93, Fos96b]. Mathematical [Rid94, Mir96b, Swa98]. Mathematics [Ste96, Che90e]. Matlab [Mir96c, Mir97]. Matlab-to-C [Mir97]. Matrices [Fos97c]. matter [Com99h]. mature [Zor97]. Maxwell [Mir99a, Nah92]. may [von94]. Mb [Wat91a]. McCluskey [Ano96-52]. Measure [MC94]. measurement [Fit91d, Gro93, Mas98, Rie92e, Rie94, Rie95b, Rie96d, Spa97]. measurements [Ede99, Mas99c]. measures [Gel96, Whi93a]. Measuring [Pfl92]. Mecha [Com94b]. mechatronics [Com94b]. Medal [Hei97]. Medalists [Ano96a, Ano97-54, Ano98-63, Ano99b].

Media [Mar97a, Bel91c, Com96c, Per94e]. Medical [Ada94a, Ada95a, Ada96i, Fie97, LK99b, Mor97d, Ros98, Fit90g, Fit91c, Jur92d, Ste93]. medicine [Aka97]. meet [Ada90c]. Meeting [Spi96]. meetings [DR98]. meets [Oli99, Per90e]. megabits [Woo90]. Memoir [Ano97b]. Memorial [Ano96-78]. memories [Ng92]. Memory [FM92, FPR^{+}92, GGJ^{+}92, JPN^{+}92, Ng92, Pri94, Sal92, DH93, Gra99]. mentor [Hos95b]. messiah [Swa98]. metal [Joh94, SOGC98]. Metcalfe [Swe96d]. Methodology [Yof96]. metrics [Mog93, Rya98]. Metropolitan [MC90]. Metropolitan-area [MC90]. Mexico [Bea97]. Michivuki [Ano96-69]. microcontroller [Gil91, VG90]. microcontrollers [Kha96]. Microelectronics [CZ90]. microgyroscope [TLN98]. Micromachines [BPM94]. micromechanics [HMGT90]. micron [Tiw94]. microsystems [PM98]. microwave [Ven90]. Migliaro [Ano96z]. MIL [Wat92a]. military [AE90, AE91, Bra99, Doo93a, Doo94a, Doo95a, Doo96a, Doo97a, Doo98, KRF97]. millennium [BDF99, Gep96c, CP98]. million [Ano99a]. Mind [Alt96, Mac91]. minds [Swe99d]. Miniature [MWP97]. Mining [Wil97]. minis [Zor90g, Zor91e]. Minting [CB97]. Minus [Fin97, Edw97]. Mir [Mar98e, Obe98]. mirror [GLS94]. Mirrors [You93]. Misdirected [Mor96]. Misguided [Day96]. missile [Day96, LP97a, Mos97, SLLP97]. Missiles [Obe99b]. missing [Jon92]. Mission [Ano97c]. misunderstood [Sin96]. Mitsumasa [Ano96-27]. mixed [Har92a]. mixed-signal [Har92a]. Mixing [BT93, Wil93]. Mo [Ano96-70]. Mobile [Rei92b, Hil99, Joh98, Mil98a, Mor98b, Oli99]. mode [Luc98d]. model [TLN98]. modeling [Gro92, Per93b]. models [Col92, Jur90d, Ven90]. modern [Ros96]. modernisation [Ros96, Sto95]. modules [Fli94, Joh90]. Molding [Hau93]. Molecule [Bry96, Bry96]. Money [And97, Kel97, Per97b]. monitoring [HSW95, Swe96a]. monopoly [Bel90c]. monster [Zor99]. Moore [Luc98e, Sch97]. Morris [Ano96-77, Ano96-78]. Morton [Ano96-73]. most [Hin99]. motorist [Jur95]. motors [Bla97]. mounts [PG96]. move [Mor98b, Zor90d]. movement [Win96]. Moving [Ath98, Dua96]. MPEG [Koe99, Noa97, Swe97b]. MPEG-2 [Noa97]. MPEG-4 [Koe99]. much [DR97, Fos97c, Mun96]. multi [Fre98]. multi-level [Fre98]. Multichip [Joh90, Dai92, Fli94]. Multikilowatts [Van96]. Multimedia [Ada93a, Col93, ZF94, Bel93e, Com94c, Koe99]. Multimedia-applications [Ada93a]. Multimedia-the [Col93]. Multinationals [Zor90d]. Multithreaded [BH95]. Murad [Ano96-124]. Muscle [Van96]. museums [Bel95d, Bel95e, Bel96d, Bel96e]. music [Par97]. musical [Bar98]. mutual [SB98a]. myths [Bel94c].

N [Ano96-78, Swe96d]. Naked [Fie97]. named [Mor91, Sel97e]. Nano [Bry96]. nanometer [Sha99]. Nanotechnology [Bry96]. NASA [Bel96b]. Nasdaq [Ber97]. national [Ada90d]. natural [Bel90e]. Navigating [Doo94c]. navigation [Get93]. necklace [DR95]. Needed [Mat90]. needs [Gom96, Jur90e]. Negotiating [Ove97]. Net [LS98, Ada92c, Ada96a, Bil96, BCW97, Lan94, LLKS95]. nets [Ada96a]. Network [Len96, Waf93, Com97a, Dix95, HPV92, Lar95, Spi94, Sto95]. Networking [Sto97, Hil91]. Networks [DR99d, Ada96h, CS93, Cru99, GB91, Ham93a, Ham93b, Jur91d, MC90, Raj94, Swe96a, Wic96b, Yal97]. networks/fuzzy [Ada96h]. Neural [Ham93a, Ada96h, Bil96, Ham93b, Jur91d, May94, Yal97]. neural-net [Bil96]. Neuromorphic [KM96]. Nevada [Bea97]. news [Bel90a]. Newslog [Ano96-79, Ano96-80, Ano96-81, Ano96-82, Ano96-83, Ano96-84, Ano96-85, Ano96-86, Ano96-87, Ano96-88, Ano96-89, Ano96-90, Ano97-55, Ano97-56, Ano97-57, Ano97-58, Ano97-59, Ano97-60, Ano97-61, Ano97-62, Ano97-63, Ano97-64, Ano97-65, Ano97-66, Ano98-64, Ano98-65, Ano98-66, Ano98-67, Ano98-68, Ano98-69, Ano98-70, Ano98-71, Ano98-72, Ano98-73, Ano98-74, Ano98-75, Ano99-56, Ano99-57, Ano99-58, Ano99-59, Ano99-60, Ano99-61, Ano99-62, Ano99-63, Ano99-64, Ano99-65, Ano99-66, Ano99-67]. Next [BDF99, Rob97, CDEJ99, Gep96c, Joh90, JS99, Mas90, Swe96e]. next-generation [CDEJ99, Joh90, Mas90]. niche [Cat90]. Nickel [SOGC98]. Nikola [Ano96-91]. NJ [BDF99, Fie97]. No [Lin96, Gep95d]. Nobutake [Ano96-103]. node [Sel97e]. noise [Ell99a]. Nokia [Gut92]. nonlinear [Ven90]. Norway [Tab96]. Nosaka [Ano96b]. nose [BLK98]. noses [NGOS98]. Nosferatu [Sel97e]. Notes [Hin96a, Hin96f, Hin96b, Hin96c, Hin96d, Hin96e, Hin97a, Hin97b]. Nuclear [Bea97, But96, CKR94, Swe94b, ZS90, Ada92e, Cro91, Fel97, Kal94, Mar98c, Swe97d, TPB95, Zor92d]. number [Luc97c]. Nurturing [MT96].

O [Ano96-104, Per98a]. object [Atw91]. object-oriented [Atw91]. observers [PA91]. o’clock [Luc97b]. Ocular [WR96]. off [GR97, LS98, Obe99d]. Often [Mos96]. Ogura [Ano96-78]. oil [Zor91a, dMA96]. Old [Zak99, Cam97, Ito90, Per91, Swe99e]. on-campus [Fal96]. once [RN97]. One [Mas99a, SB98b, FM92, HBY90, SJ93]. one-stop [HBY90]. only [Com96d]. ons [Ano91a, Com92a]. Open [PA97, Com99a]. open-source [Com99a]. Opening [Tan98]. operating [Com93c]. opportunities [IK99]. optic [DR95]. Optical [Ell96, FK91, Ast94, AF95, Wil97]. optically [Dav97]. Optics [Ell96]. options [Kin93]. optoelectronics [Che90a]. orbit [Ben98]. orbiting [Cam90]. order [Bur93, LK99a]. oriented [Atw91]. origin [Fit91a]. origins [RH97]. oscilloscope [Mur96]. oscilloscopes [RDK90, RLH93]. other [Soo94, Sur90, Wic96a]. others-much [Mun96]. our [GS97, Koe99]. outpost [Doo95b]. overhauling [Fit91b]. overhead [Fit90b]. overload [Dal90]. overnight [Ast94].

P [Ano97b]. Pack [Fuq96]. package [Her99, Mas99a]. packages [Joh90, Mir96b]. packaging [Tho97, Wat92c]. pagers [SJ93]. pains [Com96d]. paints [SLLP97]. Pan [PS98]. panacea [Car93]. panel [Wer93]. paper [Swe98c]. parallel [CS93]. parallelism [Zor92e]. Paris [Sti90a]. particles [Fay99]. partner [Soo94]. Partners [TT95]. Pascal [Mor98a]. past [Sch97]. Pasteur [Mor99a]. path [Com99a, FM92]. Patriot [Rie91a]. Patterson [Ano96-122]. Paul [Ano96-74, Wis99]. pay [DdMMNM96, Zor90e]. PC [Edw97, Ste92]. PCB [Goy94]. PCs [Com93a, Com93e, Com94a, Com94d, Ros91b, Ros92b, Voe90a]. Pederson [Per98a]. Peenemünde [Ano97b]. peep [Swe98d]. People [Mar97a, CKR94, Luc99a, Swa93, WL98, Whi93b]. performance [Byr97, DF92, LB96]. peripatetic [Zor94b]. peripherals [Kin93, Voe90b]. Personal [Kob93, San96, Col94, Com96k]. personalized [Bel93a]. Perspective [Get93, Kel97, PS98, Rod99]. Perspective/navigation [Get93]. perspectives [AL92, Gep99a]. persuasive [PA92]. pervasive [PA92]. Peter [Ano97b]. phantom [Ano90d]. Philips [Ano96-69]. Philosopher [Swe99e]. phone [Bel90f, Fis93, Flo90, Mil98a, Oli99, Pra94, Swe99a]. phones [Zor90e]. photonics [Hin92, Low97]. photovoltaic [Kap94]. Photovoltaics [BK99, Str96]. Photovolts [Lys94]. picking [Mur96]. picture [LP97b, SLLP97, WH99]. pilots [RGH90]. Pioneering [Rud96]. Pioneers [Bro99]. Piore [Ano96-52]. Piscataway [BDF99]. pitfalls [Swe98d]. place [Sto95]. placement [Dai92]. Places [Mar97a, Luc99f]. plague [Luc97d]. Planck [Fay99]. plane [DU98]. planetary [But97]. plans [Mos97]. plant [BP98]. plants [Fra95b, TPB95]. Plassche [Ano96-105]. platform [Doo96c, Dua96, TT95]. Platforms [Com93c, Gep93]. play [DU98, Par97]. playing [Ada92c]. plethora [Per90d]. Plug [DU98]. plus [Per90c]. pn [RH97]. Pocket [Com98b, SJ93, Fan97]. Poetry [Ste96]. Point [Luc98f, IO96]. points [Mur96]. Poised [JWA^{+}91]. Poison [Che99]. police [Fis92]. policy [Dav94, Day96, HS99, Mor96, Mor99b]. Polished [Fos95]. politics [HS99]. pollutes [KHSK95]. polluting [RTD97]. pollution [Gel96, LRC99, WH99]. polymers [SS92]. ponder [Mur96]. pornography [Ano97-96]. portable [PG96]. position [Hin97c]. Positioning [Get93, WL98]. Posix [Kuh91]. postmodern [Sel96c]. Power [Bal95, But96, DdMMNM96, Her99, Luc98f, RTD97, Ste96, Str96, Swe96c, Swe97e, Swe98e, Swe99f, Zor90f, Zor94c, Zor95a, Bal97, Bal91, BP98, Bur93, CKR94, Cro91, DL93, Fra95b, Fre98, Hin95, Hin96h, Ito90, Kap94, Kap99a, Leu97, Mar98c, Mar90, MT96, MA99, PR92, Rad96, Ram94, Rid94, SS96, SSH99, SB99a, SW97, Toe98, TPB95, ZF94, Zor90h, Zor91f, Zor92e, Zor92g, Zor93c]. Powerful [Fal96]. powerhouse [BP98]. powerhouses [Che90c]. pp [Ano97b, Fie97, Mor98a, Rob99, Swe98c, Wis99]. practical [Ano96-92, Ano96-93, Ano96-94, Ano97-67, Ano97-68, Ano97-69, Ano97-70, Ano98-76, Fos93, Fre98, Gra99, LP98, Sou98]. practices [Ben93, RWC91]. precollege [AW98, Che90e]. Prepackaged [Fos91]. Preparing [Pow96, Yal97, Smi92b]. prescription [Dal90, Mor96]. present [Sch97]. president [Ber97]. Press [Ano97b, BDF99, Fie97, Mor98a, Swe98c]. pretty [Fos97a]. prevention [Mor97c]. pricing [Log99]. printed [Ber99]. printed-circuit [Ber99]. priorities [Gep94a, PA91]. priority [Ram94, Sou98]. Privacy [AHP^{+}95, Sto98, Ada92a]. privatization [DSD96]. Prize [Ano96o, Ano96-42, Ano96-124]. PRML [FASN96]. Pro [Mir95]. Proactive [Tad98]. probe [Ano96-120, Ano96-121, Ano97-97, Ano97-98, Ano97-99, Ano97-100, Ano98-102, Ano99-95, Obe99d]. Problem [AHP^{+}95]. problems [Per90e]. process [LD96, SZ96]. processed [Ano90c]. Processing [Ede97, BDG96, CS93, MK90, PF96]. processor [BH95, LB96]. prodigy [Zor94b]. production [Ada92d, BC93, Fit91b, Kuk93]. products [Zor90b]. profession [Hos94a]. Profile [Hor92, KWCR92, Per92c]. Program [Asp90, Hin96f, Hin96a, Hin96b, Hin96c, Hin96d, Hin96e, Hin97a, Hin97b, Obe95, Zor92b]. programmable [Cla92]. programmer [Ell99c]. programs [EJ93, Hac93, Mir96c]. progress [Kuh91]. project [Bel90g, Mec95, Swe94a]. projection [Har99b]. proliferate [Ano90a]. proliferation [Ada92e]. Prometheus [Ano97-96]. promise [Ada91a]. promotion [Hos94b]. property [HL99]. Protecting [Mar90]. protection [Leu97]. protections [Tro99]. Proteus [Ano96z]. protocol [SZ96]. protocols [McM96]. prototyping [Com93d, Han93]. pursued [Che90c]. pursuing [Rie92b]. push [ZF94]. Pushing [YS91]. put [Mas96]. Putting [Jur90d, WS93].

Q&A [Hei97]. Quadrant [Mor99a]. Quality [Mos96, Pen99, SS96, Bel93f, Smi93]. Quantum [Ell96, HI92, Sel96c]. quantum-gravitational [Sel96c]. quarter [Doo94b]. quick [Com93d]. Quilt [Han96, Han97, Han98b]. Quilting [Han98a].

R [Ano96-52, Ano96-124, Ano97-96]. R&D [Bel90d, Bra96a, DF92, DdMMNM96, Gep94a, Gom96, Jur90e, KR90, Mar98b, MT96, Mur90c, PS98, SD94, SB98a, Yea96]. R&D-wise [Mar98b]. Race [Yof96]. races [Per91]. Radar [Bro99, But97, Fis92, Hea90]. Radio [Zor95b, HT95]. RAID [Fri96]. rail [Mor91, Sto95]. railway [Kur99]. rainbow [RT98]. RAM [GGJ^{+}92, Mas90, Wat91a]. RAMs [JPN^{+}92]. rapid [Com93d]. raring [SB98b]. rational [Noa97]. ray [Dal90, Zor92c]. rays [FK91]. Re [Cre95, Mas95, Sto98]. Re-engineering [Cre95, Mas95]. reaches [Fra95a]. reactor [Swe97a]. Reading [Mar99b, Hin96f]. Ready [All96, Mas90, SOGC98]. Real [Mar97a, Ada93b, Rya98]. real-world [Rya98]. reality [Ada93b, Com96c, Gal94, Gro92, Ano97-38]. really [San90]. realms [Pic95]. reasonable [DMP96]. rebounding [Swe97a]. Recipe [BM94]. Recipes [Mac98]. recipient [Hei97]. recognition [CMS97]. Reconciling [Mur90b]. reconnaissance [Swe97d]. record [SS99]. recording [Ham95]. recovery [dMA96]. Recycling [MH98, BM94]. Redesigning [Cam97]. reduce [Sou98]. redux [Luc98e]. referencing [St.92]. reflect [SB98a]. Reflections [Luc94a, Luc94b, Luc96a, Luc96b, Luc96d, Luc96c, Luc96e, Luc97b, Luc97c, Luc97d, Luc97e, Luc97a, Luc97f, Luc98a, Luc98b, Luc98c, Luc98d, Luc98e, Luc98f, Luc99a, Luc99b, Luc99c, Luc99d, Luc99e, Luc99f]. reform [Rud96]. Refreshing [Wat92b]. refrigerators [Per94c]. regime [MT96]. register [CMR96]. regulator [Kel97].

Reich [Ano97b]. Relationship [Obe99c, And97]. reliability [JS99, MA99, Pfl92, SS96, Wat92a]. reliability-data [JS99]. Remaking [Bry96]. Remote [Bel95c]. renaissance [Gep95b]. repeater [Ano90d]. replace [Per91]. reply [LP97b, SLLP97]. Report [Bal97, Kap90b]. repose [Fit90f]. requirements [Spi96]. Requires [Mos96]. Research [Doo95b, Bel90c, Che90d, Com93b, Fit90b, Hei97, McT96]. residential [Hol97]. Resistance [But96]. Resnik [Ano96-74]. resolution [SS99]. resonance [MWP97]. resource [Per95b]. resourcefulness [Bel90d]. resources [Cat92, Lar93]. responding [Jur90e]. responsibility [Smi92b]. restructuring [BH96, HS96b]. Rethinking [Zor92c]. reuse [Dav94]. reverberations [Fit90c]. reverse [Rau93, Zor92d]. reverse-engineered [Zor92d]. Reversing [Che90e]. Review [Bro99, Day96, Ede97, Ede99, Eng97a, Fay99, Fie97, Gad98, Kam99, Mac98, Mar97a, Mor99a, Pac99, Rob97, Rob99, Rod99, Ste98a, Swe98c, Wis99, Ano97b, Ano97-38, Ano97-96, Bai96, Bel96b, Bry96, But96, Ell96, Fos96a, Fos97f, Hel97, Mor98a, Pfl96, Ste99, Wis96]. Reviews [Bro98, Cla98, Cru99, DRE97, Ell99b, Eub98, Fos97b, Fos97c, Fos99a, Fox99, Gum99, Mar96, Mas99a, Mir98, Mir99a, Mir99b, Mor99b, Ano99-80, Ano99-81, Bil96, Dom94, EW97, Eng97c, Fos96b, Fos96c, Fos97e, Fos98b, Fos98d, Fos99c, Mir96a, Mir96c, Pad98, Pen99, Zor97]. revising [Rie91a]. revisionist [Rod98]. Revolution [Bli96, CK92, Hel99, Swe98c]. Reynold [Ano96-103]. Richard [Ano96-104]. Right [Gal96, Ung96, Hou95, Jam98, LP97b, Nor99, Yal97]. rights [Mar99b]. rise [WB97]. rises [Gep98a]. rising [Gut92]. risk [Bel93b, Fit90d]. risk-taking [Fit90d]. risks [Mar98c, MN90]. Risky [Bel96b]. rivalries [Zak99]. Rizk [Ano96-71]. Road [RW98, Ast94, Bat95, Joh93, Waf93]. Robert [Ano96-73, Swe96d]. robotics [WK90]. role [Har99b, Luc98b, RN97]. roll [Hul97]. Rolling [Kur99]. room [DSCN93]. root [Nat95]. roots [LLKS95]. rose [Gep95d]. Rothman [Fos97f]. Roundtable [Ano98-89, Com99c, Bel96c, Com96d, Com97c, Swe98a]. routes [DR99c]. routing [Dai92]. rubicon [Gro98]. ruckus [Sel96c]. Rudy [Ano96-105]. Rule [CMR96]. rules [Gre94, Shi91]. Running [Lar93]. rural [Bha94]. Russia [Obe95]. Russian [Obe99c]. Rutgers [Fie97].

S [Ano96-124, Ast95]. S-curve [Ast95]. saddle [Bal95]. safe [BC97a, PA91]. safeguards [Ada92e, Zor92d]. safety [Ada91a, Per93a]. sagging [Che90e]. sailing [Doo96d]. Sakaki [Ano96-41]. sake [Mil98d]. sale [Fil96]. Salvageability [Dil94]. Samueli [Per99b]. sanctioning [Fit90d]. Sarnoff [Ano96-41]. satellite [Doo96c]. Satellites [Mil98a, Ada91a, Cam90]. saver [Mir97]. Sayed [Ano96-42]. scale [Tho97]. scare [Fis93]. scene [Nin95]. school [Swe99e]. schools [Mir96b]. Schouhamer [Ano96-76]. Science [All96, Bel96a, Bry96, Mes98, Mor97a, Mor99a, Swe96e, Bel95d, Bel95e, Bel96d, Bel96e, Che90e, Hel98, Hel96, Hel97, Kor99, Mor99b, Mur90c]. Scientific [Swe98c]. Scientists [Fos96a]. scopes [DT92]. script [Rie91a]. sea [Mec95]. search [Per97c, Rie95a, lH98]. second [Com96d, DR99e, Wil97, Han97]. Secrets [Kal94]. securities [Wal97]. Security [Len96, Ada92a, Ada92b, BCW97, BP92, Che99, JS99]. see [PB96]. Seeing [Ano97-82, Kor93]. Seeing-eye [Ano97-82]. seek [Zor90a]. Seeking [Ada92e, Zor92d]. Seiki [Ano96-78]. select [Gil91, VG90]. Selecting [Fer91]. Semiconductor [Gep96c, SC90]. Semiconductors [Gep99g]. sense [GE95]. senses [Wic96a]. sensing [Bel95c, Zor93a]. sensors [HMGT90, McC92]. sentries [Doo97b]. serial [Fos98d, Wic97]. serious [Zor91g]. serve [SOGC98]. server [Lar94]. servers [Nat95]. service [Pet95]. services [Per90d]. Set [Eng97a, Cic95]. set-top [Cic95]. setbacks [KA94]. sets [Fos96c]. setting [Ano91b]. seventh [Nor95]. Shannon [Hor92]. Shapin [Swe98c]. shared [WB97]. shares [Hei97]. Sharing [BC97b]. sharper [Sal92]. Shattuck [Ano97-96]. ship [Doo96d]. shopping [HBY90]. shore [Doo94c]. Shoreham [Fit90f]. should [Fos99a, GC94, Zak99]. shrinking [Bel91a, Tau99]. shut [Mas96]. Shuttle [Bel98, Obe98]. Si [Cre95]. side [Soo94, Sur90]. Sides [Swe98f]. Siemens [Gut90a]. Sights [Com96e, Com96a, Mil98e, Ano96-128, Ano96-129, Ano96-130, Ano96-131, Ano96-132, Ano99-96, Ano99-97, Ano99-98, Ano99-99, Ano99-100, Ano99-101, Com96f, Com96g, Com96h, Com96i, Com96j, Com97d, Com97e, Com97f, Com97g, Com97h, Com97i, Com97j, Com98c, Com98d, Com98e, Com98f, Com98g, Com98h, Com98i, Com99d, Com99e, Com99f, Eng97b, Mil98f, Mil98g, Mil99b]. sigma [Smi93]. signal [BDG96, Goy94, Har92a, PF96, St.92]. signals [BT93]. significant [Mun96]. signs [Mil94b]. Silicon [HMGT90, Cre95, Gep98a]. simpler [Cer92]. simulate [TLN98]. simulated [Cru99, JT93]. Simulating [Swa91]. simulation [Sha99]. simulators [Cen90, Fra95b, KRF97]. Single [Mar97b, Che90f]. single-market [Che90f]. sinks [Sou98]. Site [SB99b, JB96, Ove97, Eng97a]. Sites [Ano97-101, Com97k]. Six [Smi93, Gep99a]. Six-sigma [Smi93]. skies [Mec95, PA91]. skills [Che90e, Wol92, YL94]. sky [Doo95b]. skyscrapers [Ish94]. slant [War96]. Sleuthing [Lew95]. slow [Zor94d]. Slowing [TPB95]. Small [Jul95, DMP96, Hac93, SS99, Sky94]. smaller [Ada90b]. Smart [CW94, Jur91b, Raj94, dMA96, von94, BC98, HK99, Luc97f]. smartcards [Fan97]. smartens [Mas99b]. smelling [GW98]. Smoother [Doo96d]. social [Per92b, Roc97, Swe98d]. Societal [Fit90c]. Software [Ano99-80, Ano99-81, Bro98, CS93, Com92d, Com93f, Com95c, Com97b, Com99b, Cru99, DRE97, EW97, Ell99b, Eub98, Fos96a, Fos97b, Fos97c, Fos97e, Fos98b, Fos98c, Fos98d, Fos99c, Fox99, Gab94b, Gum99, Hin96g, Hin98, HBY90, Mas99a, Mas99b, Mir96c, Mir98, Mir99a, Mir99b, Mos96, Mur90c, Pac99, Pad98, Pen99, Pfl96, Rod99, Rya98, Trl96, Trl97, Ano90a, Ano91c, Apf97, Bil96, Bra95a, Cer92, Com92c, Dai92, DU98, Dua96, Fos92, Fos93, Fos96b, Fos96c, Gab94a, Gil91, Har92a, Hou95, Jon98a, Kap92a, Kap99c, Kor93, Kor94, Kor99, Luc99b, Mat90, Mir96a, Mir96b, MP97, Mor98b, MV90, Pfl92, Rie93, Rob92, Roc97, SK92a, Smi92a, Top91]. Software/parallel [CS93]. SOI [All97]. solder [Tru98]. soldiers [KRF97]. Solid [CB99, Gep94b, Gep95c, Gep96d, Gep97f, Gep98g, Gep99g, Cha97b, FP93, SC90, Sel93, Wat91b, Wat92d, Ano96-105]. solid-state [Cha97b, Ano96-105]. solitons [Bel90e, Hau93]. Solver [Fos96c, Mir96a]. some [Bel90a, Swe98d]. sound [AC97, Bar98]. source [Com99a, GAL98, Ito90]. South [Rud96]. Soviet [Che90a, Kal94]. Space [Bel91c, Day97, Doo95b, Doo97b, Obe95, Obe98, Obe99c, Tay92, Waf93, BE91, Doo95c, LK99a, Luc94b, Mar98a, Mar98c, WK90, AAH98]. Speaking [Sel96c, Sel98a, Sel98b, Ano96-118, Ano96-119, Ano99-94, Sel96a, Sel96b, Sel97e, Sel97a, Sel97b, Sel97c, Sel97d, Sel98c, Sel98d, Sel98e, Sel98f, Sel99a, Sel99b, Sel99c, Sel99d, Sel99e]. Speakout [MW96, Mor97b, RS99b]. Spec [Apf97]. Spec-based [Apf97]. Special [Bal97, Luc99f, Bra95b]. specialities [Che90b, Jur91d, Jur92e]. specialties [Bel93e]. species [Wic96a]. Specifying [Hen97]. spectral [JJ91]. Spectrum [Ano96-29, Ano96-30, Ano96-31, Ano96-32, Ano96-33, Ano96-34, Ano96-35, Ano96-36, Ano96-37, Ano96-38, Ano96-39, Ano97-28, Ano97-29, Ano97-30, Ano97-31, Ano97-32, Ano97-33, Ano97-34, Ano97-35, Ano97-36, Ano97-37, Ano98y, Ano98z, Ano98-27, Ano98-28, Ano98-29, Ano98-30, Ano98-31, Ano98-32, Ano98-33, Ano98-34, Ano98-35, Ano98-36, Ano99-27, Ano99-28, HT95, Rad96, SPM90, Han96, Ano96-28, Ano96c, Ano96d, Ano96e, Ano96k, Ano96l, Ano96m, Ano96f, Ano96n, Ano96g, Ano96h, Ano96i, Ano96j, Ano97d, Ano97e, Ano97f, Ano97g, Ano97h, Ano97i, Ano97j, Ano97k, Ano97l, Ano97m, Ano97n, Ano97o, Ano98b, Ano98c, Ano98d, Ano98e, Ano98f, Ano98g, Ano98h, Ano98i, Ano98j, Ano98k, Ano98l, Ano98m, Ano99c, Ano99d, Ano99e, Ano99f, Ano99g, Ano99h, Ano99i, Ano99j, Ano99k, Ano99l]. Spectrum [Ano99m, Ano99n, Han97, Han98b, Hei97]. speech [CMS97]. speed [Cha97a, Faw94, GB91, GGJ^{+}92, Luc98a, RDK90]. Speeding [Kar96a]. spice [Rid98]. spin [Joh94]. spot [CCS99]. spot-checks [CCS99]. Spread [Rad96, SPM90]. springboard [Hos95a]. spur [Fit90d]. Spying [Obe99a]. squeeze [LC99]. stalls [FK91]. Standard [Ano97-83, BCD^{+}95, Dav97, YS91]. Standards [Mem98, Fit90e, Mur90b]. star [Gep98a, Gut92, Kel99]. stars [AW98, Bel95a]. start [Zor94d]. starts [Coo93]. State [Ano98-89, Com97c, Com99c, Fel97, Cha97b, Com96d, Gep94b, Gep95c, Gep96d, Gep97f, Gep98g, Gep99g, SC90, Sel93, Wat91b, Wat92d, Ano96-105]. States [BE91, Boo99, SD94]. Static [Cha97a]. Station [Doo95b, Obe98]. Statistics [Doe96]. status [Kap90b]. Staying [Fuq96, Par96, Ste98a]. steer [McA98]. Steers [Ano97-82]. Steffes [Ano96-74].

Stein [LP97b]. Steinmetz [Ano96z]. step [Kob94]. steps [MC94]. Steve [Ano96-70]. Steven [Swe98c]. still [Fit90c, MN90, NM90, Sym98]. stock [Ber97]. stop [HBY90]. Storage [Ano96-103, Ast94, AF95, Bel93e, DT92, Fri96, SC90]. Stored [Asp90]. stories [WBR91]. Storing [DL93]. Stork [Ano97-38]. storm [Lub95]. storms [KA90]. Stormy [Doo95c]. story [Ede99, Rob99]. straight [Mem97]. strange [And97, Wic96a]. Strategies [MF97, Che90c]. street [Zor90e]. streetcar [Mor91]. streets [HS98]. stress [Div96]. stressful [Swe96b]. Structuring [Sta93]. struggle [Gep95d]. struggles [SB99a]. submarine [DR95]. submicron [LY96, SHM96, VS97]. substation [Cai97]. success [Ast94, Com99h, Hos95a, Mac98, MC94, WBR91]. successes [KA94]. Suite [Eub98]. sun [Mak94]. Sung [Ano96-70]. Sung-Mo [Ano96-70]. Super [Bal97]. superchips [Doz92]. supercollider [Swe94a]. supercomputer [Gep99b]. Supercomputers [AL92, Bel92a, Com92c, CK92, Doz92, Gro92, Wat92c, Zor92e, FT93]. Supercomputing [Cat92, Doz92, KW92]. Supercomputing-Japan [KW92]. Superconducting [PC98, Mal93, MAW97]. superconductivity [RKGB92]. superconductors [HHSM93]. Superdense [AF95]. suppliers [Zor90a]. supplies [SS96]. Supplying [DMP96]. support [Che90d]. sure [Apf97]. Surge [Leu97, PC98]. Survival [Bel96c, YL94]. Surviving [Kur92]. Swayed [Mor98a]. sweep [PM98]. Switching [Hin92, Div96]. symbolism [Obe99c]. synthesis [DR91, Fox93, Rec97, Smi92a]. Synthetic [KRF97]. system [Bel90f, CCS99, Fin90, GOI98, HR96, Joh98, Jur90c, MA99, Per91, Rie92c, Sky94, WL98, Get93]. system-wide [Per91]. Systems [Ano96-40, Ada96h, CP96, Com93c, Com99a, Cox93, Gro98, Hel99, HT95, JT93, KWMH96, LP98, Mar90, MV90, Pas95, Rec92, Rec97, Rui94, SW97, TLN98].

Taaau [Ano96-124]. Table [Ano96-106, Ano96-107, Ano96-108, Ano96-109, Ano96-110, Ano96-111, Ano96-112, Ano96-113, Ano96-114, Ano96-115, Ano96-116, Ano96-117, Ano97-84, Ano97-85, Ano97-86, Ano97-87, Ano97-88, Ano97-89, Ano97-90, Ano97-91, Ano97-92, Ano97-93, Ano97-94, Ano97-95, Ano98-90, Ano98-91, Ano98-92, Ano98-93, Ano98-94, Ano98-95, Ano98-96, Ano98-97, Ano98-98, Ano98-99, Ano98-100, Ano98-101, Ano99-82, Ano99-83, Ano99-84, Ano99-85, Ano99-86, Ano99-87, Ano99-88, Ano99-89, Ano99-90, Ano99-91, Ano99-92, Ano99-93, Slo96]. Taboo [Ano97-96]. tabs [HSW95]. tack [GE95]. Tackles [Fos96c]. take [HS98, Nat95]. Takes [Mar96, Ber97, Wal97]. Taking [Hos94b, Fit90d]. talking [Kap92a]. tap [Mau94, MT92]. TAPS [Jur91d]. Teaching [Ano96-70, Ano96-122, Ano96-123, Hil91]. Teamwork [Per90c]. tech [Bar92, Bat92, Col92, DF92, Kap98b, Ric92, Ros91c, Smi92b]. Technical [Ano97a, Ano98-62, Ano99-54, Bel90f, BC99, Bel96c, Nor99, YL90, YL94]. Technically [Ano96-118, Ano96-119, Ano99-94, Sel96c, Sel96a, Sel96b, Sel97e, Sel97a, Sel97b, Sel97c, Sel97d, Sel98a, Sel98b, Sel98c, Sel98d, Sel98e, Sel98f, Sel99a, Sel99b, Sel99c, Sel99d, Sel99e, Zor97]. Technological [Fos97f, Mor99a, Gep95c, JWA^{+}91]. technologie [Com93b]. technologies [GS98, HL99, Kap94, Kas94]. Technology [AE90, AE91, Ada94b, Ano91c, Bel90g, Bel91c, Bel91b, Bel92b, Bel93c, Bel93e, Bel93d, Bel96b, BR97, BDF99, Bra99, Bro95, But96, Che91, Com92d, Com93e, Com93f, CG96, Com98a, Doo93a, Doo93b, Doo97a, Doo98, Fer98, Fin95, Fit90g, Fit91c, Fit91d, Fos98c, Gep98g, GS99, Gro93, Hel98, Hin98, Hul95, Jur91c, Jur91d, Jur92c, Jur92d, Jur92e, Jur92f, Jur93, Kap91, Kap92b, Kap93, Kap98a, Lan98, Loc95, Low92, Mas98, Mil98d, MW96, MV90, Mur90d, Per98b, Phi98, Rie92e, RS99a, Rie98e, Riv95, Ros91a, Ros91b, Ros92b, Ros98, SC90, Sar93, Sel93, Sha95, Spa97, Ste93, Sti90b, Swa93, Swe98e, Wal97, Wat91a, Wat91b, Wat92d, Zor90g, Zor90h, Zor91e, Zor91f, Zor92f, Zor92g, Zor93b, Zor93c, Zor94e, Ano90d, Ast95, Bel94b, Bel95d, Bel95e, Bel96d, Bel96e, Ber97, Bet99]. technology [Bre99, CKR94, Col93, Com97b, Com99b, Cre95, Cur93, DR99f, Fos99a, Gep99g, Gut90a, Hel97, IK99, Kap97b, Kap99b, Kor99, Kur99, Lan99, LK99b, Mal93, Mar99a, Mas99c, Mil98b, Per90c, Swe99f, Tad98, Trl97, Bel96a]. technology/electromagnetism [Mal93]. Tegopoulos [Ano96-91]. Telecom [Gom96, Apf97]. Telecommunications [Bel92b, Bel94d, DR96, Kob95, Pra94, Sti90c, Bel90b, Bel91b, Bel93d, BH92, Sai90]. telegraph [Fin90]. Telephone [Per90d, Tan98]. telephony [Joh98]. telescope [Bel95a, Bel91c]. Television [Mar97a, BBH97, Per95c]. Ten [Mur96]. terabit [Wil97]. term [Bel90c]. terminal [SCS91]. terminals [Com93g, Luc93]. terrestrial [Kas94]. Tesla [Ano96-91]. Test [Gro93, Mas98, Mas99c, Rie94, Rie95b, Spa97, Bel95b, Fit91d, Rie92e, Swe96a]. Testability [MT92]. Testing [CP96, Dav97, Fli94, HS96a, Rui94, Run99, Zor99, Bel91d, Com99h, Lau97, Lev92, Noa97, Rya98, SS91]. Tests [Rie96d, Apf97]. their [GS98, SB98a]. them [Luc93, Pel97]. themselves [Wol93]. theorem [Bri90]. Theory [Com95b]. thermal [McC92]. these [Sym98]. Thevenin [Bri90]. thick [Tho96]. Thin [All97, Tho96]. Thing [Ung96, DR97]. things [WL98]. thinking [GS97]. third [Bra96b, Han98b]. Thirty [Jon97, Jon98f, Jon98g, Jon98h, Jon98i, Jon98j, Jon98k, Jon98b, Jon98c, Jon98d, Jon98e, Jon98l, Jon99b, Jon99c, Jon99d, Jon99e, Jon99f, Jon99g, Jon99h, Jon99i, Jon99j, Jon99k, Jon99l, Jon99m, Swe97f]. Thomas [Ano96-42]. Thompson [Ano96o]. thoughts [Hei97]. threat [BCW97, JS99, Obe99b]. threats [Ada92b]. Three [Mar96, Che90c, Swe99d]. Threshold [BDF99]. Tighter [SW97]. tightrope [Cro91]. time [AAH98, Koe99, LY96, Luc97a, Mir97, Per90c, Sou98]. time-saver [Mir97]. timer [Cam97]. Times [Ste98a, Ros96]. timing [Cha97a, LHD^{+}90]. tips [Ano91b]. Tissue [Gal96]. Today [BDF99, Per94f, Bel92a]. togetherness [Gut90b]. Tomorrow [BDF99]. tone [Dix95]. Too [Cla98, DR97, CMS97, Fos97a, Whi93b]. tool [Rid98]. Tools [Faw94, HW94, Mir96b, Rie96a, Rie97f, Yof96, Ano90a, Ano90b, Ano96-43, Ano96-44, Ano96-45, Ano96-46, Ano96-47, Ano96-48, Ano96-49, Ano96-50, Ano96-51, Ano97-40, Ano97-41, Ano97-39, Ano98-37, Ano98-38, Ano98-39, Ano98-40, Ano98-41, Ano98-42, Ano98-43, Ano98-44, Ano98-45, Ano98-46, Ano98-47, Ano99-29, Ano99-38, Ano99-39, Ano99-40, Ano99-30, Ano99-31, Ano99-32, Ano99-33, Ano99-34, Ano99-35, Ano99-36, Ano99-37, Bra95c, Cen90, Eub98, Fos98c, Gep99c, Gil91, Kap92a, Mur90a, Rid94, Rie96b, Rie96c, Rie97a, Rie97b, Rie97c, Rie97d, Rie97e, Rie97g, Rie97h, Rie98a, Rob92, Sha99, Trl97]. top [Bar92, Cic95, Jur90e, PA91]. Total [Bel93f]. tour [Swa96]. towers [Rob95]. toy [Ano99-30, Ano99-31, Ano99-32, Ano99-33, Ano99-34, Ano99-35, Ano99-36, Ano99-37]. Toys [Rie96a, Rie97f, Ano96-43, Ano96-44, Ano96-45, Ano96-46, Ano96-47, Ano96-48, Ano96-49, Ano96-50, Ano96-51, Ano97-40, Ano97-41, Ano97-39, Ano98-37, Ano98-38, Ano98-39, Ano98-40, Ano98-41, Ano98-42, Ano98-43, Ano98-44, Ano98-45, Ano98-46, Ano98-47, Ano99-29, Ano99-38, Ano99-39, Ano99-40, Rie96b, Rie96c, Rie97a, Rie97b, Rie97c, Rie97d, Rie97e, Rie97g, Rie97h, Rie98a]. Traceable [Gem97]. Tracking [Jam98]. Tradeoffs [Gut90b]. trading [Wal97, Zor94d]. traffic [Ada91a, Per91, PA91, Per97c]. train [VG96]. training [KRF97, Yal97]. transcend [Ada90d]. transfer [Col92, Cur93]. transform [Nah91]. transformers [MAW97]. Transforming [MAW97]. transients [Mar90]. transistor [Ano99a, Joh94, Tau99]. transit [KHSK95]. Translating [Mir96c]. Transmission [Ano96-71, Hin93, Hin96h]. Transportation [Boe96, Bre99, Doo94d, Doo95d, Doo96e, Doo97c, Mur90e, Phi98, Che91, Doo93b, Jur92f, Mur90d]. travails [Per96]. travel [Luc96c]. Treasury [Mor97c]. Treat [Mar97a]. Trends [Gep96d, Com96b, Lar94]. Trends/developments [Gep96d, Com96b]. Trials [Gep97g, Per96]. tricks [Ito90]. Trimming [Cha97b]. Triumphs [Gep97g]. True [Gad98, Foc99]. Truly [Rid98, Foc99]. Trying [Fos97f]. tube [Bar98]. Tubes [Sym98, Bar98, CWG98]. tuning [AAH98, Gra99]. Tunnels [Ano97b]. Turbo [Ste99]. turf [Gut90b]. turning [KA94]. TV [Boo99, DR99b, Fox98, Jur92b, Per96, Sal92]. Twentieth [Fie97]. twin [Rob95]. Two [RLH93, Swe98f, Per91, Swe99d]. two-decade-old [Per91].

U.S. [MW96, Obe99c]. Uenohara [Ano96-69]. UK [Tab96]. Ultralight [Har99a]. Unbounded [Boe96]. Unclogging [Har92b]. Undergraduate [Ano96-122, Ano96-123]. unification [Nah92]. Union [Gre94]. unit [Pix96]. United [BE91, Boo99, SD94]. Universe [Zor95b]. University [Ano97b, Fie97, Swe98c, Che90d, Hil99]. Unlocking [HS96b]. unmanned [WK90]. Unnumbered [Swe98g]. Update [Kap99c]. updates [Fos96b]. upgrade [Per91]. upgraded [Fox99, Lev92]. Upgrading [Ada95b]. uphill [Gep95d]. usage [Rya98]. User [Ano90e, Bil96, Kap92c]. User-friendly [Ano90e, Bil96]. using [Gra99, RMAA91]. Utilities [Fil96, Zor91g, HBY90]. utility [SLHM94].

V [Ano96o, Ano96-124, PS92, Wil93]. vacuum [Bar98]. Value [Yea96, Eve98]. value-adding [Eve98]. vanishing [Kur99]. Vannevar [Mor98a, Zac95]. variable [Gra99]. vBNS [JNM^{+}98]. Veeravalli [Ano96o]. vehicle [Lyn93]. vehicles [Gil98, HS98, Rie92a, Rie92c, Rie92d, Rie92b, lH98]. vehicles-architecting [Rie92c]. vehicles-the [Rie92d]. vendors [Ros92a]. Venugopal [Ano96o]. Venus [But97]. verification [CK96]. Verifying [McM96]. versatility [Com92a]. verse [Kam99]. versed [Ell99c]. VHDL [Car93]. via [LC99]. Video [ARA91, Ham95, Mem98, Nat95, Dix95, JPN^{+}92, Jur92a]. videorecording [War96]. View [Jur90e, DSCN93, Ico97, Kap92c, Per94f, Sur90, Wol93]. Viewpoint [All96, Bli96, Boe96, Fuq96, Gal96, HS96a, HD96, Len96, Mos96, Par96, Van96, Wal96, Yea96, Yof96, Mor97c]. village [Pra94]. villages [Lys94]. Virtual [Ada93b, BC97b, DR98, Fos96d, Hin97c, San90, Gra98, Per97b, WB97]. visibility [BK99, Wer94]. Vision [Mar97b, Foc99, KM96]. visionary [Foc99]. visits [SB99b]. Visualization [Gal94, GE95, AL92, Bra95c, Fos97b, Fox99, Pen99, RT98]. visualized [Doe96]. Vital [Mil94b, Sym98]. Vladimir [Ano96-40]. voice [CMS97, Lan94]. Volta [MVJ96]. voltage [Hin96h]. voltages [BT93]. voltmeter [MVJ96]. Voyager [Bel90g]. vs [Gut90b]. vulnerable [Bel90c].

W [Ano96o, Ano96z, Ano96-73]. W. [Ano96-124]. Waiter [Ano96-124]. Walker [Rob99]. wallet [Per97b]. Wanlass [Rie91b]. war [Zor91d]. WARC [RTHM92]. warfare [Ada91b]. warming [Rob99]. wars [Com96k]. wartime [Ano90d]. Washington [Ada96b, Ada96c, Ada96d, Ada96e, Ada96f, Ada96g, Ada97a, Ada97b, Ada97c, Ada97d, Ada97e, Ada97f, Ada97g, Ada97h, Ada97i, Ada97j, Ada97k, Ada98a, Ada98b, Ada98c, Ada98d, Ada98e, Ada98f, Ada98g, Ada98h, Ada98i, Ada98j, Ada98k, Ada98l, Ada99a, Ada99b, Ada99c, Ada99d, Ada99e, Ada99f, Ada99g, Ada99h, Ada99i, Ada99j, Ada99k, Ada99l, Ano96-125, Ano96-126, Ano96-127]. Waste [ZS90, Bea97]. Watch [Ada96c, Ada96b, Ada96d, Ada96e, Ada96f, Ada96g, Ada97a, Ada97b, Ada97c, Ada97d, Ada97e, Ada97f, Ada97g, Ada97h, Ada97i, Ada97j, Ada97k, Ada98a, Ada98b, Ada98c, Ada98d, Ada98e, Ada98f, Ada98g, Ada98h, Ada98i, Ada98j, Ada98k, Ada98l, Ada99a, Ada99b, Ada99c, Ada99d, Ada99e, Ada99f, Ada99g, Ada99h, Ada99i, Ada99j, Ada99k, Ada99l, Ano96-53, Ano96-54, Ano96-55, Ano96-125, Ano96-126, Ano96-127, Ano98-48, Ano98-49, Ano99-41, Rie97i, Rie97j, Rie97k, Rie97l, Rie98c, Rie98d, Rie99a, Rie99b, Rie99c]. Water [Mak94, Kur92]. Watson [Ano96-123]. watts [Bha94, DdMMNM96]. wave [RLH93]. Wavelet [Aka97, BDG96]. waves [HT95, Lew95]. way [May94]. ways [Fos97f, Par97, RLH93]. Weaknesses [Zim97]. weapons [Fit91b]. weather [Doo95c, Rob97]. Web [Eng97a, Ano98-103, Ano96-128, Ano96-129, Ano96-130, Ano96-131, Ano96-132, Ano97-101, Ano99-96, Ano99-97, Ano99-98, Ano99-99, Ano99-100, Ano99-101, Com96e, Com96f, Com96a, Com96g, Com96h, Com96i, Com96j, Com97d, Com97e, Com97f, Com97g, Com97h, Com97k, Com97i, Com97j, Com98c, Com98d, Com98e, Com98f, Com98g, Com98h, Com98i, Com99d, Com99e, Com99f, Eng97b, Gep98c, JB96, Laz99, Mil98e, Mil98f, Mil98g, Mil99b, Ove97]. Websights [Com98j, Com99g]. Webware [Eng97c]. wedge [Fos98d]. Wegener [Ano97b]. well [Ell99c]. well-versed [Ell99c]. went [Hin99, Obe99d]. West [Per90e]. whatever [Hos94a]. Where [Box95]. Which [Jul94]. Whistle [Fit90h]. Whistle-blowing [Fit90h]. Who [Mor98a, Ros91c]. Wide [Eng97a, Gep98c, Per91]. Wiley [Wis99]. Will [Ano96-124, GS98, Luc96c]. Willinger [Ano96-124]. Wilson [Ano96-124, Ada95c]. Wind [Ano97b, DR99e, Jay95]. windings [MAW97]. Windows [Wal96, Str96, Zor95b]. winners [Han96, Han97, Han98b, Ste99]. Winning [MF97, Ste99, Yof96]. Wintel [MF97]. wire [Fin90, Mal93]. Wired [Ano96-133, GF98]. wireless [Col94, DR99c, Kob93, Wic96b, Hil99]. wiring [DR99e, Rad96]. wise [Mar98b, Sto97]. without [Doo96c, Ryd91]. Woe [Mar98e]. Women [Gep97g, BC99, Gep95d, Gep96a, Gep99a]. wonder [Nor95]. Wonderland [Lub97]. word [MK90]. work [Bar92, BBH97, DFC94, Ham93a, Jur90d, Mac91, Per92a, Tur91]. Workbench [San90]. workhorse [Hin96h]. Workhorses [Kha96]. Working [Ham93b, Mil98c, Ada92e]. workplace [Bar92, Bat92, DF92, Ric92, Smi92b]. works [Apf97, DP91]. workshop [Gum99]. Workstation [Com96k, Mog93, Ano91b, Ano91a, Jul94, KS92]. workstations [Ano91b, Ano91a, Com92a, Com93e, Com94a, Com94d, FT93, HPV92, HBY90, HFYT91, Kap92c, KS92, RMAA91, Ros91b, Ros92a, Ros92b, Voe90a]. World [Eng97a, Mor98a, Ste96, Bry96, Gut91, Hin99, KR90, LD96, Luc98f, Per93b, Rya98, Shi91, Spi96, Gep98c]. world-class [Shi91]. Worlds [BC97b]. worry [Zak99]. worst [Pow96]. wouldn’t [Bar90]. written [BDF99]. wrong [SLLP97]. wrought [Ros92a].

X [Com93g, Dal90, FK91, SCS91, Zor92c]. X-ray [Dal90, Zor92c]. X-rays [FK91]. x86 [Eub98].

Y2K [Per99a, Com99h, Tro99].

Yakubovich [Ano96-40]. Yale [Ano97b]. Year [Rob99, CP98, LP98]. years [Jai93, Jon97, Jon98f, Jon98g, Jon98h, Jon98i, Jon98j, Jon98k, Jon98b, Jon98c, Jon98d, Jon98e, Jon98l, Jon99b, Jon99c, Jon99d, Jon99e, Jon99f, Jon99g, Jon99h, Jon99i, Jon99j, Jon99k, Jon99l, Jon99m, Swe97f, Sym98]. Yi [Bel94a]. Yield [HS96a]. York [Mor98a, Rob99, Wis99]. young [Fow93]. you’re [Luc96d]. Yourself [Gad98].

Zachary [Mor98a]. Zadeh [Per95d].

References

Allan:1998:FTT

[AAH98] D. W. Allan, N. Ashby, and C. Hodge. Fine-tuning time in the Space Age. IEEE Spectrum, 35(3):42–51, March 1998. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anderson:1997:SD

[AC97] D. B. Anderson and M. A. Casey. The sound dimension. IEEE Spectrum, 34 (3):46–50, March 1997. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Adam:1990:CGE

[Ada90a] J. A. Adam. Competing in a global economy. IEEE Spectrum, 27(4):20–24, April 1990.

CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Adam:1990:DTS

[Ada90b] J. A. Adam. Defense: toward smaller, more deployable forces, as lethal as can be. IEEE Spectrum, 27(11): 30–33, November 1990. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Adam:1990:FLM

[Ada90c] J. A. Adam. Federal laboratories meet the marketplace. IEEE Spectrum, 27(10): 39–44, October 1990. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Adam:1990:ITN

[Ada90d] J. A. Adam. Industries transcend national boundaries. IEEE Spectrum, 27(9):26– 31, September 1990. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Adam:1991:ATC

[Ada91a] J. A. Adam. Air traffic control: satellites and artificial intelligence promise improved safety and efficiency. IEEE Spectrum, 28(2):27–32, February 1991. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Adam:1991:GLW

[Ada91b] J. A. Adam. Gulf legacy: warfare in the information age. IEEE Spectrum, 28(9): 26–33, September 1991. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Adam:1992:DSC

[Ada92a] J. A. Adam. Data security — cryptography = privacy? IEEE Spectrum, 29 (8):29–35, August 1992. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Adam:1992:DST

[Ada92b] J. A. Adam. Data security — threats and countermeasures. IEEE Spectrum, 29 (8):21–28, August 1992. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Adam:1992:EMP

[Ada92c] J. A. Adam. Electronic mail — playing on the net. IEEE Spectrum, 29(10):29, October 1992. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Adam:1992:HGC

[Ada92d] J. A. Adam. How government can help (defense production downsizing). IEEE Spectrum, 29(12):46–52, December 1992. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Adam:1992:SNS

[Ada92e] J. A. Adam. Seeking nuclear safeguards. II. Working to halt proliferation. IEEE Spectrum, 29(4):66–71, April 1992. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Adam:1993:MAI

[Ada93a] J. A. Adam. Multimediaapplications, implications. IEEE Spectrum, 30(3):24– 31, March 1993. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Adam:1993:VRR

[Ada93b] J. A. Adam. Virtual reality is for real. IEEE Spectrum, 30(10):22–29, October 1993. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Adam:1994:ME

[Ada94a] J. A. Adam. Medical electronics. IEEE Spectrum, 31 (1):70–73, January 1994. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Adam:1994:TCD

[Ada94b] J. A. Adam. Technology combats disabilities. IEEE Spectrum, 31(10):24–26, October 1994. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Adam:1995:ME

[Ada95a] J. A. Adam. Medical electronics. IEEE Spectrum, 32 (1):80–83, January 1995. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Adam:1995:UI

[Ada95b] J. A. Adam. Upgrading the Internet. IEEE Spectrum, 32(9): 24–29, September 1995. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Adam:1995:WG

[Ada95c] J. A. Adam. Wilson Greatbatch. IEEE Spectrum, 32 (3):56–61, March 1995. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Adam:1996:ANN

[Ada96a] J. Adam. Architects of the net of nets. IEEE Spectrum, 33(9): 56–63, September 1996. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Adam:1996:WWa

[Ada96b] J. Adam. Washington watch. IEEE Spectrum, 33(3):18T2– 18T4, March 1996. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Adam:1996:WWb

[Ada96c] J. Adam. Washington watch.

IEEE Spectrum, 33(7):18T1, July 1996. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Adam:1996:WWc

[Ada96d]  J. Adam.  Washington

watch. IEEE Spectrum, 33 (8):20T1, August 1996. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Adam:1996:WWd

[Ada96e] J. Adam. Washington watch.

IEEE Spectrum, 33(10):16T1– 16T2, October 1996. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Adam:1996:WWe

[Ada96f] J. Adam. Washington watch.

IEEE Spectrum, 33(11):22T2– 22T3, November 1996. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Adam:1996:WWf

[Ada96g]  J. Adam.  Washington

watch. IEEE Spectrum, 33 (12):22T2, December 1996. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Adam:1996:BKN

[Ada96h] J. A. Adam. Bart Kosko [neural networks/fuzzy systems engineer biography]. IEEE Spectrum, 33(2):58–62, February 1996. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Adam:1996:ME

[Ada96i] J. A. Adam. Medical electronics. IEEE Spectrum, 33 (1):92–95, January 1996. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Adam:1997:WWa

[Ada97a] J. Adam. Washington watch. IEEE Spectrum, 34(1):20T4– 20T6, January 1997. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Adam:1997:WWb

[Ada97b] J. Adam. Washington watch. IEEE Spectrum, 34(3):16T4– 16T6, March 1997. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Adam:1997:WWc

[Ada97c] J. Adam. Washington watch. IEEE Spectrum, 34(4):16T4– 16T6, April 1997. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Adam:1997:WWd

[Ada97d] J. Adam. Washington watch. IEEE Spectrum, 34(5):16T4, May 1997. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Adam:1997:WWe

[Ada97e] J. Adam. Washington watch.

IEEE Spectrum, 34(6):16T1– 16T4, June 1997. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Adam:1997:WWf

[Ada97f] J. Adam. Washington watch.

IEEE Spectrum, 34(7):16T6– 16T8, July 1997. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Adam:1997:WWg

[Ada97g] J. Adam. Washington watch.

IEEE Spectrum, 34(8):16T2– 16T4, August 1997. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Adam:1997:WWh

[Ada97h] J. Adam. Washington watch.

IEEE Spectrum, 34(9):24T4– 24T8, September 1997. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Adam:1997:WWi

[Ada97i] J. Adam. Washington watch.

IEEE Spectrum, 34(10):16T4– 16T8, October 1997. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Adam:1997:WWj

[Ada97j] J. Adam. Washington watch. IEEE Spectrum, 34(11):20T4– 20T5, November 1997. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Adam:1997:WWk

[Ada97k]  J. Adam.  Washington

watch. IEEE Spectrum, 34 (12):16T2, December 1997. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Adam:1998:WWa

[Ada98a] J. Adam. Washington watch. IEEE Spectrum, 35(1):16T2– 16T4, January 1998. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Adam:1998:WWb

[Ada98b] J. Adam. Washington watch. IEEE Spectrum, 35(2):16T2– 16T4, February 1998. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Adam:1998:WWc

[Ada98c]  J. Adam.  Washington

watch. IEEE Spectrum, 35 (3):24T2, March 1998. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Adam:1998:WWd

[Ada98d]  J. Adam.  Washington

watch. IEEE Spectrum, 35 (4):16T2, April 1998. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Adam:1998:WWe

[Ada98e] J. Adam. Washington watch.

IEEE Spectrum, 35(5):20T2– 20T4, May 1998. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Adam:1998:WWf

[Ada98f] J. Adam. Washington watch.

IEEE Spectrum, 35(6):20T2– 20T4, June 1998. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Adam:1998:WWg

[Ada98g] J. Adam. Washington watch.

IEEE Spectrum, 35(7):16T2– 16T4, July 1998. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Adam:1998:WWh

[Ada98h]  J. Adam.  Washington

watch. IEEE Spectrum, 35 (8):12T2, August 1998. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Adam:1998:WWi

[Ada98i] J. Adam. Washington watch. IEEE Spectrum, 35(9):16T2– 16T4, September 1998. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Adam:1998:WWj

[Ada98j] J. Adam. Washington watch. IEEE Spectrum, 35(10):16T2– 16T4, October 1998. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Adam:1998:WWk

[Ada98k] J. Adam. Washington watch. IEEE Spectrum, 35(11):16T2– 16T4, November 1998. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Adam:1998:WWl

[Ada98l]  J. Adam.  Washington

watch. IEEE Spectrum, 35 (12):16T2, December 1998. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Adam:1999:WWa

[Ada99a] J. Adam. Washington watch. IEEE Spectrum, 36(1):24T2– 24T4, January 1999. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Adam:1999:WWb

[Ada99b] J. Adam. Washington watch.

IEEE Spectrum, 36(2):16T2– 16T3, February 1999. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Adam:1999:WWc

[Ada99c] J. Adam. Washington watch.

IEEE Spectrum, 36(3):16T2– 16T3, March 1999. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Adam:1999:WWd

[Ada99d] J. Adam. Washington watch.

IEEE Spectrum, 36(4):26I2– 26I3, April 1999. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Adam:1999:WWe

[Ada99e] J. Adam. Washington watch.

IEEE Spectrum, 36(5):18I6– 18I8, May 1999. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Adam:1999:WWf

[Ada99f] J. Adam. Washington watch.

IEEE Spectrum, 36(6):26I1– 26I4, June 1999. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Adam:1999:WWg

[Ada99g] J. Adam. Washington watch. IEEE Spectrum, 36(7):16I4– 16I5, July 1999. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Adam:1999:WWh

[Ada99h] J. Adam. Washington watch. IEEE Spectrum, 36(8):16I6– 16I8, August 1999. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Adam:1999:WWi

[Ada99i] J. Adam. Washington watch. IEEE Spectrum, 36(9):26I2– 26I4, September 1999. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Adam:1999:WWj

[Ada99j] J. Adam. Washington watch. IEEE Spectrum, 36(10):26I2– 26I3, October 1999. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Adam:1999:WWk

[Ada99k] J. Adam. Washington watch. IEEE Spectrum, 36(11):26I2– 26I4, November 1999. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Adam:1999:WWl

[Ada99l]  J. Adam.  Washington

watch. IEEE Spectrum, 36 (12):24I2, December 1999. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Adam:1990:TAM

[AE90] J. A. Adam and K. Esch.

Technology ’90; aerospace and military. IEEE Spectrum, 27 (1):49–51, January 1990. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Adam:1991:TAM

[AE91] J. A. Adam and K. Esch.

Technology 1991: aerospace/ military. IEEE Spectrum, 28 (1):72–75, January 1991. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Asthana:1995:SOS

[AF95] P. Asthana and B. Finkelstein. Superdense optical storage. IEEE Spectrum, 32(8): 25–31, August 1995. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Adam:1995:PPD

[AHP^{+}95] J. A. Adam, L. J. Hoffman, S. Perrin, M. Rotenberg, M. Stansellgamm, R. N. Veeder, and W. H. Ware. The privacy problem — discussion. IEEE Spectrum, 32(12):

46–52, December 1995. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Akay:1997:WAM

[Aka97] M. Akay. Wavelet applications in medicine. IEEE Spectrum, 34(5):50–56, May 1997. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Arrott:1992:SPV

[AL92] M. Arrott and S. Latta. Supercomputers — perspectives on visualization. IEEE Spectrum, 29(9):61–65, September 1992. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Allenby:1996:VYR

[All96] B. R. Allenby. Viewpoint: Are you ready for the science of industrial ecology? IEEE Spectrum, 33(1):81, January 1996. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Alles:1997:TFS

[All97] M. L. Alles. Thin film SOI emerges. IEEE Spectrum, 34 (6):37–45, June 1997. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Altnether:1996:CMB

[Alt96] J. Altnether. Circuits of the mind [books]. IEEE Spectrum, 33(3):9–11, March 1996. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anderson:1997:MIS

[And97] H. Anderson. Money and the Internet: a strange new relationship. IEEE Spectrum, 34 (2):74–76, February 1997. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1990:BTP

[Ano90a] Anonymous. Board tools proliferate software. IEEE Spectrum, 27(11):82–85, November 1990. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1990:CCT

[Ano90b] Anonymous. The case for CASE tools. IEEE Spectrum, 27(11):78–81, November 1990. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1990:DPD

[Ano90c] Anonymous. Displaying processed data. IEEE Spectrum, 27(11):86–89, November 1990. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1990:KPR

[Ano90d]  Anonymous.  Keithley’s

phantom repeater: applying wartime technology. IEEE Spectrum, 27(12):71, December 1990. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1990:UFC

[Ano90e]  Anonymous.  User-friendly

CAE. IEEE Spectrum, 27(11): 90–92, November 1990. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1991:GEWb

[Ano91a] Anonymous. A guide to engineering workstations: add-ons for your workstation. IEEE Spectrum, 28(4):52–57, April 1991. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1991:GEWa

[Ano91b] Anonymous. A guide to engineering workstations: tips on setting up your workstation. IEEE Spectrum, 28 (4):36–37, April 1991. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1991:TS

[Ano91c] Anonymous. Technology 1991: software. IEEE Spectrum, 28 (1):36–39, January 1991. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1996:MIA

[Ano96a] Anonymous. The 1996 Medalists [IEEE awards]. IEEE Spectrum, 33(6):56–60, June 1996. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1996:AIC

[Ano96b] Anonymous. Award in International Communication — Kunishi Nosaka. IEEE Spectrum, 33(12):59, December 1996. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1996:ISFa

[Ano96c] Anonymous. IEEE Spectrum — front cover. IEEE Spectrum, 33(1):c1, January 1996. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1996:ISFb

[Ano96d] Anonymous. IEEE Spectrum — front cover. IEEE Spectrum, 33(2):c1, February 1996. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1996:ISFc

[Ano96e] Anonymous. IEEE Spectrum — front cover. IEEE Spectrum, 33(3):c1, March 1996. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1996:ISFg

[Ano96f] Anonymous. IEEE Spectrum — front cover. IEEE Spectrum, 33(7):c1, July 1996. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1996:ISFi

[Ano96g]  Anonymous.  IEEE Spec-

trum — front cover. IEEE Spectrum, 33(9):c1, September 1996. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1996:ISFj

[Ano96h]  Anonymous.  IEEE Spec-

trum — front cover. IEEE Spectrum, 33(10):c1, October 1996. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1996:ISFk

[Ano96i]  Anonymous.  IEEE Spec-

trum — front cover. IEEE Spectrum, 33(11):c1, November 1996. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1996:ISFl

[Ano96j]  Anonymous.  IEEE Spec-

trum — front cover. IEEE Spectrum, 33(12):c1, December 1996. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1996:ISFd

[Ano96k] Anonymous. IEEE Spectrum [front cover]. IEEE Spectrum, 33(4):c1, April 1996. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1996:ISFe

[Ano96l] Anonymous. IEEE Spectrum [front cover]. IEEE Spectrum, 33(5):c1, May 1996. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1996:ISFf

[Ano96m] Anonymous. IEEE Spectrum [front cover]. IEEE Spectrum, 33(6):c1, June 1996. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1996:ISFh

[Ano96n] Anonymous. IEEE Spectrum [front cover]. IEEE Spectrum, 33(8):c1, August 1996. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1996:BJT

[Ano96o]  Anonymous.  Browder J.

Thompson Prize Award — Carl W. Baum and Venugopal V. Veeravalli. IEEE Spectrum, 33(12):64, December 1996. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1996:Ca

[Ano96p]  Anonymous.  Calendar.

IEEE Spectrum, 33(1):14– 118E4, January 1996. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1996:Cb

[Ano96q]  Anonymous.  Calendar.

IEEE Spectrum, 33(2):14– 64E5, February 1996. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1996:Cc

[Ano96r]  Anonymous.  Calendar.

IEEE Spectrum, 33(3):12– 68E4, March 1996. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1996:Cd

[Ano96s]  Anonymous.  Calendar.

IEEE Spectrum, 33(4):15– 18E4, April 1996. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1996:Ce

[Ano96t]  Anonymous.  Calendar.

IEEE Spectrum, 33(7):16– 94E7, July 1996. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1996:Cf

[Ano96u]  Anonymous.  Calendar.

IEEE Spectrum, 33(8):16– 60E10, August 1996. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1996:Cg

[Ano96v]  Anonymous.  Calendar.

IEEE Spectrum, 33(9):16– 80E2, September 1996. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1996:Ch

[Ano96w]  Anonymous.  Calendar.

IEEE Spectrum, 33(10):16– 64E8, October 1996. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1996:Ci

[Ano96x] Anonymous. Calendar. IEEE Spectrum, 33(11):20–80E12, November 1996. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1996:Cj

[Ano96y]  Anonymous.  Calendar.

IEEE Spectrum, 33(12):14– 66E8, December 1996. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1996:CPS

[Ano96z] Anonymous. Charles Proteus Steinmetz Award — Marco W. Migliaro. IEEE Spectrum, 33(12):61–62, December 1996. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1996:CBA

[Ano96-27] Anonymous. Cledo Brunetti Award — Mitsumasa Koyanagi. IEEE Spectrum, 33 (12):58, December 1996. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1996:CSg

[Ano96-28] Anonymous. Coming in Spectrum. IEEE Spectrum, 33 (7):112, July 1996. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1996:CSa

[Ano96-29] Anonymous. Coming in spectrum. IEEE Spectrum, 33 (1):136, January 1996. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1996:CSb

[Ano96-30] Anonymous. Coming in spectrum. IEEE Spectrum, 33 (2):80, February 1996. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1996:CSc

[Ano96-31] Anonymous. Coming in spectrum. IEEE Spectrum, 33 (3):84, March 1996. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1996:CSd

[Ano96-32] Anonymous. Coming in spectrum. IEEE Spectrum, 33 (4):100, April 1996. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1996:CSe

[Ano96-33] Anonymous. Coming in spectrum. IEEE Spectrum, 33 (5):96, May 1996. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1996:CSf

[Ano96-34] Anonymous. Coming in spectrum. IEEE Spectrum, 33 (6):100, June 1996. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1996:CSh

[Ano96-35] Anonymous. Coming in spectrum. IEEE Spectrum, 33 (8):80, August 1996. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1996:CSi

[Ano96-36] Anonymous. Coming in spectrum. IEEE Spectrum, 33(9): 100, September 1996. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1996:CSj

[Ano96-37] Anonymous. Coming in spectrum. IEEE Spectrum, 33 (10):80, October 1996. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1996:CSk

[Ano96-38] Anonymous. Coming in spectrum. IEEE Spectrum, 33(11): 100, November 1996. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1996:CSl

[Ano96-39] Anonymous. Coming in spectrum. IEEE Spectrum, 33 (12):88, December 1996. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1996:CSAa

[Ano96-40] Anonymous. Control Systems Award — Vladimir A. Yakubovich. IEEE Spectrum, 33(12):58, December 1996. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1996:DSA

[Ano96-41] Anonymous. David Sarnoff Award — Hiroyuki Sakaki. IEEE Spectrum, 33(12):61, December 1996. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1996:DGF

[Ano96-42] Anonymous. Donald G. Fink Prize Award — Ali H. Sayed and Thomas Kailath. IEEE Spectrum, 33(12):64, December 1996. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1996:ETTa

[Ano96-43] Anonymous. EEs’ tools toys.

IEEE Spectrum, 33(2):20– 22, February 1996. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1996:ETTb

[Ano96-44]  Anonymous.  EEs’ tools

toys. IEEE Spectrum, 33 (4):83–85, April 1996. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1996:ETTc

[Ano96-45] Anonymous. EEs’ tools toys.

IEEE Spectrum, 33(5):77–80, May 1996. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1996:ETTd

[Ano96-46] Anonymous. EEs’ tools toys.

IEEE Spectrum, 33(6):85–88, June 1996. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1996:ETTe

[Ano96-47] Anonymous. EEs’ tools toys.

IEEE Spectrum, 33(7):95–97, July 1996. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1996:ETTf

[Ano96-48]  Anonymous.  EEs’ tools

toys. IEEE Spectrum, 33(8): 61–62, August 1996. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1996:ETTg

[Ano96-49] Anonymous. EEs’ tools toys. IEEE Spectrum, 33(9):84– 85, September 1996. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1996:ETTh

[Ano96-50] Anonymous. Ees’ tools toys. IEEE Spectrum, 33(10):59– 62, October 1996. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1996:ETTi

[Ano96-51] Anonymous. EEs’ tools toys. IEEE Spectrum, 33(12):65– 66, December 1996. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1996:ERP

[Ano96-52] Anonymous. Emanuel R. Piore Award — Edward J. Mc-Cluskey. IEEE Spectrum, 33 (12):61, December 1996. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1996:EWa

[Ano96-53] Anonymous. EV watch. IEEE Spectrum, 33(3):16, March 1996. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1996:EWb

[Ano96-54]  Anonymous.  EV watch.

IEEE Spectrum, 33(6):84, June 1996. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1996:EWc

[Ano96-55] Anonymous. EV watch. IEEE Spectrum, 33(12):20, December 1996. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1996:FF

[Ano96-56]  Anonymous.  Faults fail-

ures. IEEE Spectrum, 33 (2):17, February 1996. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1996:Fa

[Ano96-57] Anonymous. Forum. IEEE Spectrum, 33(1):6–8, January 1996. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1996:Fb

[Ano96-58] Anonymous. Forum. IEEE Spectrum, 33(2):6–8, February 1996. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1996:Fc

[Ano96-59] Anonymous. Forum. IEEE Spectrum, 33(3):6–8, March 1996. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1996:Fd

[Ano96-60] Anonymous. Forum. IEEE Spectrum, 33(4):6–8, April 1996. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1996:Fe

[Ano96-61] Anonymous. Forum. IEEE Spectrum, 33(5):6–8, May 1996. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1996:Fg

[Ano96-62] Anonymous. Forum. IEEE Spectrum, 33(6):6, June 1996. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1996:Fh

[Ano96-63] Anonymous. Forum. IEEE Spectrum, 33(7):7–8, July 1996. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1996:Fi

[Ano96-64] Anonymous. Forum. IEEE Spectrum, 33(8):6–8, August 1996. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1996:Fj

[Ano96-65] Anonymous. Forum. IEEE Spectrum, 33(9):6–8, September 1996. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1996:Fk

[Ano96-66] Anonymous. Forum. IEEE Spectrum, 33(10):6–8, October 1996. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1996:Fl

[Ano96-67] Anonymous. Forum. IEEE Spectrum, 33(11):6–12, November 1996. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1996:Fm

[Ano96-68] Anonymous. Forum. IEEE Spectrum, 33(12):6–8, December 1996. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1996:FPA

[Ano96-69] Anonymous. Frederik Philips Award — Michivuki Uenohara. IEEE Spectrum, 33(12): 60–61, December 1996. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1996:GTA

[Ano96-70] Anonymous. Graduate Teaching Award — Sung-Mo (Steve) Kang. IEEE Spectrum, 33(12):62, December 1996. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1996:HHE

[Ano96-71] Anonymous. Herman Halperin Electric Transmission and Distribution Award — Farouk A. M. Rizk. IEEE Spectrum, 33 (12):58, December 1996. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1996:IGA

[Ano96-72] Anonymous. IEEE guide for authors. IEEE Spectrum, 33(1):107–113, January 1996. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1996:JMA

[Ano96-73] Anonymous. Jack A. Morton Award — Robert W. Dutton. IEEE Spectrum, 33(12): 60, December 1996. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1996:JRA

[Ano96-74]  Anonymous.  Judith A.

Resnik Award — Paul G. Steffes. IEEE Spectrum, 33 (12):61, December 1996. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1996:KKC

[Ano96-75] Anonymous. Koji Kobayashi Computers and Communications Award — K. Mani Chandy. IEEE Spectrum, 33 (12):59–60, December 1996. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1996:MIC

[Ano96-76] Anonymous. Masaru Ibuka Consumer Electronics Award — Kees A. Schouhamer Immink. IEEE Spectrum, 33 (12):59, December 1996. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1996:MLA

[Ano96-77]  Anonymous.  Morris E.

Leeds Award — Mark B. Ketchen. IEEE Spectrum, 33 (12):60, December 1996. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1996:MLM

[Ano96-78] Anonymous. Morris N. Liebman Memorial Award — Seiki Ogura. IEEE Spectrum, 33 (12):60, December 1996. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1996:Na

[Ano96-79] Anonymous. Newslog. IEEE Spectrum, 33(1):1, January 1996. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1996:Nb

[Ano96-80] Anonymous. Newslog. IEEE Spectrum, 33(2):1, February 1996. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1996:Nc

[Ano96-81] Anonymous. Newslog. IEEE Spectrum, 33(3):1, March 1996. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1996:Nd

[Ano96-82] Anonymous. Newslog. IEEE Spectrum, 33(4):1, April 1996. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1996:Ne

[Ano96-83] Anonymous. Newslog. IEEE Spectrum, 33(5):1, May 1996. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1996:Nf

[Ano96-84] Anonymous. Newslog. IEEE Spectrum, 33(6):1, June 1996. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1996:Ng

[Ano96-85] Anonymous. Newslog. IEEE Spectrum, 33(7):1, July 1996. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1996:Nh

[Ano96-86] Anonymous. Newslog. IEEE Spectrum, 33(8):1, August 1996. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1996:Ni

[Ano96-87] Anonymous. Newslog. IEEE Spectrum, 33(9):1, September 1996. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1996:Nj

[Ano96-88] Anonymous. Newslog. IEEE Spectrum, 33(10):1, October 1996. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1996:Nk

[Ano96-89] Anonymous. Newslog. IEEE Spectrum, 33(11):1, November 1996. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1996:Nl

[Ano96-90] Anonymous. Newslog. IEEE Spectrum, 33(12):1, December 1996. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1996:NTA

[Ano96-91] Anonymous. Nikola Tesla Award — John A. Tegopoulos. IEEE Spectrum, 33(12):

62, December 1996. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1996:PEa

[Ano96-92] Anonymous. The practical engineer. IEEE Spectrum, 33(7):67, July 1996. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1996:PEb

[Ano96-93] Anonymous. The practical engineer. IEEE Spectrum, 33 (9):65, September 1996. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1996:PEc

[Ano96-94] Anonymous. The practical engineer. IEEE Spectrum, 33 (11):61, November 1996. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1996:RBa

[Ano96-95] Anonymous. Recent books. IEEE Spectrum, 33(3):18T6– 66T8, March 1996. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1996:RBb

[Ano96-96] Anonymous. Recent books. IEEE Spectrum, 33(5):80T2, May 1996. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1996:RBc

[Ano96-97] Anonymous. Recent books.

IEEE Spectrum, 33(7):18T2– 94T4, July 1996. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1996:RBd

[Ano96-98] Anonymous. Recent books.

IEEE Spectrum, 33(8):14– 60T3, August 1996. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1996:RBe

[Ano96-99] Anonymous. Recent books.

IEEE Spectrum, 33(9):13– 22T6, September 1996. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1996:RBf

[Ano96-100] Anonymous. Recent books.

IEEE Spectrum, 33(10):14– 64T8, October 1996. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1996:RBg

[Ano96-101] Anonymous. Recent books.

IEEE Spectrum, 33(11):22T4– 78T4, November 1996. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1996:RBh

[Ano96-102] Anonymous. Recent books. IEEE Spectrum, 33(12):13– 66T2, December 1996. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1996:RBJ

[Ano96-103] Anonymous. Reynold B. Johnson Information Storage Award — Nobutake Imamura. IEEE Spectrum, 33 (12):59, December 1996. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1996:RHK

[Ano96-104] Anonymous. Richard Harold Kaufmann Award — Marcus O. Durham. IEEE Spectrum, 33(12):59, December 1996. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1996:SSC

[Ano96-105] Anonymous. Solid-State Circuits Award — Rudy J. van de Plassche. IEEE Spectrum, 33 (12):61, December 1996. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1996:TCa

[Ano96-106] Anonymous. Table of contents. IEEE Spectrum, 33 (1):2–4, January 1996. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1996:TCb

[Ano96-107] Anonymous. Table of contents. IEEE Spectrum, 33 (2):2–4, February 1996. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1996:TCc

[Ano96-108] Anonymous. Table of contents. IEEE Spectrum, 33 (3):2–4, March 1996. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1996:TCd

[Ano96-109] Anonymous. Table of contents. IEEE Spectrum, 33 (4):2–4, April 1996. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1996:TCe

[Ano96-110] Anonymous. Table of contents. IEEE Spectrum, 33 (5):2–4, May 1996. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1996:TCf

[Ano96-111] Anonymous. Table of contents. IEEE Spectrum, 33 (6):2–4, June 1996. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1996:TCg

[Ano96-112] Anonymous. Table of contents. IEEE Spectrum, 33 (7):2–4, July 1996. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1996:TCh

[Ano96-113] Anonymous. Table of contents. IEEE Spectrum, 33 (8):2–4, August 1996. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1996:TCi

[Ano96-114] Anonymous. Table of contents. IEEE Spectrum, 33(9): 2–4, September 1996. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1996:TCj

[Ano96-115] Anonymous. Table of contents. IEEE Spectrum, 33 (10):2–4, October 1996. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1996:TCk

[Ano96-116] Anonymous. Table of contents. IEEE Spectrum, 33(11): 2–4, November 1996. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1996:TCl

[Ano96-117] Anonymous. Table of contents. IEEE Spectrum, 33(12): 2–4, December 1996. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1996:TSa

[Ano96-118] Anonymous. Technically speaking. IEEE Spectrum, 33 (2):18, February 1996. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1996:TSb

[Ano96-119] Anonymous. Technically speaking. IEEE Spectrum, 33(4):19, April 1996. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1996:PFa

[Ano96-120] Anonymous. To probe further. IEEE Spectrum, 33(1): 105–106, January 1996. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1996:PFb

[Ano96-121] Anonymous. To probe further. IEEE Spectrum, 33 (7):65, July 1996. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1996:UTAa

[Ano96-122] Anonymous. Undergraduate Teaching Awards — David A. Patterson. IEEE Spectrum, 33 (12):62, December 1996. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1996:UTAb

[Ano96-123] Anonymous. Undergraduate Teaching Awards — Karan L. Watson. IEEE Spectrum, 33 (12):62, December 1996. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1996:WRG

[Ano96-124] Anonymous. W. R. G. Baker Prize Award — Will E. Leland, Waiter Willinger, Daniel V. Wilson and Murad S. Taaau. IEEE Spectrum, 33(12):63–64, December 1996. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1996:WWa

[Ano96-125] Anonymous. Washington watch. IEEE Spectrum, 33 (2):16T1, February 1996. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1996:WWb

[Ano96-126] Anonymous. Washington watch. IEEE Spectrum, 33(4): 18T2–18T4, April 1996. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1996:WWc

[Ano96-127] Anonymous. Washington watch. IEEE Spectrum, 33(9): 22T2–22T4, September 1996. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1996:WSa

[Ano96-128] Anonymous. Web sights.

IEEE Spectrum, 33(2):65, February 1996. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1996:WSb

[Ano96-129] Anonymous. Web sights.

IEEE Spectrum, 33(4):80, April 1996. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1996:WSc

[Ano96-130] Anonymous. Web sights.

IEEE Spectrum, 33(5):76, May 1996. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1996:WSd

[Ano96-131] Anonymous. Web sights.

IEEE Spectrum, 33(6):96, June 1996. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1996:WSe

[Ano96-132] Anonymous. Web sights.

IEEE Spectrum, 33(9):95, September 1996. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1996:WI

[Ano96-133] Anonymous. Wired for interactivity. IEEE Spectrum, 33(4):21, April 1996. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1997:TFA

[Ano97a] Anonymous. 1997 Technical Field Awards. IEEE Spectrum, 34(8):70–76, August 1997. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1997:AUR

[Ano97b] Anonymous. Aerodynamics under the Reich: review of The Peenemünde Wind Tunnels: A Memoir, Wegener, Peter P., Yale University Press, New Haven, Conn., 1996, 187 pp, $30. IEEE Spectrum, 34 (4):10–12, April 1997. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1997:BMD

[Ano97c] Anonymous. Ballistic mission defense: It’s back. IEEE Spectrum, 34(9):26–27, September 1997. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1997:ISFa

[Ano97d] Anonymous. IEEE Spectrum — front cover. IEEE Spectrum, 34(1):c1, January 1997. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1997:ISFb

[Ano97e] Anonymous. IEEE Spectrum — front cover. IEEE Spectrum, 34(2):c1, February 1997. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1997:ISFc

[Ano97f] Anonymous. IEEE Spectrum — front cover. IEEE Spectrum, 34(3):c1, March 1997. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1997:ISFd

[Ano97g] Anonymous. IEEE Spectrum — front cover. IEEE Spectrum, 34(4):c1, April 1997. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1997:ISFe

[Ano97h] Anonymous. IEEE Spectrum — front cover. IEEE Spectrum, 34(5):c1, May 1997. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1997:ISFf

[Ano97i] Anonymous. IEEE Spectrum — front cover. IEEE Spectrum, 34(6):c1, June 1997. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1997:ISFg

[Ano97j] Anonymous. IEEE Spectrum — front cover. IEEE Spectrum, 34(7):c1, July 1997. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1997:ISFh

[Ano97k] Anonymous. IEEE Spectrum — front cover. IEEE Spectrum, 34(8):c1, August 1997. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1997:ISFi

[Ano97l]  Anonymous.  IEEE Spec-

trum — front cover. IEEE Spectrum, 34(9):c1, September 1997. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1997:ISFj

[Ano97m]  Anonymous.  IEEE Spec-

trum — front cover. IEEE Spectrum, 34(10):c1, October 1997. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1997:ISFk

[Ano97n]  Anonymous.  IEEE Spec-

trum — front cover. IEEE Spectrum, 34(11):c1, November 1997. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1997:ISFl

[Ano97o]  Anonymous.  IEEE Spec-

trum — front cover. IEEE Spectrum, 34(12):c1, December 1997. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1997:Ca

[Ano97p]  Anonymous.  Calendar.

IEEE Spectrum, 34(1):15– 104E10, January 1997. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1997:Cb

[Ano97q]  Anonymous.  Calendar.

IEEE Spectrum, 34(2):13– 80E11, February 1997. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1997:Cc

[Ano97r]  Anonymous.  Calendar.

IEEE Spectrum, 34(3):14– 76E16, March 1997. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1997:Cd

[Ano97s]  Anonymous.  Calendar.

IEEE Spectrum, 34(4):16– 56E2, April 1997. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1997:Ce

[Ano97t]  Anonymous.  Calendar.

IEEE Spectrum, 34(5):14– 72E4, May 1997. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1997:Cf

[Ano97u]  Anonymous.  Calendar.

IEEE Spectrum, 34(6):15– 72E8, June 1997. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1997:Cg

[Ano97v]  Anonymous.  Calendar.

IEEE Spectrum, 34(7):10– 16E4, July 1997. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1997:Ch

[Ano97w] Anonymous. Calendar. IEEE Spectrum, 34(8):14–80E8, August 1997. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1997:Ci

[Ano97x]  Anonymous.  Calendar.

IEEE Spectrum, 34(9):10– 96E6, September 1997. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1997:Cj

[Ano97y]  Anonymous.  Calendar.

IEEE Spectrum, 34(10):16– 72E4, October 1997. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1997:Ck

[Ano97z] Anonymous. Calendar. IEEE Spectrum, 34(11):19–20E12, November 1997. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1997:Cl

[Ano97-27] Anonymous. Calendar. IEEE Spectrum, 34(12):16–16E16, December 1997. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1997:CSa

[Ano97-28] Anonymous. Coming in spectrum. IEEE Spectrum, 34 (1):130, January 1997. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1997:CSb

[Ano97-29] Anonymous. Coming in spectrum. IEEE Spectrum, 34 (2):98, February 1997. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1997:CSc

[Ano97-30] Anonymous. Coming in spectrum. IEEE Spectrum, 34 (4):70, April 1997. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1997:CSd

[Ano97-31] Anonymous. Coming in spectrum. IEEE Spectrum, 34 (5):90, May 1997. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1997:CSe

[Ano97-32] Anonymous. Coming in spectrum. IEEE Spectrum, 34 (6):90, June 1997. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1997:CSf

[Ano97-33] Anonymous. Coming in spectrum. IEEE Spectrum, 34 (7):86, July 1997. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1997:CSg

[Ano97-34] Anonymous. Coming in spectrum. IEEE Spectrum, 34 (8):98, August 1997. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1997:CSh

[Ano97-35] Anonymous. Coming in spectrum. IEEE Spectrum, 34(9): 110, September 1997. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1997:CSi

[Ano97-36] Anonymous. Coming in spectrum. IEEE Spectrum, 34 (10):94, October 1997. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1997:CSj

[Ano97-37] Anonymous. Coming in spectrum. IEEE Spectrum, 34 (12):94, December 1997. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1997:DBD

[Ano97-38] Anonymous. Deep blue’s dad [review of Hal’s legacy: 2001 computer as dream and reality (Stork, D. G.; 1997). IEEE Spectrum, 34(7):12–14, July 1997. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1997:ETTc

[Ano97-39] Anonymous. EE’s tools and toys. IEEE Spectrum, 34(10): 70–72, October 1997. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1997:ETTa

[Ano97-40] Anonymous. EEs’ tools toys. IEEE Spectrum, 34(2):81– 84, February 1997. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1997:ETTb

[Ano97-41] Anonymous. EEs’ tools toys. IEEE Spectrum, 34(9):94– 96, September 1997. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1997:Fa

[Ano97-42] Anonymous. Forum. IEEE Spectrum, 34(1):6–8, January 1997. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1997:Fb

[Ano97-43] Anonymous. Forum. IEEE Spectrum, 34(2):6–8, February 1997. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1997:Fc

[Ano97-44] Anonymous. Forum. IEEE Spectrum, 34(3):6–8, March 1997. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1997:Fd

[Ano97-45] Anonymous. Forum. IEEE Spectrum, 34(4):6–8, April 1997. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1997:Fe

[Ano97-46] Anonymous. Forum. IEEE Spectrum, 34(5):6, May 1997. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1997:Ff

[Ano97-47] Anonymous. Forum. IEEE Spectrum, 34(6):6, June 1997. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1997:Fg

[Ano97-48] Anonymous. Forum. IEEE Spectrum, 34(7):6–8, July 1997. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1997:Fh

[Ano97-49] Anonymous. Forum. IEEE Spectrum, 34(8):6–8, August 1997. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1997:Fi

[Ano97-50] Anonymous. Forum. IEEE Spectrum, 34(9):6–8, September 1997. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1997:Fj

[Ano97-51] Anonymous. Forum. IEEE Spectrum, 34(10):6–8, October 1997. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1997:Fk

[Ano97-52] Anonymous. Forum. IEEE Spectrum, 34(11):6–10, November 1997. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1997:Fl

[Ano97-53] Anonymous. Forum. IEEE Spectrum, 34(12):6–8, December 1997. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1997:IAM

[Ano97-54] Anonymous. IEEE awards — the 1997 Medalists. IEEE Spectrum, 34(6):24–30, June 1997. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1997:Na

[Ano97-55] Anonymous. Newslog. IEEE Spectrum, 34(1):1, January 1997. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1997:Nb

[Ano97-56] Anonymous. Newslog. IEEE Spectrum, 34(2):1, February 1997. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1997:Nc

[Ano97-57] Anonymous. Newslog. IEEE Spectrum, 34(3):1, March 1997. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1997:Nd

[Ano97-58] Anonymous. Newslog. IEEE Spectrum, 34(4):1, April 1997. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1997:Ne

[Ano97-59] Anonymous. Newslog. IEEE Spectrum, 34(5):1, May 1997. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1997:Nf

[Ano97-60] Anonymous. Newslog. IEEE Spectrum, 34(6):1, June 1997. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1997:Ng

[Ano97-61] Anonymous. Newslog. IEEE Spectrum, 34(7):1, July 1997. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1997:Nh

[Ano97-62] Anonymous. Newslog. IEEE Spectrum, 34(8):1, August 1997. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1997:Ni

[Ano97-63] Anonymous. Newslog. IEEE Spectrum, 34(9):1, September 1997. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1997:Nj

[Ano97-64] Anonymous. Newslog. IEEE Spectrum, 34(10):1, October 1997. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1997:Nk

[Ano97-65] Anonymous. Newslog. IEEE Spectrum, 34(11):1, November 1997. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1997:Nl

[Ano97-66] Anonymous. Newslog. IEEE Spectrum, 34(12):1, December 1997. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1997:PEa

[Ano97-67] Anonymous. The practical engineer. IEEE Spectrum, 34 (3):59, March 1997. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1997:PEb

[Ano97-68] Anonymous. The practical engineer. IEEE Spectrum, 34(5):59, May 1997. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1997:PEc

[Ano97-69] Anonymous. The practical engineer. IEEE Spectrum, 34(7):57, July 1997. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1997:PEd

[Ano97-70] Anonymous. The practical engineer. IEEE Spectrum, 34 (9):79, September 1997. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1997:RBa

[Ano97-71] Anonymous. Recent books.

IEEE Spectrum, 34(1):20T8– 20T14, January 1997. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1997:RBb

[Ano97-72] Anonymous. Recent books.

IEEE Spectrum, 34(3):11– 76T10, March 1997. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1997:RBc

[Ano97-73] Anonymous. Recent books.

IEEE Spectrum, 34(4):16T8– 56T5, April 1997. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1997:RBd

[Ano97-74] Anonymous. Recent books.

IEEE Spectrum, 34(5):16T8– 72T4, May 1997. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1997:RBe

[Ano97-75] Anonymous. Recent books. IEEE Spectrum, 34(6):14– 72T4, June 1997. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1997:RBf

[Ano97-76] Anonymous. Recent books. IEEE Spectrum, 34(7):15– 16T10, July 1997. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1997:RBg

[Ano97-77] Anonymous. Recent books. IEEE Spectrum, 34(8):12– 80T4, August 1997. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1997:RBh

[Ano97-78] Anonymous. Recent books. IEEE Spectrum, 34(9):24T10– 96T4, September 1997. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1997:RBi

[Ano97-79] Anonymous. Recent books. IEEE Spectrum, 34(10):14– 72T2, October 1997. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1997:RBj

[Ano97-80] Anonymous. Recent books.

IEEE Spectrum, 34(11):18– 20T14, November 1997. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1997:RBk

[Ano97-81] Anonymous. Recent books.

IEEE Spectrum, 34(12):15– 64T4, December 1997. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1997:SEC

[Ano97-82] Anonymous. Seeing-eye cane steers the blind. IEEE Spectrum, 34(10):17, October 1997. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1997:SD

[Ano97-83] Anonymous. The standard diaries. IEEE Spectrum, 34(11): 49–61, November 1997. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1997:TCa

[Ano97-84] Anonymous. Table of contents. IEEE Spectrum, 34 (1):2–4, January 1997. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1997:TCb

[Ano97-85] Anonymous. Table of contents. IEEE Spectrum, 34 (2):2–4, February 1997. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1997:TCc

[Ano97-86] Anonymous. Table of contents. IEEE Spectrum, 34 (3):2–4, March 1997. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1997:TCd

[Ano97-87] Anonymous. Table of contents. IEEE Spectrum, 34 (4):2–4, April 1997. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1997:TCe

[Ano97-88] Anonymous. Table of contents. IEEE Spectrum, 34 (5):2–4, May 1997. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1997:TCf

[Ano97-89] Anonymous. Table of contents. IEEE Spectrum, 34 (6):2–4, June 1997. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1997:TCg

[Ano97-90] Anonymous. Table of contents. IEEE Spectrum, 34 (7):2–4, July 1997. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1997:TCh

[Ano97-91] Anonymous. Table of contents. IEEE Spectrum, 34 (8):2–4, August 1997. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1997:TCi

[Ano97-92] Anonymous. Table of contents. IEEE Spectrum, 34(9): 2–4, September 1997. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1997:TCj

[Ano97-93] Anonymous. Table of contents. IEEE Spectrum, 34 (10):2–4, October 1997. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1997:TCk

[Ano97-94] Anonymous. Table of contents. IEEE Spectrum, 34(11): 2–4, November 1997. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1997:TCl

[Ano97-95] Anonymous. Table of contents. IEEE Spectrum, 34(12): 2–4, December 1997. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1997:TTR

[Ano97-96] Anonymous. Taboo or not taboo? [review of Forbidden knowledge: from Prometheus to pornography, (Shattuck, R.; 1996). IEEE Spectrum, 34 (7):14–15, July 1997. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1997:PFa

[Ano97-97] Anonymous. To probe further. IEEE Spectrum, 34(1): 103–104, January 1997. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1997:PFb

[Ano97-98] Anonymous. To probe further. IEEE Spectrum, 34(2): 77–80, February 1997. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1997:PFc

[Ano97-99] Anonymous. To probe further. IEEE Spectrum, 34 (3):51, March 1997. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1997:PFd

[Ano97-100] Anonymous. To probe further. IEEE Spectrum, 34(9): 69, September 1997. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1997:WS

[Ano97-101] Anonymous. Web sites. IEEE Spectrum, 34(10):88, October 1997. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1998:B

[Ano98a] Anonymous. Books. IEEE Spectrum, 35(2):10–12, February 1998. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1998:ISFa

[Ano98b] Anonymous. IEEE Spectrum — front cover. IEEE Spectrum, 35(1):c1, January 1998. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1998:ISFb

[Ano98c] Anonymous. IEEE Spectrum — front cover. IEEE Spectrum, 35(2):c1, February 1998. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1998:ISFc

[Ano98d] Anonymous. IEEE Spectrum — front cover. IEEE Spectrum, 35(3):c1, March 1998.

CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1998:ISFd

[Ano98e] Anonymous. IEEE Spectrum — front cover. IEEE Spectrum, 35(4):c1, April 1998. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1998:ISFe

[Ano98f] Anonymous. IEEE Spectrum — front cover. IEEE Spectrum, 35(5):c1, May 1998. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1998:ISFf

[Ano98g] Anonymous. IEEE Spectrum — front cover. IEEE Spectrum, 35(6):c1, June 1998. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1998:ISFg

[Ano98h] Anonymous. IEEE Spectrum — front cover. IEEE Spectrum, 35(7):c1, July 1998. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1998:ISFh

[Ano98i] Anonymous. IEEE Spectrum — front cover. IEEE Spectrum, 35(8):c1, August 1998. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1998:ISFi

[Ano98j]  Anonymous.  IEEE Spec-

trum — front cover. IEEE Spectrum, 35(9):c1, September 1998. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1998:ISFj

[Ano98k]  Anonymous.  IEEE Spec-

trum — front cover. IEEE Spectrum, 35(10):c1, October 1998. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1998:ISFk

[Ano98l]  Anonymous.  IEEE Spec-

trum — front cover. IEEE Spectrum, 35(11):c1, November 1998. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1998:ISFl

[Ano98m]  Anonymous.  IEEE Spec-

trum — [front cover]. IEEE Spectrum, 35(12):c1, December 1998. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1998:Ca

[Ano98n]  Anonymous.  Calendar.

IEEE Spectrum, 35(1):12– 108E4, January 1998. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1998:Cb

[Ano98o]  Anonymous.  Calendar.

IEEE Spectrum, 35(2):13– 68E7, February 1998. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1998:Cc

[Ano98p]  Anonymous.  Calendar.

IEEE Spectrum, 35(3):17– 68E7, March 1998. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1998:Cd

[Ano98q]  Anonymous.  Calendar.

IEEE Spectrum, 35(4):16– 72E4, April 1998. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1998:Ce

[Ano98r]  Anonymous.  Calendar.

IEEE Spectrum, 35(5):14– 60E4, May 1998. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1998:Cf

[Ano98s]  Anonymous.  Calendar.

IEEE Spectrum, 35(6):17– 20E4, June 1998. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1998:Cg

[Ano98t] Anonymous. Calendar. IEEE Spectrum, 35(8):11–76E4, August 1998. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1998:Ch

[Ano98u]  Anonymous.  Calendar.

IEEE Spectrum, 35(9):10– 80E7, September 1998. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1998:Ci

[Ano98v]  Anonymous.  Calendar.

IEEE Spectrum, 35(10):16– 80E2, October 1998. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1998:Cj

[Ano98w]  Anonymous.  Calendar.

IEEE Spectrum, 35(11):16– 60E3, November 1998. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1998:Ck

[Ano98x]  Anonymous.  Calendar.

IEEE Spectrum, 35(12):13– 64E6, December 1998. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1998:CSa

[Ano98y] Anonymous. Coming in spectrum. IEEE Spectrum, 35 (1):130, January 1998. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1998:CSb

[Ano98z] Anonymous. Coming in spectrum. IEEE Spectrum, 35 (2):86, February 1998. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1998:CSc

[Ano98-27] Anonymous. Coming in spectrum. IEEE Spectrum, 35 (3):86, March 1998. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1998:CSd

[Ano98-28] Anonymous. Coming in spectrum. IEEE Spectrum, 35 (4):90, April 1998. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1998:CSe

[Ano98-29] Anonymous. Coming in spectrum. IEEE Spectrum, 35 (5):78, May 1998. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1998:CSf

[Ano98-30] Anonymous. Coming in spectrum. IEEE Spectrum, 35 (6):98, June 1998. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1998:CSg

[Ano98-31] Anonymous. Coming in spectrum. IEEE Spectrum, 35 (7):94, July 1998. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1998:CSh

[Ano98-32] Anonymous. Coming in spectrum. IEEE Spectrum, 35 (8):94, August 1998. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1998:CSi

[Ano98-33] Anonymous. Coming in spectrum. IEEE Spectrum, 35 (9):98, September 1998. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1998:CSj

[Ano98-34] Anonymous. Coming in spectrum. IEEE Spectrum, 35 (10):102, October 1998. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1998:CSk

[Ano98-35] Anonymous. Coming in spectrum. IEEE Spectrum, 35 (11):86, November 1998. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1998:CSl

[Ano98-36] Anonymous. Coming in spectrum. IEEE Spectrum, 35 (12):98, December 1998. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1998:ETTa

[Ano98-37] Anonymous. EE’s tools and toys. IEEE Spectrum, 35(2): 67–68, February 1998. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1998:ETTb

[Ano98-38] Anonymous. EE’s tools and toys. IEEE Spectrum, 35 (3):66–68, March 1998. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1998:ETTc

[Ano98-39] Anonymous. EE’s tools and toys. IEEE Spectrum, 35 (4):69–71, April 1998. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1998:ETTd

[Ano98-40] Anonymous. EE’s tools and toys. IEEE Spectrum, 35 (5):61–64, May 1998. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1998:ETTe

[Ano98-41] Anonymous. EE’s tools and toys. IEEE Spectrum, 35 (6):82–84, June 1998. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1998:ETTf

[Ano98-42] Anonymous. EE’s tools and toys. IEEE Spectrum, 35 (7):77–80, July 1998. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1998:ETTg

[Ano98-43] Anonymous. EE’s tools and toys. IEEE Spectrum, 35(8): 87–89, August 1998. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1998:ETTh

[Ano98-44] Anonymous. EE’s tools and toys. IEEE Spectrum, 35(9): 81–83, September 1998. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1998:ETTi

[Ano98-45] Anonymous. EE’s tools and toys. IEEE Spectrum, 35(10): 81–84, October 1998. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1998:ETTj

[Ano98-46] Anonymous. EE’s tools and toys. IEEE Spectrum, 35(11): 57–60, November 1998. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1998:ETTk

[Ano98-47] Anonymous. EEs’ tools toys. IEEE Spectrum, 35(12):65– 68, December 1998. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1998:EWa

[Ano98-48] Anonymous. EV watch. IEEE Spectrum, 35(3):22–23, March 1998. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1998:EWb

[Ano98-49] Anonymous. EV watch. IEEE Spectrum, 35(12):16, December 1998. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1998:Fa

[Ano98-50] Anonymous. Forum. IEEE Spectrum, 35(1):6, January 1998. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1998:Fb

[Ano98-51] Anonymous. Forum. IEEE Spectrum, 35(2):6–8, February 1998. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1998:Fc

[Ano98-52] Anonymous. Forum. IEEE Spectrum, 35(3):6–8, March 1998. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1998:Fd

[Ano98-53] Anonymous. Forum. IEEE Spectrum, 35(4):6–10, April 1998. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1998:Fe

[Ano98-54] Anonymous. Forum. IEEE Spectrum, 35(5):6–8, May 1998. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1998:Ff

[Ano98-55] Anonymous. Forum. IEEE Spectrum, 35(6):6–8, June 1998. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1998:Fg

[Ano98-56] Anonymous. Forum. IEEE Spectrum, 35(7):6–8, July 1998. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1998:Fh

[Ano98-57] Anonymous. Forum. IEEE Spectrum, 35(8):6–7, August 1998. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1998:Fi

[Ano98-58] Anonymous. Forum. IEEE Spectrum, 35(9):6–8, September 1998. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1998:Fj

[Ano98-59] Anonymous. Forum. IEEE Spectrum, 35(10):6–10, October 1998. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1998:Fk

[Ano98-60] Anonymous. Forum. IEEE Spectrum, 35(11):6–10, November 1998. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1998:Fl

[Ano98-61] Anonymous. Forum. IEEE Spectrum, 35(12):6–8, December 1998. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1998:ITF

[Ano98-62] Anonymous. The IEEE 1998 Technical Field Awards. IEEE Spectrum, 35(8):52–61, August 1998. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1998:IMA

[Ano98-63]  Anonymous.  The IEEE

Medalists [awards]. IEEE Spectrum, 35(6):62–67, June 1998. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1998:Na

[Ano98-64] Anonymous. Newslog. IEEE Spectrum, 35(1):1, January 1998. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1998:Nb

[Ano98-65] Anonymous. Newslog. IEEE Spectrum, 35(2):1, February 1998. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1998:Nc

[Ano98-66] Anonymous. Newslog. IEEE Spectrum, 35(3):1, March 1998. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1998:Nd

[Ano98-67] Anonymous. Newslog. IEEE Spectrum, 35(4):1, April 1998. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1998:Ne

[Ano98-68] Anonymous. Newslog. IEEE Spectrum, 35(5):1, May 1998. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1998:Nf

[Ano98-69] Anonymous. Newslog. IEEE Spectrum, 35(6):1, June 1998.

CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1998:Ng

[Ano98-70] Anonymous. Newslog. IEEE Spectrum, 35(7):1, July 1998. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1998:Nh

[Ano98-71] Anonymous. Newslog. IEEE Spectrum, 35(8):1, August 1998. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1998:Ni

[Ano98-72] Anonymous. Newslog. IEEE Spectrum, 35(9):1, September 1998. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1998:Nj

[Ano98-73] Anonymous. Newslog. IEEE Spectrum, 35(10):1, October 1998. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1998:Nk

[Ano98-74] Anonymous. Newslog. IEEE Spectrum, 35(11):1, November 1998. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1998:Nl

[Ano98-75] Anonymous. Newslog. IEEE Spectrum, 35(12):1, December 1998. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1998:PE

[Ano98-76] Anonymous. The practical engineer. IEEE Spectrum, 35(7):57, July 1998. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1998:RBa

[Ano98-77] Anonymous. Recent books.

IEEE Spectrum, 35(1):16T6– 108T4, January 1998. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1998:RBb

[Ano98-78] Anonymous. Recent books.

IEEE Spectrum, 35(2):12– 16T8, February 1998. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1998:RBc

[Ano98-79] Anonymous. Recent books.

IEEE Spectrum, 35(3):16– 68T2, March 1998. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1998:RBd

[Ano98-80] Anonymous. Recent books.

IEEE Spectrum, 35(4):15– 72T2, April 1998. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1998:RBe

[Ano98-81] Anonymous. Recent books. IEEE Spectrum, 35(5):20T5– 60T4, May 1998. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1998:RBf

[Ano98-82] Anonymous. Recent books. IEEE Spectrum, 35(6):16– 20T8, June 1998. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1998:RBg

[Ano98-83] Anonymous. Recent books. IEEE Spectrum, 35(7):76T3– 76T6, July 1998. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1998:RBh

[Ano98-84] Anonymous. Recent books. IEEE Spectrum, 35(8):12T4– 76T2, August 1998. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1998:RBi

[Ano98-85] Anonymous. Recent books. IEEE Spectrum, 35(9):16T6– 16T8, September 1998. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1998:RBj

[Ano98-86] Anonymous. Recent books.

IEEE Spectrum, 35(10):16T6– 80T2, October 1998. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1998:RBk

[Ano98-87] Anonymous. Recent books.

IEEE Spectrum, 35(11):14– 60T6, November 1998. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1998:RBl

[Ano98-88] Anonymous. Recent books.

IEEE Spectrum, 35(12):12– 16T4, December 1998. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1998:SIR

[Ano98-89] Anonymous. State of the Internet: Roundtable 4.0 [communications]. IEEE Spectrum, 35(10):69–79, October 1998. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1998:TCa

[Ano98-90] Anonymous. Table of contents. IEEE Spectrum, 35 (1):2–4, January 1998. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1998:TCb

[Ano98-91] Anonymous. Table of contents. IEEE Spectrum, 35 (2):2–4, February 1998. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1998:TCc

[Ano98-92] Anonymous. Table of contents. IEEE Spectrum, 35 (3):2–4, March 1998. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1998:TCd

[Ano98-93] Anonymous. Table of contents. IEEE Spectrum, 35 (4):2–4, April 1998. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1998:TCe

[Ano98-94] Anonymous. Table of contents. IEEE Spectrum, 35 (5):2–4, May 1998. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1998:TCf

[Ano98-95] Anonymous. Table of contents. IEEE Spectrum, 35 (6):2–4, June 1998. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1998:TCg

[Ano98-96] Anonymous. Table of contents. IEEE Spectrum, 35 (7):2–4, July 1998. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1998:TCh

[Ano98-97] Anonymous. Table of contents. IEEE Spectrum, 35 (8):2–4, August 1998. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1998:TCi

[Ano98-98] Anonymous. Table of contents. IEEE Spectrum, 35(9): 2–4, September 1998. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1998:TCj

[Ano98-99] Anonymous. Table of contents. IEEE Spectrum, 35 (10):2–4, October 1998. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1998:TCk

[Ano98-100] Anonymous. Table of contents. IEEE Spectrum, 35(11): 2–4, November 1998. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1998:TCl

[Ano98-101] Anonymous. Table of contents. IEEE Spectrum, 35(12): 2–4, December 1998. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1998:PF

[Ano98-102] Anonymous. To probe further. IEEE Spectrum, 35(1): 104–105, January 1998. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1998:WAD

[Ano98-103] Anonymous. What’s ahead for design on the web. IEEE Spectrum, 35(9):53–63, September 1998. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1999:MTI

[Ano99a] Anonymous. The 100-million transistor IC. IEEE Spectrum, 36(7):22–24, July 1999. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1999:IM

[Ano99b] Anonymous. The 1999 IEEE Medalists. IEEE Spectrum, 36 (6):62–68, June 1999. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1999:ISFa

[Ano99c] Anonymous. IEEE Spectrum — [front cover]. IEEE Spectrum, 36(1):c1, January 1999. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1999:ISFb

[Ano99d] Anonymous. IEEE Spectrum — [front cover]. IEEE Spectrum, 36(2):c1, February 1999. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1999:ISFc

[Ano99e] Anonymous. IEEE Spectrum — [front cover]. IEEE Spectrum, 36(3):c1, March 1999. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1999:ISFd

[Ano99f] Anonymous. IEEE Spectrum — [front cover]. IEEE Spectrum, 36(4):c1, April 1999. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1999:ISFe

[Ano99g] Anonymous. IEEE Spectrum — [front cover]. IEEE Spectrum, 36(5):c1, May 1999. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1999:ISFf

[Ano99h] Anonymous. IEEE Spectrum — [front cover]. IEEE Spectrum, 36(6):c1, June 1999. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1999:ISFg

[Ano99i] Anonymous. IEEE Spectrum — [front cover]. IEEE Spectrum, 36(7):c1, July 1999. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1999:ISFh

[Ano99j] Anonymous. IEEE Spectrum — [front cover]. IEEE Spectrum, 36(8):c1, August 1999. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1999:ISFi

[Ano99k]  Anonymous.  IEEE Spec-

trum — [front cover]. IEEE Spectrum, 36(9):c1, September 1999. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1999:ISFj

[Ano99l]  Anonymous.  IEEE Spec-

trum — [front cover]. IEEE Spectrum, 36(10):c1, October 1999. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1999:ISFk

[Ano99m]  Anonymous.  IEEE Spec-

trum — [front cover]. IEEE Spectrum, 36(11):c1, November 1999. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1999:ISFl

[Ano99n]  Anonymous.  IEEE Spec-

trum — [front cover]. IEEE Spectrum, 36(12):c1, December 1999. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1999:Ca

[Ano99o]  Anonymous.  Calendar.

IEEE Spectrum, 36(1):14– 112E2, January 1999. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1999:Cb

[Ano99p]  Anonymous.  Calendar.

IEEE Spectrum, 36(2):14– 64E7, February 1999. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1999:Cc

[Ano99q]  Anonymous.  Calendar.

IEEE Spectrum, 36(3):13– 64E7, March 1999. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1999:Cd

[Ano99r]  Anonymous.  Calendar.

IEEE Spectrum, 36(4):19– 78R9, April 1999. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1999:Ce

[Ano99s] Anonymous. Calendar. IEEE Spectrum, 36(5):18–66R15, May 1999. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1999:Cf

[Ano99t] Anonymous. Calendar. IEEE Spectrum, 36(6):23–82R13, June 1999. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1999:Cg

[Ano99u] Anonymous. Calendar. IEEE Spectrum, 36(7):16–88R16, July 1999. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1999:Ch

[Ano99v] Anonymous. Calendar. IEEE Spectrum, 36(8):16–72R4, August 1999. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1999:Ci

[Ano99w]  Anonymous.  Calendar.

IEEE Spectrum, 36(9):20– 26I8, September 1999. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1999:Cj

[Ano99x]  Anonymous.  Calendar.

IEEE Spectrum, 36(10):20– 74R9, October 1999. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1999:Ck

[Ano99y]  Anonymous.  Calendar.

IEEE Spectrum, 36(11):19– 26, November 1999. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1999:Cl

[Ano99z]  Anonymous.  Calendar.

IEEE Spectrum, 36(12):14– 24I8, December 1999. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1999:CSa

[Ano99-27] Anonymous. Coming in spectrum. IEEE Spectrum, 36 (2):82, February 1999. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1999:CSb

[Ano99-28] Anonymous. Coming in spectrum. IEEE Spectrum, 36 (3):82, March 1999. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1999:ETTa

[Ano99-29] Anonymous. EE’s tools and toys. IEEE Spectrum, 36(1): 132–135, January 1999. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1999:ETTe

[Ano99-30] Anonymous. EEs’ tools toy. IEEE Spectrum, 36(5):74–77, May 1999. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1999:ETTf

[Ano99-31]  Anonymous.  EEs’ tools

toy. IEEE Spectrum, 36(6): 100–103, June 1999. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1999:ETTg

[Ano99-32] Anonymous. EEs’ tools toy. IEEE Spectrum, 36(7):82–87, July 1999. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1999:ETTh

[Ano99-33]  Anonymous.  EEs’ tools

toy. IEEE Spectrum, 36(8): 70–72, August 1999. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1999:ETTi

[Ano99-34] Anonymous. EEs’ tools toy. IEEE Spectrum, 36(9):87– 90, September 1999. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1999:ETTj

[Ano99-35] Anonymous. EEs’ tools toy.

IEEE Spectrum, 36(10):70– 74, October 1999. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1999:ETTk

[Ano99-36] Anonymous. EEs’ tools toy.

IEEE Spectrum, 36(11):78– 81, November 1999. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1999:ETTl

[Ano99-37] Anonymous. EEs’ tools toy.

IEEE Spectrum, 36(12):78– 79, December 1999. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1999:ETTb

[Ano99-38] Anonymous. EEs’ tools toys.

IEEE Spectrum, 36(2):62– 63, February 1999. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1999:ETTc

[Ano99-39]  Anonymous.  EEs’ tools

toys. IEEE Spectrum, 36 (3):60–64, March 1999. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1999:ETTd

[Ano99-40]  Anonymous.  EEs’ tools

toys. IEEE Spectrum, 36 (4):86–89, April 1999. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1999:EW

[Ano99-41] Anonymous. EV watch. IEEE Spectrum, 36(3):17–18, March 1999. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1999:Fa

[Ano99-42] Anonymous. Forum. IEEE Spectrum, 36(1):6–8, January 1999. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1999:Fb

[Ano99-43] Anonymous. Forum. IEEE Spectrum, 36(2):6–8, February 1999. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1999:Fc

[Ano99-44] Anonymous. Forum. IEEE Spectrum, 36(3):6–8, March 1999. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1999:Fd

[Ano99-45] Anonymous. Forum. IEEE Spectrum, 36(4):6–12, April 1999. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1999:Fe

[Ano99-46] Anonymous. Forum. IEEE Spectrum, 36(5):6–8, May 1999. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1999:Ff

[Ano99-47] Anonymous. Forum. IEEE Spectrum, 36(6):6–12, June 1999. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1999:Fg

[Ano99-48] Anonymous. Forum. IEEE Spectrum, 36(7):6–8, July 1999. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1999:Fh

[Ano99-49] Anonymous. Forum. IEEE Spectrum, 36(8):6–13, August 1999. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1999:Fi

[Ano99-50] Anonymous. Forum. IEEE Spectrum, 36(9):6–14, September 1999. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1999:Fj

[Ano99-51] Anonymous. Forum. IEEE Spectrum, 36(10):6–12, October 1999. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1999:Fk

[Ano99-52] Anonymous. Forum. IEEE Spectrum, 36(11):6–14, November 1999. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1999:Fl

[Ano99-53] Anonymous. Forum. IEEE Spectrum, 36(12):6–8, December 1999. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1999:ITF

[Ano99-54] Anonymous. The IEEE Technical Field Awards. IEEE Spectrum, 36(8):57–69, August 1999. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1999:I

[Ano99-55]  Anonymous.  Innovations.

IEEE Spectrum, 36(10):24– 26, October 1999. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1999:Na

[Ano99-56] Anonymous. Newslog. IEEE Spectrum, 36(1):1, January 1999. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1999:Nb

[Ano99-57] Anonymous. Newslog. IEEE Spectrum, 36(2):1, February 1999. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1999:Nc

[Ano99-58] Anonymous. Newslog. IEEE Spectrum, 36(3):1, March 1999. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1999:Nd

[Ano99-59] Anonymous. Newslog. IEEE Spectrum, 36(4):1, April 1999. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1999:Ne

[Ano99-60] Anonymous. Newslog. IEEE Spectrum, 36(5):1, May 1999. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1999:Nf

[Ano99-61] Anonymous. Newslog. IEEE Spectrum, 36(6):1, June 1999. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1999:Ng

[Ano99-62] Anonymous. Newslog. IEEE Spectrum, 36(7):1, July 1999. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1999:Nh

[Ano99-63] Anonymous. Newslog. IEEE Spectrum, 36(8):1, August 1999. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1999:Ni

[Ano99-64] Anonymous. Newslog. IEEE Spectrum, 36(9):1, September 1999. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1999:Nj

[Ano99-65] Anonymous. Newslog. IEEE Spectrum, 36(10):1, October 1999. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1999:Nk

[Ano99-66] Anonymous. Newslog. IEEE Spectrum, 36(11):1, November 1999. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1999:Nl

[Ano99-67] Anonymous. Newslog. IEEE Spectrum, 36(12):1, December 1999. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1999:RBa

[Ano99-68] Anonymous. Recent books. IEEE Spectrum, 36(1):13– 24T8, January 1999. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1999:RBb

[Ano99-69] Anonymous. Recent books. IEEE Spectrum, 36(2):16T4– 16T7, February 1999. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1999:RBc

[Ano99-70] Anonymous. Recent books.

IEEE Spectrum, 36(3):64T4, March 1999. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1999:RBd

[Ano99-71] Anonymous. Recent books.

IEEE Spectrum, 36(4):26I6– 26I8, April 1999. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1999:RBe

[Ano99-72] Anonymous. Recent books.

IEEE Spectrum, 36(5):66R5– 66R6, May 1999. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1999:RBf

[Ano99-73] Anonymous. Recent books.

IEEE Spectrum, 36(6):26I8, June 1999. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1999:RBg

[Ano99-74] Anonymous. Recent books.

IEEE Spectrum, 36(7):16I8, July 1999. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1999:RBh

[Ano99-75] Anonymous. Recent books. IEEE Spectrum, 36(8):72R8, August 1999. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1999:RBi

[Ano99-76] Anonymous. Recent books. IEEE Spectrum, 36(9):17– 90R15, September 1999. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1999:RBj

[Ano99-77] Anonymous. Recent books. IEEE Spectrum, 36(10):26I8, October 1999. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1999:RBk

[Ano99-78] Anonymous. Recent books. IEEE Spectrum, 36(11):26I6– 26I8, November 1999. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1999:RBl

[Ano99-79] Anonymous. Recent books. IEEE Spectrum, 36(12):24I4– 24I6, December 1999. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1999:SRa

[Ano99-80] Anonymous. Software reviews. IEEE Spectrum, 36 (4):73–78, April 1999. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1999:SRb

[Ano99-81] Anonymous. Software reviews. IEEE Spectrum, 36 (10):88, October 1999. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1999:TCa

[Ano99-82] Anonymous. Table of contents. IEEE Spectrum, 36 (1):2–4, January 1999. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1999:TCb

[Ano99-83] Anonymous. Table of contents. IEEE Spectrum, 36 (2):2–4, February 1999. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1999:TCc

[Ano99-84] Anonymous. Table of contents. IEEE Spectrum, 36 (3):2–4, March 1999. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1999:TCd

[Ano99-85] Anonymous. Table of contents. IEEE Spectrum, 36 (4):2–4, April 1999. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1999:TCe

[Ano99-86] Anonymous. Table of contents. IEEE Spectrum, 36 (5):2–4, May 1999. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1999:TCf

[Ano99-87] Anonymous. Table of contents. IEEE Spectrum, 36 (6):2–4, June 1999. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1999:TCg

[Ano99-88] Anonymous. Table of contents. IEEE Spectrum, 36 (7):2–4, July 1999. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1999:TCh

[Ano99-89] Anonymous. Table of contents. IEEE Spectrum, 36 (8):2–4, August 1999. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1999:TCi

[Ano99-90] Anonymous. Table of contents. IEEE Spectrum, 36(9): 2–4, September 1999. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1999:TCj

[Ano99-91] Anonymous. Table of contents. IEEE Spectrum, 36 (10):2–4, October 1999. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1999:TCk

[Ano99-92] Anonymous. Table of contents. IEEE Spectrum, 36(11): 2–4, November 1999. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1999:TCl

[Ano99-93] Anonymous. Table of contents. IEEE Spectrum, 36(12): 2–4, December 1999. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1999:TS

[Ano99-94]  Anonymous.  Technically

speaking. IEEE Spectrum, 36 (12):25, December 1999. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1999:PF

[Ano99-95] Anonymous. To probe further. IEEE Spectrum, 36(1): 108–109, January 1999. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1999:WSa

[Ano99-96]  Anonymous.  Web sights.

IEEE Spectrum, 36(4):28, April 1999. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1999:WSb

[Ano99-97]  Anonymous.  Web sights.

IEEE Spectrum, 36(6):99, June 1999. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1999:WSc

[Ano99-98]  Anonymous.  Web sights.

IEEE Spectrum, 36(7):93, July 1999. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1999:WSd

[Ano99-99]  Anonymous.  Web sights.

IEEE Spectrum, 36(9):101, September 1999. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1999:WSe

[Ano99-100] Anonymous. Web sights. IEEE Spectrum, 36(10):91, October 1999. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1999:WSf

[Ano99-101] Anonymous. Web sights.

IEEE Spectrum, 36(12):82– 84, December 1999. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Apfelbaum:1997:SBT

[Apf97] L. Apfelbaum. Spec-based tests make sure telecom software works. IEEE Spectrum, 34(11):77–83, November 1997. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Ang:1991:VCM

[ARA91] P. H. Ang, P. A. Ruetz, and D. Auld. Video compression makes big gains. IEEE Spectrum, 28(10):16–19, October 1991. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Aspray:1990:BBS

[Asp90] William Aspray. Back to basics: The stored program concept. IEEE Spectrum, 27(9): 51, September 1990. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Aspray:1997:II

[Asp97] William Aspray. Inventing the Internet. IEEE Spectrum, 34 (2):12, February 1997. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Asthana:1994:LRO

[Ast94] P. Asthana. A long road to overnight success [optical disc storage]. IEEE Spectrum, 31(10):60–66, October 1994. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Asthana:1995:JTC

[Ast95] P. Asthana. Jumping the technology S-curve. IEEE Spectrum, 32(6):49–54, June 1995. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Atherton:1998:MJF

[Ath98] R. W. Atherton. Moving Java to the factory. IEEE Spectrum, 35(12):18–23, December 1998. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Atwood:1991:COO

[Atw91] T. M. Atwood. The case for object-oriented databases. IEEE Spectrum, 28(2):44– 47, February 1991. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anagnostopoulos:1998:FGS

[AW98] C. N. Anagnostopoulos and L. A. Williams. Few gold stars for precollege education. IEEE Spectrum, 35(4):18–26, April 1998. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Bail:1996:HBR

[Bai96] W. Bail. Out of humor [book review]. IEEE Spectrum, 33 (8):11–14, August 1996. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Balk:1991:EAP

[Bal91] A. Balk. East Asia’s power crescent. IEEE Spectrum, 28 (6):26–29, June 1991. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Baliga:1995:PIS

[Bal95] B. J. Baliga. Power ICs in the saddle. IEEE Spectrum, 32(7):34–40, July 1995. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Bell:1996:C

[BAL96] Trudy E. Bell, J. A. Adam, and S. J. Lowe. Communications. IEEE Spectrum, 33 (1):30–41, January 1996. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Balachandran:1997:SPS

[Bal97]  U. Balachandran.  Super

power(special report). IEEE Spectrum, 34(7):18–19, July 1997. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Barnett:1990:FWH

[Bar90] R. D. Barnett. The frequency that wouldn’t die hydroelectric generators. IEEE Spectrum, 27(11):120–121, November 1990. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Barclay:1992:DHT

[Bar92] D. Barclay. Diversity in the high-tech workplace: commitment from the top makes it work. IEEE Spectrum, 29 (6):24–27, June 1992. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Barbour:1998:CST

[Bar98] E. Barbour. The cool sound of tubes [vacuum tube musical applications]. IEEE Spectrum, 35(8):24–35, August 1998. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Batten:1992:DHT

[Bat92] E. Batten. Diversity in the high-tech workplace: forums for diversity. IEEE Spectrum, 29(6):27–29, June 1992. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Bates:1995:GFH

[Bat95] B. Bates. Getting a Ford HEV on the road. IEEE Spectrum, 32(7):22–25, July 1995. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Bhatt:1997:DTM

[BBH97] B. Bhatt, D. Birks, and D. Hermreck. Digital television: making it work. IEEE Spectrum, 34(10):19–28, October 1997. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Borrus:1993:BLE

[BC93] M. G. Borrus and S. S. Cohen.

Beyond lean: an essay (mass production). IEEE Spectrum, 30(9):67–68, September 1993. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Baldwin:1997:LS

[BC97a] R. W. Baldwin and C. V.

Chang. Locking the e-safe. IEEE Spectrum, 34(2):40– 46, February 1997. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Braham:1997:SVW

[BC97b] R. Braham and Richard Comerford. Sharing virtual worlds. IEEE Spectrum, 34 (3):18–19, March 1997. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Badami:1998:HAG

[BC98] V. V. Badami and N. W. Chbat. Home appliances get smart. IEEE Spectrum, 35 (8):36–43, August 1998. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Bretz:1999:TWC

[BC99] E. A. Bretz and L. Ceppert. Technical women on corporate boards. IEEE Spectrum, 36 (2):50–56, February 1999. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Basile:1995:UHS

[BCD^{+}95] C. Basile, A. P. Cavallerano, M. S. Deiss, R. Keeler, J. S. Lim, W. C. Luplow, W. H. Paik, E. Petajan, R. Rast, G. Reitmeier, T. R. Smith, and C. Todd. The US HDTV standard. IEEE Spectrum, 32 (4):36–45, April 1995. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Bishop:1997:TNI

[BCW97] M. Bishop, S. Cheung, and C. Wee. The threat from the net [Internet security]. IEEE Spectrum, 34(8):56–63, August 1997. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Bell:1999:TNM

[BDF99] Trudy E. Bell, D. Dooling, and J. Fouke. Threshold of the new millennium [book excerpt from Engineering Tomorrow: Today’s Technology Experts Envision the Next Century (IEEE Press, Piscataway, NJ, 2000) written by Trudy E. Bell and Dave Dooling, and edited by Janie Fouke]. IEEE Spectrum, 36(10):59–64, October 1999. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Bruce:1996:WAS

[BDG96] Andrew Bruce, David Donoho, and Hong-Ye Gao. Wavelet analysis [for signal processing]. IEEE Spectrum, 33(10): 26–35, October 1996. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Bell:1991:USS

[BE91] Trudy E. Bell and K. Esch.

The United States in space. IEEE Spectrum, 28(8):18– 20, August 1991. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Beard:1997:NWD

[Bea97] J. Beard. Nuclear waste disposal in New Mexico and Nevada. IEEE Spectrum, 34 (11):33–40, November 1997. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Bell:1990:ESB

[Bel90a] Trudy E. Bell. ‘90s employment: some bad news, but some good. IEEE Spectrum, 27(12):32–43, December 1990. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Bell:1990:AT

[Bel90b]  Trudy E. Bell.  Applica-

tions ’90: telecommunications. IEEE Spectrum, 27(2): 31–33, February 1990. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Bell:1990:MCL

[Bel90c] Trudy E. Bell. From monopoly to competition: long-term research is vulnerable. IEEE Spectrum, 27(10):46–50, October 1990. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Bell:1990:JKC

[Bel90d] Trudy E. Bell. Just keeping current entails ingenuity, resourcefulness, and improvisation (R&D). IEEE Spectrum, 27(10):69–73, October 1990. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Bell:1990:LAL

[Bel90e] Trudy E. Bell. Light that acts like ‘natural bits’ (solitons). IEEE Spectrum, 27 (8):56–57, August 1990. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Bell:1990:TCD

[Bel90f] Trudy E. Bell. Technical challenges to a decentralized phone system. IEEE Spectrum, 27(9):32–37, September 1990. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Bell:1990:TME

[Bel90g] Trudy E. Bell. Technology ’90: the main event (Voyager 2 project). IEEE Spectrum, 27 (1):28–29, January 1990. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Bell:1991:ISC

[Bel91a] Trudy E. Bell. Incredible shrinking computers. IEEE Spectrum, 28(5):37–41, May 1991. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Bell:1991:TT

[Bel91b]  Trudy E. Bell.  Technol-

ogy 1991: telecommunications. IEEE Spectrum, 28(1): 44–47, January 1991. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Bell:1991:TME

[Bel91c] Trudy E. Bell. Technology 1991: the media event Hubble Space Telescope. IEEE Spectrum, 28(1):31, January 1991. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Bell:1991:WBB

[Bel91d] Trudy E. Bell. When bust is best (destructive testing). IEEE Spectrum, 28 (3):56–59, March 1991. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Bell:1992:SBT

[Bel92a] Trudy E. Bell. Supercomputers — beyond today’s supercomputers. IEEE Spectrum, 29(9):72–75, September 1992. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Bell:1992:TT

[Bel92b]  Trudy E. Bell.  Technol-

ogy 1992 — telecommunications. IEEE Spectrum, 29(1): 36–38, January 1992. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Bell:1993:BPB

[Bel93a] Trudy E. Bell. Bicycles on a personalized basis FMS. IEEE Spectrum, 30(9):32– 35, September 1993. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Bell:1993:JR

[Bel93b]  Trudy E. Bell.  Jobs at

risk. IEEE Spectrum, 30(8): 18–31, August 1993. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Bell:1993:TNT

[Bel93c] Trudy E. Bell. Technology 1993 — new technology directions. IEEE Spectrum, 30 (1):81, January 1993. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Bell:1993:TT

[Bel93d]  Trudy E. Bell.  Technol-

ogy 1993 — telecommunications. IEEE Spectrum, 30(1): 38–41, January 1993. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Bell:1993:TSM

[Bel93e] Trudy E. Bell. Technology 1993 — the specialties (multimedia, magnetic storage, education). IEEE Spectrum, 30 (1):80, January 1993. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Bellefeuille:1993:TQM

[Bel93f] J. Bellefeuille. Total quality management. IEEE Spectrum, 30(9):47–50, September 1993. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Bell:1994:AYC

[Bel94a] Trudy E. Bell. Alfred Yi Cho [biography]. IEEE Spectrum, 31(10):70–73, October 1994. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Bell:1994:AFI

[Bel94b]  Trudy E. Bell.  Analysis

and forecast issue: technology 1994. IEEE Spectrum, 31 (1):20–21, January 1994. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Bell:1994:ELF

[Bel94c] Trudy E. Bell. Engineering layoffs: facts and myths. IEEE Spectrum, 31(11):16– 25, November 1994. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Bell:1994:T

[Bel94d] Trudy E. Bell. Telecommunications. IEEE Spectrum, 31 (1):22–25, January 1994. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Bell:1995:EST

[Bel95a] Trudy E. Bell. Electronics and the stars [telescope control]. IEEE Spectrum, 32(8): 16–24, August 1995. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Bell:1995:HTK

[Bel95b] Trudy E. Bell. The HDTV ‘test kitchens’. IEEE Spectrum, 32(4):46–49, April 1995. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Bell:1995:RS

[Bel95c] Trudy E. Bell. Remote sensing. IEEE Spectrum, 32(3): 24–31, March 1995. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Bell:1995:USTa

[Bel95d] Trudy E. Bell. US science and technology museums — I. IEEE Spectrum, 32(9):50– 71, September 1995. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Bell:1995:USTb

[Bel95e] Trudy E. Bell. US science and technology museums — II. IEEE Spectrum, 32(10): 48–71, October 1995. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Bell:1996:BIA

[Bel96a] Trudy E. Bell. The Beckman Institute for Advanced Science and Technology. IEEE Spectrum, 33(11):54–60, November 1996. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Bell:1996:CLD

[Bel96b] Trudy E. Bell. The Challenger Launch Decision: Risky Technology, Culture, and Deviance at NASA [book review]. IEEE Spectrum, 33 (6):10–84T6, June 1996. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Bell:1996:ERS

[Bel96c] Trudy E. Bell. Employment roundtable: Survival calls for more than technical fitness [careers]. IEEE Spectrum, 33 (3):20–31, March 1996. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Bell:1996:ISTa

[Bel96d] Trudy E. Bell. International science and technology museums. I. IEEE Spectrum, 33 (4):39–49, April 1996. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Bell:1996:ISTb

[Bel96e] Trudy E. Bell. International science and technology museums. II. IEEE Spectrum, 33 (6):74–83, June 1996. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Bell:1998:LSB

[Bel98] Trudy E. Bell. Lessons from the shuttle [books]. IEEE Spectrum, 35(6):14–16, June 1998. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Bell:1999:DDD

[Bel99] A. E. Bell. The dynamic digital disk. IEEE Spectrum, 36(10):28–35, October 1999. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Bendz:1993:IAM

[Ben93] D. J. Bendz. International actions (manufacturing practices). IEEE Spectrum, 30(9): 63–66, September 1993. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Benedetto:1998:ECI

[Ben98] J. M. Benedetto. Economyclass ion-defying ICs in orbit. IEEE Spectrum, 35(3): 36–41, March 1998. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Berkeley:1997:NTF

[Ber97] A. R. Berkeley. Nasdaq’s technology floor: its president takes stock. IEEE Spectrum, 34(2):66–67, February 1997. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Berrie:1999:DDP

[Ber99] J. Berrie. The defensive design of printed-circuit boards.

IEEE Spectrum, 36(9):76– 81, September 1999. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Betts:1999:ETA

[Bet99] K. S. Betts. The environment [1999 technology analysis and forecast]. IEEE Spectrum, 36(1):104–107, January 1999. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Bodson:1992:WLG

[BH92] D. Bodson and E. Harris. When the lines go down (telecommunications). IEEE Spectrum, 29(3):40–44, March 1992. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Byrd:1995:MPA

[BH95] G. T. Byrd and M. A. Holliday. Multithreaded processor architectures. IEEE Spectrum, 32(8):38–46, August 1995. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Barkovich:1996:CNC

[BH96] B. R. Barkovich and D. V. Hawk. Charting a new course in California [electricity market restructuring]. IEEE Spectrum, 33(7):26–31, July 1996. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Bhattacharya:1994:HWR

[Bha94] T. K. Bhattacharya. Homemade watts for rural India. IEEE Spectrum, 31(10): 44–46, October 1994. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Bilmes:1996:UFN

[Bil96]  J. Bilmes.  User-friendly

neural-net design [software reviews]. IEEE Spectrum, 33 (2):63, February 1996. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Bisat:1995:BE

[Bis95] M. T. Bisat. A broad education. IEEE Spectrum, 32 (12):6, December 1995. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Boothroyd:1993:DA

[BK93] G. Boothroyd and W. Knight.

Design for assembly. IEEE Spectrum, 30(9):53–55, September 1993. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Benner:1999:PGG

[BK99] J. P. Benner and L. Kazmerski. Photovoltaics gaining greater visibility. IEEE Spectrum, 36(9):34–42, September 1999. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Blackett:1996:GGC

[Bla96] R. K. Blackett. As good as gold [circuit design]. IEEE Spectrum, 33(6):68–71, June 1996. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Blaugher:1997:LCH

[Bla97] R. D. Blaugher. Low-calorie, high-energy generators and motors. IEEE Spectrum, 34 (7):36–42, July 1997. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Blineau:1996:VDR

[Bli96] J. Blineau. Viewpoint: The digital revolution is complete. IEEE Spectrum, 33 (1):49, January 1996. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Baltes:1998:ENL

[BLK98] H. Baltes, D. Lange, and A. Koll. The electronic nose in Lilliput. IEEE Spectrum, 35(9):35–38, September 1998. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Biddle:1994:RR

[BM94] M. B. Biddle and R. Mann. Recipe for recycling. IEEE Spectrum, 31(8):22–24, August 1994. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Boehmer:1996:VTU

[Boe96] L. S. Boehmer. Viewpoint:

Transportation unbounded. IEEE Spectrum, 33(1):84– 85, January 1996. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Booth:1999:DTU

[Boo99] S. A. Booth. Digital TV in the United States. IEEE Spectrum, 36(3):39–41, March 1999. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Borsook:1994:DC

[Bor94] P. Borsook. Data communications. IEEE Spectrum, 31 (1):26–29, January 1994. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Boxer:1995:WBC

[Box95] A. Boxer. Where buses cannot go. IEEE Spectrum, 32 (2):41–45, February 1995. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Bowles:1992:DSB

[BP92] J. B. Bowles and C. E.

Pelaez. Data security — bad code. IEEE Spectrum, 29(8): 36–40, August 1992. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Bechtold:1998:BNP

[BP98] K. Bechtold and M. Pokojski. Berlin’s new powerhouse [combined cycle power plant]. IEEE Spectrum, 35 (3):52–57, March 1998. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Bryzek:1994:MM

[BPM94] J. Bryzek, K. Peterson, and W. McCulley. Micromachines on the march. IEEE Spectrum, 31(5):20–31, May 1994. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Bell:1997:CTA

[BR97] Trudy E. Bell and M. J. Riezenman. Communications [technology analysis and forecast]. IEEE Spectrum, 34(1): 27–37, January 1997. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Braham:1995:AS

[Bra95a] R. Braham. Application software. IEEE Spectrum, 32(1): 66–70, January 1995. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Braham:1995:DBC

[Bra95b] R. Braham. The digital backlot [cinema special effects]. IEEE Spectrum, 32(7):51–63, July 1995. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Braham:1995:MVN

[Bra95c] R. Braham. Math and visualization: new tools, new frontiers. IEEE Spectrum, 32(11): 19–28, November 1995. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Braham:1996:AFR

[Bra96a] R. Braham. The Air Force R&D balancing act. IEEE Spectrum, 33(3):45–51, March 1996. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Braham:1996:BTD

[Bra96b] R. Braham. Beyond the third dimension [books]. IEEE Spectrum, 33(11):18, November 1996. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Braham:1996:CE

[Bra96c] R. Braham. Consumer electronics. IEEE Spectrum, 33 (1):46–50, January 1996. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Braham:1998:FF

[Bra98]  R. Braham.  Faults fail-

ures. IEEE Spectrum, 35(8): 83–84, August 1998. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Braham:1999:AMT

[Bra99] R. Braham. Aerospace and military [technology 1999 analysis and forecast]. IEEE Spectrum, 36(1):73–78, January 1999. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Bretz:1999:TTA

[Bre99] E. A. Bretz. Transportation [1999 technology analysis and forecast]. IEEE Spectrum, 36(1):98–103, January 1999. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Brittain:1990:TT

[Bri90] J. E. Brittain. Thevenin’s theorem. IEEE Spectrum, 27 (3):42, March 1990. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Brody:1995:TA

[Bro95] G. Brody. Technology for all. IEEE Spectrum, 32(12): 8–??, December 1995. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Brockhurst:1998:SR

[Bro98] F. C. Brockhurst. Software reviews. IEEE Spectrum, 35 (2):16, February 1998. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Brown:1999:PRB

[Bro99] L. Brown. Pioneers of radar [book review]. IEEE Spectrum, 36(11):17–18, November 1999. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Brunet:1994:HLL

[Bru94] C. J. Brunet. Hybridizing the local loop. IEEE Spectrum, 31(6):28–32, June 1994. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Bryzek:1996:NES

[Bry96] J. Bryzek. Nano: The Emerging Science of Nanotechnology: Remaking the world — Molecule by Molecule [book review]. IEEE Spectrum, 33 (6):8–11, June 1996. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Bretz:1999:ACC

[BS99] E. A. Bretz and W. Sweet. Addressing the coal conundrum. IEEE Spectrum, 36(12):40– 41, December 1999. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Brust:1993:MSV

[BT93] L. Brust and M.-S. Tsay. Mixing signals and voltages on chip. IEEE Spectrum, 30(8): 40–43, August 1993. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Buc:1995:BE

[Buc95] S. M. Buc. A broad education. IEEE Spectrum, 32 (12):6, December 1995. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Burwen:1993:KOP

[Bur93] R. S. Burwen. Kilowatts on order (power amplifiers). IEEE Spectrum, 30(2):32–37, February 1993. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Butler:1996:RNT

[But96] L. J. Butler. Resistance to new technology: Nuclear power, information technology, and biotechnology [book review]. IEEE Spectrum, 33(6):8, June 1996. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Butrica:1997:CVP

[But97] A. J. Butrica. In conjunction with Venus [planetary radar astronomy]. IEEE Spectrum, 34(12):31–38, December 1997. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Byrne:1997:EJP

[Byr97] K. Byrne. Evaluating jukebox performance. IEEE Spectrum, 34(7):70–73, July 1997. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Caird:1997:ISA

[Cai97] K. Caird. Integrating substation automation. IEEE Spectrum, 34(8):64–69, August 1997. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Campanella:1990:CSO

[Cam90] S. J. Campanella. Communications satellites: orbiting into the 90s. IEEE Spectrum, 27 (8):49–52, August 1990. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Camenzind:1997:ROT

[Cam97] H. R. Camenzind. Redesigning the old 555 [timer circuit]. IEEE Spectrum, 34(9): 80–85, September 1997. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Carroll:1993:VPH

[Car93] M. Carroll. VHDL — panacea or hype? IEEE Spectrum, 30(6):34–37, June 1993. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cates:1990:GAF

[Cat90] R. Cates. Gallium arsenide finds a new niche. IEEE Spectrum, 27(4):25–28, April 1990. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Catlett:1992:SBR

[Cat92] C. E. Catlett. Supercomputing — balancing resources. IEEE Spectrum, 29(9):48– 55, September 1992. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Chaum:1997:MEC

[CB97] D. Chaum and S. Brands. ‘Minting’ electronic cash. IEEE Spectrum, 34(2):30– 34, February 1997. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Crawford:1999:SFF

[CB99] R. H. Crawford and J. J. Beaman. Solid freeform fabrication. IEEE Spectrum, 36(2): 34–43, February 1999. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Carreira:1999:FIS

[CCS99] J. V. Carreira, D. Costa, and J. G. Silva. Fault injection spot-checks computer system dependability. IEEE Spectrum, 36(8):50–55, August 1999. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Collins:1999:DEG

[CDEJ99] J. C. Collins, J. Dunn, P. Emer, and M. Johnson.

Data express gigabit junction with the next-generation Internet. IEEE Spectrum, 36(2): 18–25, February 1999. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cendes:1990:ESC

[Cen90] Z. J. Cendes. EM simulators = CAE tools. IEEE Spectrum, 27(11):73–77, November 1990. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cerchione:1992:ESD

[Cer92] R. G. Cerchione. Engineering software — data acquisition made simpler. IEEE Spectrum, 29(11):66–71, November 1992. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Comerford:1996:T

[CG96] Richard Comerford and L. Geppert. Technology 1996. IEEE Spectrum, 33(1):27–29, January 1996. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Chang:1997:STA

[Cha97a] L. L. Chang. Static timing analysis of high-speed boards. IEEE Spectrum, 34 (3):67–74, March 1997. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Chao:1997:TAC

[Cha97b] R. L. Chao. Trimming analog circuits with solid-state devices. IEEE Spectrum, 34(10): 63–67, October 1997. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Chappell:1999:FAI

[Cha99] B. Chappell. The fine art of IC design. IEEE Spectrum, 36(7):30–34, July 1999. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Chen:1990:ASO

[Che90a] K. T. Chen. Applications ’90: Soviet optoelectronics. IEEE Spectrum, 27(2):44–45, February 1990. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Chen:1990:ASE

[Che90b] K. T. Chen. Applications ’90: the specialities (electrical engineering advances). IEEE Spectrum, 27(2):46–47, February 1990. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Chen:1990:CSP

[Che90c] K. T. Chen. Contrasting strategies are pursued by big three economic powerhouses. IEEE Spectrum, 27(10):76– 78, October 1990. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Chen:1990:HUR

[Che90d] K. T. Chen. Harnessing university research for competitiveness, industry support. IEEE Spectrum, 27(10):73– 76, October 1990. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Chen:1990:RSP

[Che90e] K. T. Chen. Reversing sagging precollege skills in mathematics and science. IEEE Spectrum, 27(12):44–48, December 1990. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Chen:1990:SME

[Che90f] K. T. Chen. The single-market engineer. IEEE Spectrum, 27 (6):47–49, June 1990. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Chen:1991:TT

[Che91] K. T. Chen. Technology 1991: transportation. IEEE Spectrum, 28(1):69–71, January 1991. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Chen:1999:PJD

[Che99] E. Chen. Poison Java [data security]. IEEE Spectrum, 36 (8):38–43, August 1999. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Chrysler:1990:CMT

[Chr90] M. Chrysler. Canon: more than just concern. IEEE Spectrum, 27(11):113–116, November 1990. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Ciciora:1995:IST

[Cic95] W. S. Ciciora. Inside the settop box. IEEE Spectrum, 32 (4):70–75, April 1995. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cybenko:1992:SRE

[CK92] G. Cybenko and D. J. Kuck. Supercomputers — revolution or evolution? IEEE Spectrum, 29(9):39–41, September 1992. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Clarke:1996:CAV

[CK96] E. M. Clarke and R. P. Kurshan. Computer-aided verification. IEEE Spectrum, 33 (6):61–67, June 1996. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Carpenter:1999:DID

[CK99] B. E. Carpenter and D. D. Kandlur. Diversifying Internet delivery. IEEE Spectrum, 36(11):57–61, November 1999. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Chidambaram:1994:NTP

[CKR94] R. Chidambaram, A. Kakodkar, and P. Rodriguez. Nuclear technology: power to the people. IEEE Spectrum, 31 (3):36–39, March 1994. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Clark:1992:FPL

[Cla92] T. R. Clark. Fitting programmable logic. IEEE Spectrum, 29(3):36–39, March 1992. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Clark:1998:TIT

[Cla98] A. Clark. Too intelligent, or too artificial, or both? [book reviews]. IEEE Spectrum, 35(10):12–14, October 1998. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Caporossi:1996:RCR

[CMR96] D. Caporossi, F. E. Marschner, and S. Read. Rule checking at the register level. IEEE Spectrum, 33(6):72–73, June 1996. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Comerford:1997:VCH

[CMS97] Richard Comerford, J. Makhoul, and R. Schwartz. The voice of the computer is heard in the land (and it listens too!) [speech recognition].

IEEE Spectrum, 34(12):39– 47, December 1997. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cole:1992:DLM

[Col92] B. Cole. DOE labs: models for tech transfer. IEEE Spectrum, 29(12):53–57, December 1992. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cole:1993:MTF

[Col93] B. Cole. Multimedia-the technology framework. IEEE Spectrum, 30(3):32–39, March 1993. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Colmenares:1994:FPW

[Col94] N. J. Colmenares. The FCC on personal wireless. IEEE Spectrum, 31(5):39–46, May 1994. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Comerford:1992:EWA

[Com92a] Richard Comerford. Engineering workstations — add-ons add versatility. IEEE Spectrum, 29(4):46–51, April 1992. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Comerford:1992:HDD

[Com92b] Richard Comerford. How DEC developed Alpha. IEEE Spectrum, 29(7):26–31, July 1992. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Comerford:1992:SSB

[Com92c] Richard Comerford. Supercomputers — software on the brink. IEEE Spectrum, 29(9): 34–38, September 1992. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Comerford:1992:TS

[Com92d] Richard Comerford. Technology 1992 — software. IEEE Spectrum, 29(1):30–32, January 1992. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Comerford:1993:EP

[Com93a] Richard Comerford. Engineers and PCs. IEEE Spectrum, 30 (5):40–42, May 1993. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Comerford:1993:HTR

[Com93b] Richard Comerford. La haute technologie (research initiatives). IEEE Spectrum, 30 (3):55–60, March 1993. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Comerford:1993:PNC

[Com93c] Richard Comerford. Platforms: the new contenders — operating systems go head-to-head. IEEE Spectrum, 30(12):

23–25, December 1993. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Comerford:1993:QLR

[Com93d] Richard Comerford. A quick look at rapid prototyping. IEEE Spectrum, 30(9):28– 29, September 1993. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Comerford:1993:TPW

[Com93e] Richard Comerford. Technology 1993 — PCs and workstations. IEEE Spectrum, 30 (1):26–29, January 1993. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Comerford:1993:TS

[Com93f] Richard Comerford. Technology 1993 — software. IEEE Spectrum, 30(1):30–33, January 1993. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Comerford:1993:XEC

[Com93g] Richard Comerford. X, as in expediency (computer terminals). IEEE Spectrum, 30 (5):60–61, May 1993. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Comerford:1994:EWP

[Com94a]  Richard Comerford.  En-

gineering workstations and PCs. IEEE Spectrum, 31 (4):45, April 1994. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Comerford:1994:MWM

[Com94b] Richard Comerford. Mecha… what? [mechatronics]. IEEE Spectrum, 31(8):46–49, August 1994. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Comerford:1994:MD

[Com94c] Richard Comerford. The multimedia drive. IEEE Spectrum, 31(4):77–83, April 1994. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Comerford:1994:PW

[Com94d] Richard Comerford. PCs and workstations. IEEE Spectrum, 31(1):35–37, January 1994. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Comerford:1995:CFA

[Com95a] Richard Comerford. Coping with a flexible architecture. IEEE Spectrum, 32 (4):27–32, April 1995. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Comerford:1995:CTC

[Com95b] Richard Comerford. Cornell Theory Center. IEEE Spectrum, 32(12):64–??, December 1995. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Comerford:1995:SE

[Com95c] Richard Comerford. Software engineering. IEEE Spectrum, 32(1):62–65, January 1995. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Comerford:1996:WSc

[Com96a] R. Comerford. Web sights. IEEE Spectrum, 33(7):93–94, July 1996. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Comerford:1996:CTD

[Com96b] Richard Comerford. Computers [trends/developments]. IEEE Spectrum, 33(1):42– 45, January 1996. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Comerford:1996:IMI

[Com96c] Richard Comerford. Interactive media: an Internet reality. IEEE Spectrum, 33 (4):29–32, April 1996. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Comerford:1996:OGP

[Com96d] Richard Comerford. It’s only growing pains … The second annual roundtable on the state of the Internet. IEEE Spectrum, 33(9):46–55, September 1996. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Comerford:1996:WSa

[Com96e] Richard Comerford. Web sights. IEEE Spectrum, 33 (1):116, January 1996. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Comerford:1996:WSb

[Com96f] Richard Comerford. Web sights. IEEE Spectrum, 33 (3):65, March 1996. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Comerford:1996:WSd

[Com96g] Richard Comerford. Web sights. IEEE Spectrum, 33 (8):75, August 1996. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Comerford:1996:WSe

[Com96h] Richard Comerford. Web sights. IEEE Spectrum, 33 (10):65, October 1996. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Comerford:1996:WSf

[Com96i] Richard Comerford. Web sights. IEEE Spectrum, 33 (11):77–78, November 1996. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Comerford:1996:WSg

[Com96j] Richard Comerford. Web sights. IEEE Spectrum, 33 (12):24, December 1996. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Comerford:1996:WWG

[Com96k] Richard Comerford. Workstation wars get personal. IEEE Spectrum, 33(6):43–47, June 1996. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Comerford:1997:BDN

[Com97a] Richard Comerford. The battle for the desktop [network computers]. IEEE Spectrum, 34(5):20–28, May 1997. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Comerford:1997:SET

[Com97b] Richard Comerford. Software engineering [technology analysis and forecast]. IEEE Spectrum, 34(1):65–69, January 1997. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Comerford:1997:SIR

[Com97c] Richard Comerford. State of the Internet: roundtable 3.0. IEEE Spectrum, 34(10): 29–38, October 1997. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Comerford:1997:WSa

[Com97d] Richard Comerford. Web sights. IEEE Spectrum, 34 (1):105, January 1997. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Comerford:1997:WSb

[Com97e] Richard Comerford. Web sights. IEEE Spectrum, 34 (3):80, March 1997. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Comerford:1997:WSc

[Com97f] Richard Comerford. Web sights. IEEE Spectrum, 34 (4):65, April 1997. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Comerford:1997:WSd

[Com97g] Richard Comerford. Web sights. IEEE Spectrum, 34 (5):85, May 1997. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Comerford:1997:WSe

[Com97h] Richard Comerford. Web sights. IEEE Spectrum, 34 (6):84, June 1997. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Comerford:1997:WSg

[Com97i] Richard Comerford. Web sights. IEEE Spectrum, 34(9): 107, September 1997. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Comerford:1997:WSh

[Com97j] Richard Comerford. Web sights. IEEE Spectrum, 34 (12):89, December 1997. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Comerford:1997:WSf

[Com97k] Richard Comerford. Web sites. IEEE Spectrum, 34 (7):79, July 1997. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Comerford:1998:CTA

[Com98a] Richard Comerford. Computers [technology 1998 analysis and forecast]. IEEE Spectrum, 35(1):43–47, January 1998. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Comerford:1998:PCI

[Com98b] Richard Comerford. Pocket computers ignite OS battle. IEEE Spectrum, 35(5):43–48, May 1998. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Comerford:1998:WSa

[Com98c] Richard Comerford. Web sights. IEEE Spectrum, 35 (1):125, January 1998. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Comerford:1998:WSb

[Com98d] Richard Comerford. Web sights. IEEE Spectrum, 35 (3):69, March 1998. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Comerford:1998:WSc

[Com98e] Richard Comerford. Web sights. IEEE Spectrum, 35 (4):83, April 1998. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Comerford:1998:WSd

[Com98f] Richard Comerford. Web sights. IEEE Spectrum, 35 (5):71, May 1998. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Comerford:1998:WSe

[Com98g] Richard Comerford. Web sights. IEEE Spectrum, 35 (7):76, July 1998. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Comerford:1998:WSf

[Com98h] Richard Comerford. Web sights. IEEE Spectrum, 35 (8):76, August 1998. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Comerford:1998:WSg

[Com98i] Richard Comerford. Web sights. IEEE Spectrum, 35 (9):80, September 1998. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Comerford:1998:W

[Com98j] Richard Comerford. Websights. IEEE Spectrum, 35 (12):64, December 1998. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Comerford:1999:POS

[Com99a] Richard Comerford. The path to open-source systems. IEEE Spectrum, 36(5):25–31, May 1999. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Comerford:1999:SET

[Com99b] Richard Comerford. Software engineering [technology 1999 analysis and forecast]. IEEE Spectrum, 36(1):84–87, January 1999. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Comerford:1999:SIR

[Com99c] Richard Comerford. State of the Internet: Roundtable 5.0. IEEE Spectrum, 36(10): 42–50, October 1999. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Comerford:1999:WSa

[Com99d] Richard Comerford. Web sights. IEEE Spectrum, 36 (1):131, January 1999. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Comerford:1999:WSb

[Com99e] Richard Comerford. Web sights. IEEE Spectrum, 36 (5):66, May 1999. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Comerford:1999:WSc

[Com99f] Richard Comerford. Web sights. IEEE Spectrum, 36 (8):89, August 1999. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Comerford:1999:W

[Com99g] Richard Comerford. Websights. IEEE Spectrum, 36 (2):77, February 1999. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Comerford:1999:YTD

[Com99h] Richard Comerford. Y2K testing: does success matter? IEEE Spectrum, 36(7):71–73, July 1999. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Coopersmith:1993:FFS

[Coo93] J. Coopersmith. Facsimile’s false starts. IEEE Spectrum, 30(2):46–49, February 1993. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cox:1992:FF

[Cox92] E. Cox. Fuzzy fundamentals. IEEE Spectrum, 29(10): 58–61, October 1992. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cox:1993:AFS

[Cox93] E. Cox. Adaptive fuzzy systems. IEEE Spectrum, 30(2): 27–31, February 1993. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Chandramouli:1996:TSC

[CP96] R. Chandramouli and S. Pateras. Testing systems on a chip. IEEE Spectrum, 33(11): 42–47, November 1996. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Comerford:1998:BYM

[CP98] Richard Comerford and T. Perry.

Brooding on the Year 2000 [Millennium bug]. IEEE Spectrum, 35(6):68–73, June 1998. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cressler:1995:RES

[Cre95] J. D. Cressler. Re-engineering silicon: Si–Ge heterojunction bipolar technology. IEEE Spectrum, 32(3):49–55, March 1995. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cross:1991:JNP

[Cro91] M. Cross. Japan’s nuclear power tightrope. IEEE Spectrum, 28(4):77–80, April 1991. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cruz:1999:CNS

[Cru99] R. L. Cruz. Communications networks simulated [software reviews]. IEEE Spectrum, 36 (8):86, August 1999. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cagan:1993:SPP

[CS93] L. D. Cagan and A. H. Sherman. Software/parallel processing — Linda on the networks. IEEE Spectrum, 30 (12):31–35, December 1993. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Curran:1993:ETT

[Cur93] L. J. Curran. ‘A’ for effort (technology transfer education). IEEE Spectrum, 30 (2):50–52, February 1993. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Collier:1994:SCS

[CW94] W. C. Collier and R. J. Weiland. Smart cars, smart highways. IEEE Spectrum, 31 (4):27–33, April 1994. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Chalamala:1998:FFT

[CWG98] B. R. Chalamala, Yi Wei, and B. E. Gnade. FED up with fat tubes. IEEE Spectrum, 35 (4):42–51, April 1998. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Chen:1990:MC

[CZ90] K. T. Chen and Glenn Zorpette. Microelectronics and computers. IEEE Spectrum, 27(6):32–33, June 1990. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Dai:1992:ESM

[Dai92] Wayne Wei-Ming Dai. Engineering software — multichip routing and placement.

IEEE Spectrum, 29(11):61– 62, November 1992. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Dallas:1990:DPX

[Dal90] W. J. Dallas. A digital prescription for X-ray overload. IEEE Spectrum, 27 (4):33–36, April 1990. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Davis:1994:APR

[Dav94] T. Davis. Adopting a policy of reuse. IEEE Spectrum, 31(6):44–48, June 1994. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Davidson:1997:TLO

[Dav97] S. Davidson. Testing LANs optically to the Gigabit Ethernet standard. IEEE Spectrum, 34(9):86–90, September 1997. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Day:1996:MMP

[Day96] D. A. Day. Misguided missile policy [book review]. IEEE Spectrum, 33(4):11, April 1996. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Day:1997:MS

[Day97] D. A. Day. Manna from space? IEEE Spectrum, 34(4):13, April 1997. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Drummond:1996:PRM

[DdMMNM96] M. Drummond, J. d. M. M. Neto, and M. P. Meirelles. Power R&D: more watts, but less to pay. IEEE Spectrum, 33 (7):60–62, July 1996. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Deyhimy:1995:GAJ

[Dey95] I. Deyhimy. Gallium arsenide joins the giants. IEEE Spectrum, 32(2):33–40, February 1995. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

DiTomaso:1992:DHT

[DF92] N. DiTomaso and G. F. Farris. Diversity in the hightech workplace: diversity and performance in R&D. IEEE Spectrum, 29(6):21–24, June 1992. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

DiTomaso:1994:DDW

[DFC94] N. DiTomaso, G. F. Farris, and R. Cordero. Degrees and diversity at work. IEEE Spectrum, 31(4):38–42, April 1994. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Dipert:1993:FMG

[DH93] B. Dipert and L. Hebert.

Flash memory goes mainstream. IEEE Spectrum, 30(10):48–52, October 1993. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Dillon:1994:SBL

[Dil94] P. S. Dillon. Salvageability becoming law. IEEE Spectrum, 31(8):18–21, August 1994. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Divan:1996:LSS

[Div96] D. Divan. Low stress switching for efficiency. IEEE Spectrum, 33(12):33–39, December 1996. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Dixit:1995:LVD

[Dix95] S. Dixit. A look at the video dial-tone network. IEEE Spectrum, 32(4):64–65, April 1995. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

DeWinkel:1993:SPC

[DL93] C. DeWinkel and J. D. Lamoree. Storing power for critical loads. IEEE Spectrum, 30(6):38–42, June 1993. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Dagnelie:1996:TAE

[DM96] G. Dagnelie and R. W. Massof. Toward an artificial eye. IEEE Spectrum, 33(5):20–29, May 1996. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

daRocha:1996:SOR

[dMA96] A. E. da Rocha, C. K. Morooka, and L. Alegre. Smart oil recovery. IEEE Spectrum, 33(7):48–51, July 1996. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

DAjuz:1996:SSL

[DMP96] A. D’Ajuz, M. L. B. Martinez, and F. A. F. Pazo Blanco. Supplying small loads at reasonable cost. IEEE Spectrum, 33(6):34–35, June 1996. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Doering:1996:SV

[Doe96] E. R. Doering. Statistics visualized. IEEE Spectrum, 33(6):18, June 1996. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Doering:1997:GLD

[Doe97] E. R. Doering. Good-looking data analysis charts. IEEE Spectrum, 34(4):14, April 1997. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Dominach:1994:DRD

[Dom94] R. F. Dominach. Design reviews at a distance. IEEE Spectrum, 31(6):39–40, June 1994. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Dooling:1993:TAM

[Doo93a]  D. Dooling.  Technology

1993 — aerospace and military. IEEE Spectrum, 30(1): 72–75, January 1993. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Dooling:1993:TT

[Doo93b] D. Dooling. Technology 1993 — transportation. IEEE Spectrum, 30(1):68–71, January 1993. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Dooling:1994:AM

[Doo94a] D. Dooling. Aerospace and military. IEEE Spectrum, 31 (1):66–69, January 1994. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Dooling:1994:LQC

[Doo94b] D. Dooling. L+25: a quarter century after the Apollo landing. IEEE Spectrum, 31 (7):16–29, July 1994. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Dooling:1994:NCS

[Doo94c] D. Dooling. Navigating close to shore. IEEE Spectrum, 31(12):24–31, December 1994. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Dooling:1994:T

[Doo94d]  D. Dooling.  Transporta-

tion. IEEE Spectrum, 31(1): 62–65, January 1994. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Dooling:1995:AM

[Doo95a] D. Dooling. Aerospace and military. IEEE Spectrum, 32 (1):76–79, January 1995. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Dooling:1995:ROB

[Doo95b] D. Dooling. Research outpost beyond the sky [International Space Station]. IEEE Spectrum, 32(10):28–33, October 1995. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Dooling:1995:SWS

[Doo95c] D. Dooling. Stormy weather in space. IEEE Spectrum, 32 (6):64–72, June 1995. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Dooling:1995:T

[Doo95d]  D. Dooling.  Transporta-

tion. IEEE Spectrum, 32(1): 71–75, January 1995. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Dooling:1996:AM

[Doo96a] D. Dooling. Aerospace and military. IEEE Spectrum, 33 (1):87–91, January 1996. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Dooling:1996:CEB

[Doo96b] D. Dooling. The continuous electron beam accelerator facility. IEEE Spectrum, 33(12): 55–57, December 1996. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Dooling:1996:LCM

[Doo96c] D. Dooling. Launcher without a country [marine satellite launch platform]. IEEE Spectrum, 33(10):18–25, October 1996. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Dooling:1996:SSS

[Doo96d] D. Dooling. Smoother sailing [ship design and control]. IEEE Spectrum, 33(8): 52–59, August 1996. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Dooling:1996:T

[Doo96e]  D. Dooling.  Transporta-

tion. IEEE Spectrum, 33(1): 82–86, January 1996. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Dooling:1997:AMT

[Doo97a] D. Dooling. Aerospace and military [technology development]. IEEE Spectrum, 34 (1):89–93, January 1997. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Dooling:1997:SS

[Doo97b]  D. Dooling.  Space sen-

tries. IEEE Spectrum, 34(9): 50–59, September 1997. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Dooling:1997:T

[Doo97c]  D. Dooling.  Transporta-

tion. IEEE Spectrum, 34(1): 84–88, January 1997. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Dooling:1998:AMT

[Doo98] D. Dooling. Aerospace and military [technology 1998 analysis and forecast]. IEEE Spectrum, 35(1):90–94, January 1998. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Dozier:1992:SSS

[Doz92] H. Dozier. Supercomputers — superchips for supercomputing. IEEE Spectrum, 29(9): 66–68, September 1992. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Delisle:1991:GDC

[DP91] D. Delisle and L. Pelamougues. A guide to data communications: B-ISDN and how it works. IEEE Spectrum, 28(8):39–42, August 1991. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Damiano:1991:LSA

[DR91] R. Damiano and D. S. Reeves.

Logic synthesis for ASICs. IEEE Spectrum, 28(11):26– 33, November 1991. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Denniston:1995:GNF

[DR95] F. J. Denniston and P. K.

Runge. The glass necklace [FLAG submarine fibre-optic link]. IEEE Spectrum, 32(10): 24–27, October 1995. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Dutta-Roy:1996:TB

[DR96] A. Dutta-Roy. Telecommunications [brazil]. IEEE Spectrum, 33(6):36–42, June 1996.

CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Dutta-Roy:1997:TMG

[DR97] A. Dutta-Roy. Too much of a good thing? IEEE Spectrum, 34(6):8–14, June 1997. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Dutta-Roy:1998:VMD

[DR98] A. Dutta-Roy. Virtual meetings with desktop conferencing. IEEE Spectrum, 35 (7):47–56, July 1998. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Dutta-Roy:1999:BHI

[DR99a] A. Dutta-Roy. Bring home the Internet. IEEE Spectrum, 36 (3):32–38, March 1999. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Dutta-Roy:1999:CJT

[DR99b] A. Dutta-Roy. Cable: it’s not just for TV. IEEE Spectrum, 36(5):53–59, May 1999. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Dutta-Roy:1999:FWR

[DR99c] A. Dutta-Roy. Fixed wireless routes for Internet access. IEEE Spectrum, 36(9): 61–69, September 1999. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Dutta-Roy:1999:NH

[DR99d] A. Dutta-Roy. Networks for homes. IEEE Spectrum, 36 (12):26–33, December 1999. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Dutta-Roy:1999:SWW

[DR99e] A. Dutta-Roy. A second wind for wiring [ADSL]. IEEE Spectrum, 36(9):52–60, September 1999. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Dutta-Roya:1999:CTF

[DR99f] A. Dutta-Roya. Computers [1999 technology forecast and analysis]. IEEE Spectrum, 36 (1):46–51, January 1999. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Dutta-Roy:1997:SR

[DRE97] A. Dutta-Roy and D. Ellis.

Software reviews. IEEE Spectrum, 34(12):66–67, December 1997. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Dutta-Roy:1996:GLI

[DRS96] A. Dutta-Roy and N. M.

Segoshi. Going on line with the Internet. IEEE Spectrum, 33(7):54–59, July 1996. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

DeFanti:1993:RV

[DSCN93] T. A. DeFanti, D. J. Sandin, and C. Cruz-Neira. A ‘room’ with a ‘view’. IEEE Spectrum, 30(10):30–33, October 1993. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Dos-Santos-Dionisio:1996:MPC

[DSD96] F. Dos-Santos-Dionisio. Making privatization constitutional. IEEE Spectrum, 33 (6):38–39, June 1996. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Dye:1992:DSS

[DT92] T. A. Dye and E. Teose. Digital storage scopes advance. IEEE Spectrum, 29(2): 38–41, February 1992. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Doherty:1998:PPS

[DU98] C. Doherty and T. Uslander. Plug and play software hones car and plane manufacture. IEEE Spectrum, 35(10): 39–42, October 1998. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Duane:1995:BE

[Dua95] M. Duane. A broad education. IEEE Spectrum, 32 (12):6, December 1995. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Duarte:1996:MSG

[Dua96] C. H. C. Duarte. Moving software to a global platform. IEEE Spectrum, 33(7):40–43, July 1996. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Eyre:1999:DCC

[EB99] J. Eyre and J. Bier. DSPs court the consumer. IEEE Spectrum, 36(3):47–53, March 1999. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Eden:1997:DIP

[Ede97] M. Eden. Digital image processing [book review]. IEEE Spectrum, 34(3):10–11, March 1997. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Eden:1999:BRS

[Ede99] M. Eden. Book review: The story of electrical and magnetic measurements: from 500 B.C. to the 1940’s. IEEE Spectrum, 36(8):14–15, August 1999. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Edwards:1997:BPM

[Edw97] P. N. Edwards. From baggage to the PC, minus the hype. IEEE Spectrum, 34(2):

10–12, February 1997. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Ehmann:1993:MPM

[EJ93] K. F. Ehmann and P. C. Jones. Master’s programs and more (manufacturing education). IEEE Spectrum, 30(9): 76–81, September 1993. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Eldridge:1993:FF

[Eld93] J. Eldridge. Filing in a flash. IEEE Spectrum, 30(10): 53–54, October 1993. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Elliott:1996:OCQ

[Ell96] D. S. Elliott. Optical coherence and quantum optics [book review]. IEEE Spectrum, 33(9):12–13, September 1996. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Elliott:1999:NAN

[Ell99a] S. J. Elliott. Down with noise [active noise control]. IEEE Spectrum, 36(6):54–61, June 1999. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Ellis:1999:ADC

[Ell99b] D. Ellis. Adaptable 3-D clipart adorns documents [software reviews]. IEEE Spectrum, 36(8):84, August 1999. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Ellis:1999:WVP

[Ell99c] D. Ellis. The well-versed programmer. IEEE Spectrum, 36 (2):12–13, February 1999. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Engler:1996:FF

[Eng96] C. E. Engler. Faults failures.

IEEE Spectrum, 33(8):60, August 1996. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Engler:1997:BRH

[Eng97a] C. Engler. Book review: How to Set Up and Maintain a World Wide Web Site. IEEE Spectrum, 34(3):11, March 1997. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Engler:1997:WS

[Eng97b] C. E. Engler. Web sights.

IEEE Spectrum, 34(8):81– 82, August 1997. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Engler:1997:WR

[Eng97c] C. E. Engler. Webware reviews. IEEE Spectrum, 34(1): 104T2–104T7, January 1997. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Engler:1998:FF

[Eng98] C. E. Engler. Faults failures. IEEE Spectrum, 35 (4):72, April 1998. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Eurich:1990:EG

[ER90] J. P. Eurich and G. Roth. EDIF grows up. IEEE Spectrum, 27(11):68–72, November 1990. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Esch:1990:AD

[Esc90] K. Esch. Aerospace and defense. IEEE Spectrum, 27 (6):30–32, June 1990. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Eubanks:1998:SRS

[Eub98] B. Eubanks. Software reviews suite of tools helps beget x86 applications. IEEE Spectrum, 35(10):17, October 1998. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Even:1998:FVA

[Eve98] S. Even. Four value-adding algorithms. IEEE Spectrum, 35(5):33–38, May 1998. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Ellis:1997:SR

[EW97] D. Ellis and C. F. Weggel.

Software reviews. IEEE Spectrum, 34(8):91–93, August 1997. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Falcao:1996:PCC

[Fal96] D. M. Falcao. Powerful oncampus computing for industry. IEEE Spectrum, 33(6):32, June 1996. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Fancher:1997:YPS

[Fan97] C. H. Fancher. In your pocket: smartcards. IEEE Spectrum, 34(2):47–53, February 1997. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Fisher:1996:PDB

[FASN96] G. D. Fisher, W. L. Abbott, J. L. Sonntag, and R. Nesin. PRML detection boosts hard-disk drive capacity. IEEE Spectrum, 33(11): 70–76, November 1996. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Fawcett:1994:TSF

[Faw94] B. K. Fawcett. Tools to speed FPGA development. IEEE Spectrum, 31(11):88– 94, November 1994. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Fayyazuddin:1999:BRG

[Fay99] A. Fayyazuddin. Book review: Gravity, particles, and the Planck length. IEEE Spectrum, 36(6):14–16, June 1999. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Feffer:1999:IIL

[Fef99] L. B. Feffer. Insisting on the impossible: the life of Edwin Land. IEEE Spectrum, 36(5):14–16, May 1999. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Felix:1997:SNE

[Fel97] F. Felix. State of the nuclear economy. IEEE Spectrum, 34(11):29–32, November 1997. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Ferguson:1991:SMC

[Fer91] W. E. Ferguson. Selecting math coprocessors. IEEE Spectrum, 28(7):38–41, July 1991. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Ferrone:1998:ETA

[Fer98] R. Ferrone. The environment [technology 1998 analysis and forecast]. IEEE Spectrum, 35 (1):95–99, January 1998. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Fielder:1997:BRN

[Fie97] John Fielder. Book review:

Naked to the Bone: Medical Imaging in the Twentieth Century, Kevles, Bettyann Holtzman, Rutgers University Press, New Brunswick, NJ, 378 pp., $35.95. IEEE Spectrum, 34(12):14–15, December 1997. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Filho:1996:USB

[Fil96] A. V. Filho. Utilities for sale [brazil]. IEEE Spectrum, 33 (6):29–33, June 1996. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Finn:1990:CBW

[Fin90] B. Finn. A continent bound by wire (telegraph system). IEEE Spectrum, 27(8):58–59, August 1990. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Fineman:1995:TA

[Fin95] J. J. Fineman. Technology for all. IEEE Spectrum, 32(12): 8–??, December 1995. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Finvers:1997:DMG

[Fin97] I. Finvers. Diagramming minus the gruntwork. IEEE Spectrum, 34(6):65, June 1997. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Fisher:1992:IPR

[Fis92] P. D. Fisher. Improving on police radar. IEEE Spectrum, 29(7):38–43, July 1992. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Fischetti:1993:CPS

[Fis93] M. Fischetti. The cellular phone scare. IEEE Spectrum, 30(6):43–47, June 1993. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Fitzgerald:1990:AI

[Fit90a] K. Fitzgerald. Applications ’90: instrumentation. IEEE Spectrum, 27(2):37–38, February 1990. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Fitzgerald:1990:CRG

[Fit90b] K. Fitzgerald. Contract research gains adherents for its flexibility and low overhead. IEEE Spectrum, 27(10): 58–60, October 1990. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Fitzgerald:1990:EFJ

[Fit90c] K. Fitzgerald. Electromagnetic fields: the jury’s still out. 2. Societal reverberations. IEEE Spectrum, 27 (8):27–32, August 1990. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Fitzgerald:1990:ERT

[Fit90d] K. Fitzgerald. Encouraging risk-taking, sanctioning failures are helping spur creativity. IEEE Spectrum, 27(10): 67–69, October 1990. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Fitzgerald:1990:GS

[Fit90e] K. Fitzgerald. Global standards. IEEE Spectrum, 27 (6):44–46, June 1990. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Fitzgerald:1990:SR

[Fit90f] K. Fitzgerald. Shoreham in repose. IEEE Spectrum, 27 (5):46–48, May 1990. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Fitzgerald:1990:TME

[Fit90g] K. Fitzgerald. Technology ’90: medical electronics. IEEE Spectrum, 27(1):52–54, January 1990. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Fitzgerald:1990:WBA

[Fit90h]  K. Fitzgerald.  Whistle-

blowing: not always a losing game (engineers). IEEE Spectrum, 27(12):49–52, December 1990. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Fitzgerald:1991:EMOb

[Fit91a] K. Fitzgerald. Energy management — origin of a culture. IEEE Spectrum, 28 (3):45–47, March 1991. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Fitzgerald:1991:EMOa

[Fit91b]  K. Fitzgerald.  Energy

management — overhauling weapons production. IEEE Spectrum, 28(3):39–44, March 1991. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Fitzgerald:1991:TME

[Fit91c]  K. Fitzgerald.  Technol-

ogy 1991: medical electronics. IEEE Spectrum, 28(1): 76–78, January 1991. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Fitzgerald:1991:TTM

[Fit91d] K. Fitzgerald. Technology 1991: test and measurement. IEEE Spectrum, 28(1): 56–58, January 1991. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Flores:1991:OLS

[FK91] G. E. Flores and B. Kirkpatrick. Optical lithography stalls X-rays. IEEE Spectrum, 28(10):24–27, October 1991. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Flint:1994:TMM

[Fli94] A. Flint. Testing multichip modules. IEEE Spectrum, 31 (3):59–62, March 1994. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Flory:1990:GBB

[Flo90] D. Flory. The great blue box phone frauds. IEEE Spectrum, 27(11):117–119, November 1990. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Farmwald:1992:MFP

[FM92] M. P. Farmwald and D. Mooring. Memory — a fast path to one memory. IEEE Spectrum, 29(10):50–51, October 1992. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Focel:1999:TVT

[Foc99] D. B. Focel. A true visionary’s truly fantastic vision. IEEE Spectrum, 36(7):9–13, July 1999. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Forden:1997:AL

[For97] G. E. Forden. The airborne laser. IEEE Spectrum, 34(9): 40–49, September 1997. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Foster:1991:PM

[Fos91] Kenneth R. Foster. Prepackaged math. IEEE Spectrum, 28(11):44–50, November 1991. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Foster:1992:ESM

[Fos92] Kenneth R. Foster. Engineering software — math and graphics. IEEE Spectrum, 29 (11):72, November 1992. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Foster:1993:ESA

[Fos93]  Kenneth R. Foster.  En-

gineering software — ‘abstract’ math made practical. IEEE Spectrum, 30(11):42– 46, November 1993. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Foster:1995:PE

[Fos95] Kenneth R. Foster. Polished electromagnetics. IEEE Spectrum, 32(12):58–59, December 1995. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Foster:1996:HSE

[Fos96a] Kenneth R. Foster. Handbook of software for engineers and scientists [book review]. IEEE Spectrum, 33(9): 10–12, September 1996. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Foster:1996:MMU

[Fos96b] Kenneth R. Foster. Major math updates [software reviews]. IEEE Spectrum, 33 (8):64–66, August 1996. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Foster:1996:STL

[Fos96c] Kenneth R. Foster. Solver tackles large sets of equations [software reviews]. IEEE Spectrum, 33(12):22, December 1996. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Foster:1996:VI

[Fos96d] Kenneth R. Foster. Virtual instruments. IEEE Spectrum, 33(4):18, April 1996. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Foster:1997:HPT

[Fos97a] Kenneth R. Foster. HiQ, and pretty, too. IEEE Spectrum, 34(6):65–66, June 1997. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Foster:1997:IVS

[Fos97b]  Kenneth R. Foster.  Im-

proved visualization [software reviews]. IEEE Spectrum, 34 (2):14–15, February 1997. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Foster:1997:MMM

[Fos97c] Kenneth R. Foster. Matrices and much, much more [software reviews]. IEEE Spectrum, 34(2):14, February 1997. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Foster:1997:NML

[Fos97d] Kenneth R. Foster. New math for labview. IEEE Spectrum, 34(6):66, June 1997. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Foster:1997:SR

[Fos97e] Kenneth R. Foster. Software reviews. IEEE Spectrum, 34(10):68–69, October 1997. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Foster:1997:TIB

[Fos97f] Kenneth R. Foster. Trying to have it both ways [review of Beginnings Count: The Technological Imperative in American Health Care (Rothman, D.). IEEE Spectrum, 34(12):

12–14, December 1997. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Foster:1998:EHF

[Fos98a] K. Foster. Engineering’s human face. IEEE Spectrum, 35(10):14–15, October 1998. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Foster:1998:SR

[Fos98b] Kenneth R. Foster. Software reviews. IEEE Spectrum, 35 (12):92, December 1998. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Foster:1998:STT

[Fos98c] Kenneth R. Foster. Software tools [technology 1998 analysis and forecast]. IEEE Spectrum, 35(1):52–56, January 1998. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Foster:1998:SWS

[Fos98d] Kenneth R. Foster. Software wedge for serial data [software reviews]. IEEE Spectrum, 35 (6):91–92, June 1998. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Foster:1999:GLW

[Fos99a] Kenneth R. Foster. The golem at large: what you should know about technology [book reviews]. IEEE Spectrum, 36 (4):16–18, April 1999. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Foster:1999:MI

[Fos99b] Kenneth R. Foster. Math on the Internet. IEEE Spectrum, 36(4):36–40, April 1999. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Foster:1999:SR

[Fos99c] Kenneth R. Foster. Software reviews. IEEE Spectrum, 36(12):85–88, December 1999. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Fowler:1993:DYE

[Fow93] C. A. Fowler. DOs and DON’Ts for young EEs. IEEE Spectrum, 30(10):59–61, October 1993. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Fowler:1994:DAG

[Fow94] C. A. Fowler. Defense acquisition: grab the ax. IEEE Spectrum, 31(10):55–59, October 1994. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Fox:1993:HLS

[Fox93] J. R. Fox. A higher level of synthesis (CAD). IEEE Spectrum, 30(3):43–47, March 1993. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Fox:1995:DDE

[Fox95] B. Fox. The digital dawn in Europe [HDTV]. IEEE Spectrum, 32(4):50–53, April 1995. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Fox:1998:DTC

[Fox98] B. Fox. Digital TV comes down to Earth. IEEE Spectrum, 35(10):23–29, October 1998. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Fox:1999:SRI

[Fox99] P. Fox. Software reviews — IDL data visualization broadly upgraded. IEEE Spectrum, 36(6):93–94, June 1999. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Forbes:1993:SFC

[FP93] C. A. Forbes and J. F.

Pierre. The solid fuel-cell future. IEEE Spectrum, 30(10): 40–44, October 1993. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Foss:1992:MFI

[FPR^{+}92] R. C. Foss, B. Prince, R. Rodgers, D. B. Gustavson, D. V. James, G. Stone, and S. Kempainen. Memory — fast interfaces for DRAMs.

IEEE Spectrum, 29(10):54– 57, October 1992. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Frankel:1995:IRM

[Fra95a] D. Frankel. ISDN reaches the market. IEEE Spectrum, 32 (6):20–25, June 1995. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Fray:1995:CSF

[Fra95b] R. Fray. Compact simulators for fossil-fueled power plants. IEEE Spectrum, 32(2): 46–51, February 1995. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Frankel:1996:EEI

[Fra96] C. Frankel. The environment [electronics industry aspects]. IEEE Spectrum, 33 (1):76–81, January 1996. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Frenkil:1998:PEM

[Fre98] J. Frenkil. The practical engineer [a multi-level approach to low-power IC design]. IEEE Spectrum, 35(2):54–60, February 1998. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Friedman:1996:RKG

[Fri96] M. B. Friedman. RAID keeps going and going and … [magnetic disk storage]. IEEE Spectrum, 33(4):73–79, April 1996. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Furtney:1993:WS

[FT93] M. Furtney and G. Taylor.

Of workstations and supercomputers. IEEE Spectrum, 30(5):64–68, May 1993. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Fujita:1996:DCC

[Fuj96] M. Fujita. Debugging a communications chip. IEEE Spectrum, 33(6):64, June 1996. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Fuqua:1996:VGA

[Fuq96] D. Fuqua. Viewpoint: Getting ahead of the pack — and staying ahead. IEEE Spectrum, 33(1):90, January 1996. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Gabel:1994:AS

[Gab94a] D. A. Gabel. Application software. IEEE Spectrum, 31(1): 42–45, January 1994. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Gabel:1994:SE

[Gab94b] D. A. Gabel. Software engineering. IEEE Spectrum, 31 (1):38–41, January 1994. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Gaddy:1998:YTB

[Gad98] C. D. Gaddy. To yourself be true (book review). IEEE Spectrum, 35(1):8–11, January 1998. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Gallagher:1994:VLR

[Gal94] R. S. Gallagher. Visualization: the look of reality. IEEE Spectrum, 31(11):48–55, November 1994. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Galletti:1996:VLD

[Gal96] P. M. Galletti. Viewpoint: Let’s do tissue engineering right. IEEE Spectrum, 33 (1):94, January 1996. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Giboney:1998:ILS

[GAL98] K. S. Giboney, L. B. Aronson, and B. E. Lemoff. The ideal light source for datanets. IEEE Spectrum, 35(2):43– 53, February 1998. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Gerla:1991:GDC

[GB91] M. Gerla and J. A. Bannister. A guide to data communications: high-speed local-area networks. IEEE Spectrum, 28 (8):26–31, August 1991. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Gracon:1994:LSC

[GC94] H. Gracon and M. Clark. Layoffs should come last. IEEE Spectrum, 31(5):52–55, May 1994. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Gershon:1995:VNT

[GE95] N. Gershon and S. G. Eick. Visualization’s new tack: making sense of information. IEEE Spectrum, 32(11):38– 40, November 1995. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Gellings:1996:ECP

[Gel96] C. W. Gellings. Electrotechnology cleans up [pollution control measures]. IEEE Spectrum, 33(2):46–50, February 1996. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Gemmell:1997:TC

[Gem97] P. S. Gemmell. Traceable ecash. IEEE Spectrum, 34(2): 35–37, February 1997. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Geppert:1993:PNC

[Gep93] L. Geppert. Platforms: the new contenders — not your father’s CPU. IEEE Spectrum, 30(12):20–23, December 1993. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Geppert:1994:IRN

[Gep94a] L. Geppert. Industrial R&D: the new priorities. IEEE Spectrum, 31(9):30–36, September 1994. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Geppert:1994:SS

[Gep94b] L. Geppert. Solid state. IEEE Spectrum, 31(1):50–53, January 1994. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Geppert:1995:CC

[Gep95a] L. Geppert. Chipmaking in China. IEEE Spectrum, 32 (12):36–??, December 1995. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Geppert:1995:ERE

[Gep95b]  L. Geppert.  Educating

the renaissance engineer. IEEE Spectrum, 32(9):39– 43, September 1995. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Geppert:1995:SST

[Gep95c] L. Geppert. Solid state [1995 technological forecast]. IEEE Spectrum, 32(1):35–39, January 1995. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Geppert:1995:USN

[Gep95d] L. Geppert. The uphill struggle: no rose garden for women in engineering. IEEE Spectrum, 32(5):40–50, May 1995. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Geppert:1996:EWE

[Gep96a] L. Geppert. Entrepreneurial women engineers. IEEE Spectrum, 33(8):28–37, August 1996. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Geppert:1996:FDG

[Gep96b] L. Geppert. Fresh dates are giving computers indigestion [faults and failures]. IEEE Spectrum, 33(5):75, May 1996. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Geppert:1996:SLN

[Gep96c] L. Geppert. Semiconductor lithography for the next millennium. IEEE Spectrum, 33 (4):33–38, April 1996. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Geppert:1996:SST

[Gep96d]  L. Geppert.  Solid state

[trends/developments]. IEEE Spectrum, 33(1):51–55, January 1996. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Geppert:1997:FFa

[Gep97a]  L. Geppert.  Faults fail-

ures. IEEE Spectrum, 34 (2):85, February 1997. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Geppert:1997:FFb

[Gep97b]  L. Geppert.  Faults fail-

ures. IEEE Spectrum, 34 (5):17, May 1997. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Geppert:1997:Ia

[Gep97c]  L. Geppert.  Innovations.

IEEE Spectrum, 34(6):76, June 1997. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Geppert:1997:Ib

[Gep97d]  L. Geppert.  Innovations.

IEEE Spectrum, 34(8):16, August 1997. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Geppert:1997:Ic

[Gep97e]  L. Geppert.  Innovations.

IEEE Spectrum, 34(12):65, December 1997. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Geppert:1997:SSD

[Gep97f] L. Geppert. Solid state [development forecast]. IEEE Spectrum, 34(1):55–59, January 1997. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Geppert:1997:TTW

[Gep97g] L. Geppert. Trials & triumphs: Women engineers in Europe. IEEE Spectrum, 34 (8):44–55, August 1997. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Geppert:1998:ESS

[Gep98a] L. Geppert. Europe’s silicon star rises. IEEE Spectrum, 35(10):34–38, October 1998. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Geppert:1998:HFD

[Gep98b] L. Geppert. High-flying DSP architectures. IEEE Spectrum, 35(11):53–56, November 1998. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Geppert:1998:IDW

[Gep98c] L. Geppert. IC design on the World Wide Web. IEEE Spectrum, 35(6):45–50, June 1998. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Geppert:1998:Ia

[Gep98d]  L. Geppert.  Innovations.

IEEE Spectrum, 35(3):24, March 1998. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Geppert:1998:Ib

[Gep98e]  L. Geppert.  Innovations.

IEEE Spectrum, 35(5):72, May 1998. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Geppert:1998:Ic

[Gep98f]  L. Geppert.  Innovations.

IEEE Spectrum, 35(9):18, September 1998. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Geppert:1998:SST

[Gep98g] L. Geppert. Solid state [technology 1998 analysis and forecast]. IEEE Spectrum, 35(1): 23–28, January 1998. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Geppert:1999:CPS

[Gep99a] L. Geppert. Careers perspectives of six women engineers. IEEE Spectrum, 36(12): 69–77, December 1999. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Geppert:1999:CSB

[Gep99b] L. Geppert. Champagne supercomputer on a beer budget. IEEE Spectrum, 36(5):19–22, May 1999. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Geppert:1999:DTA

[Gep99c] L. Geppert. Design tools for analog and digital ICs. IEEE Spectrum, 36(4):41–48, April 1999. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Geppert:1999:FFa

[Gep99d]  L. Geppert.  Faults fail-

ures. IEEE Spectrum, 36 (3):76–77, March 1999. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Geppert:1999:FFb

[Gep99e]  L. Geppert.  Faults fail-

ures. IEEE Spectrum, 36 (6):92, June 1999. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Geppert:1999:I

[Gep99f]  L. Geppert.  Innovations.

IEEE Spectrum, 36(8):79, August 1999. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Geppert:1999:SSS

[Gep99g] L. Geppert. Solid state [semiconductors. 1999 technology analysis and forecast]. IEEE Spectrum, 36(1):52–56, January 1999. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Getting:1993:PNG

[Get93] I. A. Getting. Perspective/ navigation — the Global Positioning System. IEEE Spectrum, 30(12):36–38, December 1993. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Ginsberg:1998:WCI

[GF98] R. B. Ginsberg and Kenneth R. Foster. The wired classroom [Internet-based education]. IEEE Spectrum, 35 (8):44–51, August 1998. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Gjessing:1992:MRL

[GGJ^{+}92] S. Gjessing, D. B. Gustavson, D. V. James, G. Stone, and H. Wiggers. Memory — a RAM link for high speed. IEEE Spectrum, 29(10):52– 53, October 1992. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Gilmour:1991:HST

[Gil91] P. S. Gilmour. How to select tools for microcontroller software. IEEE Spectrum, 28(2): 37–39, February 1991. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Gilchrist:1998:FCF

[Gil98] T. Gilchrist. Fuel cells to the fore [electric vehicles]. IEEE Spectrum, 35(11):35– 40, November 1998. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Gourley:1994:DM

[GLS94] P. L. Gourley, K. L. Lear, and R. P. Schneider. A different mirror. IEEE Spectrum, 31 (8):31–33, August 1994. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Gershon:1993:EDD

[GM93] N. D. Gershon and C. G.

Miller. Environment — dealing with the data deluge. IEEE Spectrum, 30(7):28–32, July 1993. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Gilliland:1992:CDE

[GMKJ92] M. W. Gilliland, P. Mac-Corquodale, J. P. Kash, and A. Jameton. Conversion and the defense engineer. IEEE Spectrum, 29(12):26– 28, December 1992. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Gal-Oz:1998:ABB

[GOI98] S. Gal-Oz and M. V. Isaacs. Automate the big bottleneck in embedded system design. IEEE Spectrum, 35(8): 62–67, August 1998. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Gomes:1996:TRF

[Gom96] L. E. C. Gomes. Telecom R&D focuses on immediate needs. IEEE Spectrum, 33(7):63–65, July 1996. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Gunshor:1993:BLH

[GON93] R. L. Gunshor, N. Otsuka, and A. V. Nurmiko. Blue lasers on the horizon. IEEE Spectrum, 30(5):28–33, May 1993. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Goyal:1994:MSI

[Goy94] R. Goyal. Managing signal integrity [PCB design]. IEEE Spectrum, 31(3):54–58, March 1994. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Grudinin:1997:HEL

[GR97] N. Grudinin and I. Roytelman. Heading off emergencies in large electric grids. IEEE Spectrum, 34(4):42–47, April 1997. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Gray:1998:VF

[Gra98] S. Gray. In virtual fashion. IEEE Spectrum, 35(2): 18–25, February 1998. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Gray:1999:PEF

[Gra99] K. Gray. The practical engineer — fine-tuning memory macros using variable internal delays. IEEE Spectrum, 36 (8):44–49, August 1999. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Green:1994:CUR

[Gre94] M. Green. Complying with union rules. IEEE Spectrum, 31(4):34–37, April 1994. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Grossman:1992:SMR

[Gro92] M. Grossman. Supercomputers — modeling reality. IEEE Spectrum, 29(9):56– 60, September 1992. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Grossman:1993:TTM

[Gro93] M. Grossman. Technology 1993 — test and measurement. IEEE Spectrum, 30(1): 50–57, January 1993. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Grothenhuis:1998:CER

[Gro98] K. Grothenhuis. Crossing the euro rubicon [information systems]. IEEE Spectrum, 35(10):30–33, October 1998. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Geppert:1997:HIC

[GS97] L. Geppert and W. Sweet. http://: it has changed everything, especially our engineering thinking. IEEE Spectrum, 34(1):23–26, January 1997. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Geppert:1998:BWL

[GS98] L. Geppert and W. Sweet. Breakthroughs will leave their mark on many key technologies. IEEE Spectrum, 35(1): 19–22, January 1998. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Geppert:1999:TAF

[GS99] L. Geppert and W. Sweet. Technology 1999 — analysis forcast. IEEE Spectrum, 36 (1):25–28, January 1999. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Gumas:1999:DWC

[Gum99] C. C. Gumas. The DSP workshop: a compelling case [software reviews]. IEEE Spectrum, 36(8):80–81, August 1999. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Gupta:1994:AGG

[Gup94] M. S. Gupta. In academe’s grungy groves [engineering education]. IEEE Spectrum, 31 (3):16, March 1994. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Guterl:1990:SMM

[Gut90a] F. Guterl. Siemens: making marketing as important as technology. IEEE Spectrum, 27(12):57–60, December 1990. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Guterl:1990:TTV

[Gut90b] F. Guterl. Tradeoffs: turf vs. togetherness (ESPRIT). IEEE Spectrum, 27(6):36–38, June 1990. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Guterl:1991:EBC

[Gut91] F. Guterl. Ericsson bets on a cellular world. IEEE Spectrum, 28(2):48–51, February 1991. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Guterl:1992:NRS

[Gut92] F. Guterl. Nokia’s rising star flickers (manufacturing firm). IEEE Spectrum, 29(5):48–51, May 1992. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Gopel:1998:DS

[GW98] W. Gopel and T. Weiss. Design for smelling. IEEE Spectrum, 35(9):32–34, September 1998. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Hackwood:1993:EPH

[Hac93] S. Hackwood. Extension programs help small US firms. IEEE Spectrum, 30(9):73– 75, September 1993. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Hammerstrom:1993:NNW

[Ham93a] D. Hammerstrom. Neural networks at work. IEEE Spectrum, 30(6):26–32, June 1993. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Hammerstrom:1993:WNN

[Ham93b] D. Hammerstrom. Working with neural networks. IEEE Spectrum, 30(7):46–53, July 1993. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Hamalainen:1995:VRG

[Ham95] J. Hamalainen. Video recording goes digital. IEEE Spectrum, 32(4):76–79, April 1995. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Hancock:1993:PHF

[Han93] D. Hancock. ‘Prototyping’ the Hubble fix. IEEE Spectrum, 30(10):34–39, October 1993. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Hantman:1996:WLI

[Han96] N. T. Hantman. And the winners are … The IEEE Spectrum Quilt Block Contest. IEEE Spectrum, 33 (10):57, October 1996. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Hantman:1997:WLS

[Han97] N. T. Hantman. And the winners are … Spectrum’s Second Quilt Block Contest. IEEE Spectrum, 34(12):48– 51, December 1997. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Hantman:1998:DAQ

[Han98a] N. Hantman. Digital aids to quilting design. IEEE Spectrum, 35(7):13, July 1998. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Hantman:1998:WLS

[Han98b] N. T. Hantman. And the winners are … Spectrum’s Third Quilt Block Contest.

IEEE Spectrum, 35(12):60– 63, December 1998. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Harjani:1992:ESD

[Har92a]  R. Harjani.  Engineering

software — designing mixedsignal ICs. IEEE Spectrum, 29(11):49–51, November 1992. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Hartman:1992:UDC

[Har92b] D. Hartman. Unclogging distributed computing. IEEE Spectrum, 29(5):36–39, May 1992. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Harned:1999:ULI

[Har99a] N. Harned. Ultralight lithography [IC manufacture]. IEEE Spectrum, 36(7):35–40, July 1999. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Harriott:1999:NRB

[Har99b] L. R. Harriott. A new role for e-beam: electron projection. IEEE Spectrum, 36(7):41–45, July 1999. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Haus:1993:MLS

[Hau93] H. A. Haus. Molding light into solitons. IEEE Spectrum, 30 (3):48–53, March 1993. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Hayashi:1993:IM

[Hay93] H. Hayashi. Intelligent manufacturing. IEEE Spectrum, 30(9):82–84, September 1993. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Hoffman:1990:SWO

[HBY90] S. Hoffman and G. Ben-Yaacov. Software workstations: one-stop shopping for utilities. IEEE Spectrum, 27 (12):53–56, December 1990. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Heyns:1996:VHC

[HD96] M. Heyns and G. DeClerk.

Viewpoint: How clean is ‘clean enough’? IEEE Spectrum, 33(1):52–53, January 1996. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Headrick:1990:LHH

[Hea90] J. M. Headrick. Looking over the horizon (hf radar). IEEE Spectrum, 27(7):36–39, July 1990. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Heilmeier:1997:GHH

[Hei97] G. H. Heilmeier. George H. Heilmeier — the recipient of the 1997 IEEE Medal of Honor shares thoughts about research management in this Q&A with Spectrum magazine. IEEE Spectrum, 34 (6):18–23, June 1997. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Hellemans:1996:BS

[Hel96] A. Hellemans. Basic science. IEEE Spectrum, 33(1): 101–104, January 1996. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Hellemans:1997:BST

[Hel97] A. Hellemans. Basic science [1996 technology review]. IEEE Spectrum, 34(1): 70–73, January 1997. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Helemans:1998:BST

[Hel98] A. Helemans. Basic science [technology 1998 analysis and forecast]. IEEE Spectrum, 35(1):100–103, January 1998. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Hellestrand:1999:RSE

[Hel99] G. R. Hellestrand. The revolution in systems engineering. IEEE Spectrum, 36(9): 43–51, September 1999. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Hendriks:1997:SCD

[Hen97] P. Hendriks. Specifying communications DACs. IEEE Spectrum, 34(7):58–69, July 1997. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Herrell:1999:PP

[Her99] D. Herrell. Power to the package. IEEE Spectrum, 36 (7):46–53, July 1999. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Horning:1991:GEW

[HFYT91] R. J. Horning, M. Forsyth, J. Yetter, and L. J. Thayer. A guide to engineering workstations: how ICs impact workstations. IEEE Spectrum, 28 (4):58–62, April 1991. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Hammond:1993:DS

[HHSM93] R. B. Hammond, G. L. Hey-Shipton, and G. L. Matthaei. Designing with superconductors. IEEE Spectrum, 30 (4):34–39, April 1993. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Hess:1992:AQL

[HI92] K. Hess and G. J. Iafrate. Approaching the quantum limit.

IEEE Spectrum, 29(7):44–49, July 1992. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Hills:1991:TIN

[Hil91] A. Hills. Teaching information networking. IEEE Spectrum, 28(10):22–23, October 1991. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Hills:1999:WAM

[Hil99] A. Hills. Wireless Andrew [mobile computing for university campus]. IEEE Spectrum, 36(6):49–53, June 1999. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Hines:1990:AAD

[Hin90] J. R. Hines. Affordable analog design. IEEE Spectrum, 27(11):60–67, November 1990. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Hinton:1992:SP

[Hin92] H. S. Hinton. Switching to photonics. IEEE Spectrum, 29 (2):42–45, February 1992. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Hingorani:1993:FAT

[Hin93] N. G. Hingorani. Flexible AC transmission. IEEE Spectrum, 30(4):40–45, April 1993. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Hingorani:1995:ICP

[Hin95] N. G. Hingorani. Introducing custom power. IEEE Spectrum, 32(6):41–48, June 1995. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Hines:1996:PNa

[Hin96a]  J. R. Hines.  Program

notes. IEEE Spectrum, 33 (1):122, January 1996. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Hines:1996:PNb

[Hin96b] J. R. Hines. Program notes.

IEEE Spectrum, 33(5):81–82, May 1996. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Hines:1996:PNc

[Hin96c] J. R. Hines. Program notes.

IEEE Spectrum, 33(7):100, July 1996. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Hines:1996:PNd

[Hin96d]  J. R. Hines.  Program

notes. IEEE Spectrum, 33 (9):23, September 1996. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Hines:1996:PNe

[Hin96e]  J. R. Hines.  Program

notes. IEEE Spectrum, 33 (11):23, November 1996. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Hines:1996:PNJ

[Hin96f] J. R. Hines. Program notes: Java: jive?; key klatch; good reading. IEEE Spectrum, 33 (3):15, March 1996. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Hines:1996:SE

[Hin96g] J. R. Hines. Software engineering. IEEE Spectrum, 33 (1):60–64, January 1996. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Hingorani:1996:HVD

[Hin96h] N. G. Hingorani. High-voltage DC transmission: a power electronics workhorse. IEEE Spectrum, 33(4):63–72, April 1996. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Hines:1997:PNa

[Hin97a] J. R. Hines. Program notes. IEEE Spectrum, 34(9):20– 22, September 1997. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Hines:1997:PNb

[Hin97b]  J. R. Hines.  Program

notes. IEEE Spectrum, 34 (11):20, November 1997. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Hines:1997:VMJ

[Hin97c] J. R. Hines. Virtual machines jockey for position. IEEE Spectrum, 34(7):16, July 1997. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Hines:1998:SET

[Hin98] J. Hines. Software engineering [technology 1998 analysis and forecast]. IEEE Spectrum, 35 (1):48–51, January 1998. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Hinman:1999:ILH

[Hin99] R. T. Hinman. Infinite loop:

How Apple, the world’s most insanely great computer went insane. IEEE Spectrum, 36 (5):16–17, May 1999. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Hunt:1993:KCB

[HJ93] E. R. Hunt and G. Johnson. Keeping chaos at bay. IEEE Spectrum, 30(11):32– 36, November 1993. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Hammond:1999:MSC

[HK99] J. S. Hammond and R. L. Keeney. Making smart choices in engineering. IEEE Spectrum, 36(11):71–76, November 1999. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Herbert:1999:EIP

[HL99] J. R. Herbert and M. Loui. The ethics of intelligectual property and the new information technologies. IEEE Spectrum, 36(8):29–37, August 1999. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Howe:1990:SMS

[HMGT90] R. T. Howe, R. S. Muller, K. J. Gabriel, and W. S. N. Trimmer. Silicon micromechanics: sensors and actuators on a chip. IEEE Spectrum, 27(7):29–31, July 1990. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Holliday:1997:RG

[Hol97] C. R. Holliday. The residential gateway. IEEE Spectrum, 34(5):29–31, May 1997. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Horgan:1992:CSP

[Hor92] J. Horgan. Claude E. Shannon [profile]. IEEE Spectrum, 29(4):72–75, April 1992. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Hoschette:1994:LGW

[Hos94a]  J. A. Hoschette.  Look-

ing good — whatever happens [engineering profession]. IEEE Spectrum, 31(12):42– 44, December 1994. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Hoschette:1994:TCY

[Hos94b]  J. A. Hoschette.  Taking

charge of your promotion. IEEE Spectrum, 31(11):96– 99, November 1994. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Hoschette:1995:HSS

[Hos95a] J. A. Hoschette. The inhouse springboard to success [company job advertisements, career development]. IEEE Spectrum, 32(3):62–64, March 1995. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Hoschette:1995:MH

[Hos95b] J. A. Hoschette. A mentor in hand. IEEE Spectrum, 32(2): 56–58, February 1995. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

House:1995:CRS

[Hou95] R. House. Choosing the right software for data acquisition. IEEE Spectrum, 32(5):24–26, May 1995. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Hickerson:1992:EWM

[HPV92] L. B. Hickerson, C. S. Pervier, and P. Valdes. Engineering workstations — managing network workstations. IEEE Spectrum, 29(4):55–56, April 1992. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Hunt:1996:BSC

[HR96] M. Hunt and J. A. Rowson. Blocking in a system on a chip. IEEE Spectrum, 33(11): 35–41, November 1996. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Hawkins:1996:VDY

[HS96a] Charles F. Hawkins and Jerry M. Soden. Viewpoint: Dealing with yield losses in I_{DDQ} testing. IEEE Spectrum, 33(1):68–69, January 1996. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Hunt:1996:UGE

[HS96b] S. Hunt and G. Shuttleworth. Unlocking the grid [electricity industry restructuring]. IEEE Spectrum, 33(7):20–25, July 1996. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Hermance:1998:HEV

[HS98] D. Hermance and S. Sasaki.

Hybrid electric vehicles take to the streets. IEEE Spectrum, 35(11):48–52, November 1998. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Hood:1999:EPP

[HS99] M. Hood and W. Sweet. Energy policy and politics in China. IEEE Spectrum, 36 (11):34–38, November 1999. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Hoshen:1995:KTC

[HSW95] J. Hoshen, J. Sennott, and M. Winkler. Keeping tabs on criminals [electronic monitoring]. IEEE Spectrum, 32(2): 26–32, February 1995. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Horne:1995:NWN

[HT95] W. D. Horne and R. M. Taylor. New waves for new systems [radio frequency spectrum allocation]. IEEE Spectrum, 32(10):72–77, October 1995. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Hull:1995:TA

[Hul95]  S. Hull.  Technology for

all. IEEE Spectrum, 32(12): 6–??, December 1995. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Hull:1997:FR

[Hul97] J. R. Hull. Flywheels on a roll. IEEE Spectrum, 34(7):20–25, July 1997. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Hopf:1994:TID

[HW94] S. Hopf and T. Warneke. Tools interact in design analyses. IEEE Spectrum, 31(6):41–43, June 1994. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Icove:1997:CCI

[Ico97] D. J. Icove. Collaring the cybercrook: an investigator’s view. IEEE Spectrum, 34 (6):31–36, June 1997. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Iyer:1999:EDT

[IK99] S. S. Iyer and H. L. Kalter. Embedded DRAM technology: opportunities and challenges. IEEE Spectrum, 36 (4):56–64, April 1999. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Inacio:1996:DDF

[IO96] C. Inacio and D. Ombres.

The DSP decision: fixed point or floating? IEEE Spectrum, 33(9):72–74, September 1996. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Ishii:1994:ES

[Ish94]  T. Ishii.  Elevators for

skyscrapers. IEEE Spectrum, 31(9):42–46, September 1994. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Itoh:1990:NTO

[Ito90] N. Itoh. New tricks for an old power source (fuel cells). IEEE Spectrum, 27(9): 40–43, September 1990. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jaikumar:1993:YC

[Jai93] R. Jaikumar. 200 years to CIM. IEEE Spectrum, 30(9): 26–27, September 1993. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

James:1998:TRC

[Jam98] J. V. James. Tracking the right clues with exploratory data analysis. IEEE Spectrum, 35(7):58–65, July 1998. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jayadev:1995:HW

[Jay95] J. Jayadev. Harnessing the wind. IEEE Spectrum, 32(11): 78–83, November 1995. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jankovic:1996:EWS

[JB96] G. Jankovic and L. Black. Engineering a Web site. IEEE Spectrum, 33(11):62– 69, November 1996. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jangi:1991:ESA

[JJ91] S. Jangi and Y. Jain. Embedding spectral analysis in equipment. IEEE Spectrum, 28(2):40–43, February 1991. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jamison:1998:VYF

[JNM^{+}98] J. Jamison, R. Nicklas, G. Miller, K. Thompson, R. Wilder, L. Cunningham, and C. Song. vBNS: not your father’s Internet. IEEE Spectrum, 35(7):38–46, July 1998. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Johnson:1990:MMN

[Joh90]  R. R. Johnson.  Multi-

chip modules: next-generation packages. IEEE Spectrum, 27 (3):34–36, March 1990. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Johnson:1993:HRM

[Joh93]  R. F. Johnson.  Hitting

the road with manufacturing know-how (education). IEEE Spectrum, 30(9):81, September 1993. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Johnson:1994:AMS

[Joh94] M. Johnson. The all-metal spin transistor. IEEE Spectrum, 31(5):47–51, May 1994. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Johnston:1998:EFM

[Joh98] W. Johnston. Europe’s future mobile telephony system. IEEE Spectrum, 35(10): 49–53, October 1998. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jones:1992:CME

[Jon92] C. Jones. CASE’s missing elements. IEEE Spectrum, 29 (6):38–41, June 1992. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jones:1997:TYA

[Jon97] W. D. Jones. Thirty years ago …. IEEE Spectrum, 34(12):64T10–64T12, December 1997. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jones:1998:BDS

[Jon98a] C. Jones. Bad days for software. IEEE Spectrum, 35(9): 47–52, September 1998. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jones:1998:TYAg

[Jon98b] W. Jones. Thirty years ago …. IEEE Spectrum, 35 (8):76T4–76T6, August 1998. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jones:1998:TYAh

[Jon98c] W. Jones. Thirty years ago …. IEEE Spectrum, 35(9): 80T2–80T4, September 1998. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jones:1998:TYAi

[Jon98d] W. Jones. Thirty years ago …. IEEE Spectrum, 35(10): 80T4–80T6, October 1998. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jones:1998:TYAj

[Jon98e] W. Jones. Thirty years ago …. IEEE Spectrum, 35(11): 60T11, November 1998. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jones:1998:TYAa

[Jon98f] W. D. Jones. Thirty years ago …. IEEE Spectrum, 35 (1):108T8, January 1998. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jones:1998:TYAb

[Jon98g] W. D. Jones. Thirty years ago …. IEEE Spectrum, 35 (2):68t2–68t5, February 1998. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jones:1998:TYAc

[Jon98h] W. D. Jones. Thirty years ago …. IEEE Spectrum, 35(3): 68T4–68T7, March 1998. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jones:1998:TYAd

[Jon98i] W. D. Jones. Thirty years ago …. IEEE Spectrum, 35 (4):72T3, April 1998. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jones:1998:TYAe

[Jon98j] W. D. Jones. Thirty years ago …. IEEE Spectrum, 35(5): 60T7–60T8, May 1998. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jones:1998:TYAf

[Jon98k] W. D. Jones. Thirty years ago …. IEEE Spectrum, 35(6): 84T2–84T6, June 1998. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jones:1998:TYAk

[Jon98l] W. D. Jones. Thirty years ago …. IEEE Spectrum, 35(12): 64T3–64T6, December 1998. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jones:1999:I

[Jon99a]  W. D. Jones.  Innova-

tions. IEEE Spectrum, 36 (1):112, January 1999. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jones:1999:TYAa

[Jon99b] W. D. Jones. Thirty years ago …. IEEE Spectrum, 36(1): 112T2–112T4, January 1999. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jones:1999:TYAb

[Jon99c] W. D. Jones. Thirty years ago …. IEEE Spectrum, 36(2): 64T2–64T5, February 1999. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jones:1999:TYAc

[Jon99d] W. D. Jones. Thirty years ago …. IEEE Spectrum, 36(3): 16T6–16T8, March 1999. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jones:1999:TYAd

[Jon99e] W. D. Jones. Thirty years ago …. IEEE Spectrum, 36 (4):78R10–78R13, April 1999. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jones:1999:TYAe

[Jon99f] W. D. Jones. Thirty years ago …. IEEE Spectrum, 36 (5):66R2, May 1999. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jones:1999:TYAf

[Jon99g] W. D. Jones. Thirty years ago …. IEEE Spectrum, 36(6): 82R2–82R4, June 1999. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jones:1999:TYAg

[Jon99h] W. D. Jones. Thirty years ago …. IEEE Spectrum, 36(7): 88R14–88R15, July 1999. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jones:1999:TYAh

[Jon99i] W. D. Jones. Thirty years ago …. IEEE Spectrum, 36(8): 72R9–72R11, August 1999. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jones:1999:TYAi

[Jon99j] W. D. Jones. Thirty years ago …. IEEE Spectrum, 36(9): 90R2–90R7, September 1999. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jones:1999:TYAj

[Jon99k] W. D. Jones. Thirty years ago …. IEEE Spectrum, 36(10): 74R11–74R13, October 1999. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jones:1999:TYAk

[Jon99l] W. D. Jones. Thirty years ago …. IEEE Spectrum, 36(11): 82R2–82R6, November 1999. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jones:1999:TYAl

[Jon99m] W. D. Jones. Thirty years ago …. IEEE Spectrum, 36(12): 88R2–88R11, December 1999. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jones:1992:MNE

[JPN^{+}92] F. Jones, B. Prince, R. Norwood, J. Hartigan, W. C. Vogley, C. A. Hart, and D. Bondurant. Memory — a new era of fast dynamic RAMs (for video applications). IEEE Spectrum, 29(10):43–45, October 1992. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jones:1999:NTG

[JS99] D. A. Jones and R. L. Skelton. The next threat to grid reliability-data security. IEEE Spectrum, 36(6):46–48, June 1999. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jayaraman:1993:MSS

[JT93] R. Jayaraman and P. Toole.

Manufacturing systems simulated. IEEE Spectrum, 30(9): 60–62, September 1993. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Juliussen:1994:WLE

[Jul94] E. Juliussen. Which low-end workstation? IEEE Spectrum, 31(4):51–54, April 1994. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Juliussen:1995:SC

[Jul95] E. Juliussen. Small computers. IEEE Spectrum, 32(1):

44–47, January 1995. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Juliussen:1997:C

[Jul97] E. Juliussen. Computers. IEEE Spectrum, 34(1):49– 54, January 1997. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jurgen:1990:ACE

[Jur90a]  R. K. Jurgen.  Applica-

tions ’90: consumer electronics. IEEE Spectrum, 27(2): 39–41, February 1990. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jurgen:1990:FAH

[Jur90b] R. K. Jurgen. Friendly adversaries help US companies. IEEE Spectrum, 27(5):50–51, May 1990. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jurgen:1990:GCM

[Jur90c] R. K. Jurgen. GlobalR&D for a 21st century manufacturing system. IEEE Spectrum, 27(11):122–124, November 1990. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jurgen:1990:PEW

[Jur90d] R. K. Jurgen. Putting electronics to work in the 1991 car models. IEEE Spectrum, 27(12):72–75, December 1990. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jurgen:1990:VTR

[Jur90e] R. K. Jurgen. View from the top: responding to the business divisions’ needs (R&D). IEEE Spectrum, 27(10):34– 39, October 1990. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jurgen:1991:CDH

[Jur91a] R. K. Jurgen. The challenges of digital HDTV. IEEE Spectrum, 28(4):28–30, April 1991. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jurgen:1991:SCH

[Jur91b] R. K. Jurgen. Smart cars and highways go global. IEEE Spectrum, 28(5):26–36, May 1991. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jurgen:1991:TCE

[Jur91c]  R. K. Jurgen.  Technol-

ogy 1991: consumer electronics. IEEE Spectrum, 28(1): 65–68, January 1991. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jurgen:1991:TSN

[Jur91d] R. K. Jurgen. Technology 1991: the specialities neural networks, TAPS, integrated document interchange. IEEE Spectrum, 28(1):79, January 1991. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jurgen:1992:DV

[Jur92a]  R. K. Jurgen.  Digital

video. IEEE Spectrum, 29 (3):24–30, March 1992. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jurgen:1992:GBT

[Jur92b] R. K. Jurgen. Good-bye to TV ghosts. IEEE Spectrum, 29(7):50–52, July 1992. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jurgen:1992:TCE

[Jur92c] R. K. Jurgen. Technology 1992 — consumer electronics. IEEE Spectrum, 29(1): 52–54, January 1992. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jurgen:1992:TME

[Jur92d]  R. K. Jurgen.  Technol-

ogy 1992 — medical electronics. IEEE Spectrum, 29(1): 61–62, January 1992. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jurgen:1992:TS

[Jur92e] R. K. Jurgen. Technology 1992 — the specialities. IEEE Spectrum, 29(1):63, January 1992. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jurgen:1992:TT

[Jur92f] R. K. Jurgen. Technology 1992 — transportation. IEEE Spectrum, 29(1):55–57, January 1992. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jurgen:1993:TCE

[Jur93] R. K. Jurgen. Technology 1993 — consumer electronics. IEEE Spectrum, 30(1): 65–67, January 1993. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jurgen:1995:EM

[Jur95] R. K. Jurgen. The electronic motorist. IEEE Spectrum, 32 (3):37–48, March 1995. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jurgen:1996:BDA

[Jur96] R. K. Jurgen. Broadcasting with digital audio. IEEE Spectrum, 33(3):52–59, March 1996. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jurgen:1991:PTL

[JWA^{+}91] R. K. Jurgen, G. Watson, J. A. Adam, T. E. Bell, K. Esch, K. Fitzgerald, Glenn Zorpette, and D. L. DiMaria. Poised for technological leadership. IEEE Spectrum, 28 (6):30–42, June 1991. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kappernman:1990:BGS

[KA90] J. G. Kappernman and V. D. Albertson. Bracing for the geomagnetic storms. IEEE Spectrum, 27(3):27–33, March 1990. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kodali:1994:EDT

[KA94] V. P. Kodali and V. K. Aatre. Electronics and defense: turning setbacks into successes. IEEE Spectrum, 31 (3):47–50, March 1994. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kalinin:1994:SSN

[Kal94] A. V. Kalinin. Secrets of the Soviet nuclear complex. IEEE Spectrum, 31(5):32–38, May 1994. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kam:1999:NCV

[Kam99] M. Kam. New chapter and verse [book review]. IEEE Spectrum, 36(2):10–12, February 1999. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kaplan:1990:AIE

[Kap90a] Gadi Kaplan. Applications ’90: industrial electronics. IEEE Spectrum, 27(2):42– 43, February 1990. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kaplan:1990:ESR

[Kap90b] Gadi Kaplan. EC ’92: status report. IEEE Spectrum, 27(6):22–24, June 1990. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kaplan:1991:TIE

[Kap91] Gadi Kaplan. Technology 1991: industrial electronics. IEEE Spectrum, 28(1): 59–60, January 1991. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kaplan:1992:EST

[Kap92a] Gadi Kaplan. Engineering software — talking about tools. IEEE Spectrum, 29(11): 32–33, November 1992. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kaplan:1992:TIE

[Kap92b] Gadi Kaplan. Technology 1992 — industrial electronics. IEEE Spectrum, 29(1): 47–48, January 1992. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kaplen:1992:EWU

[Kap92c]  G. Kaplen.  Engineering

workstations — the user’s view. IEEE Spectrum, 29 (4):30–31, April 1992. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kaplan:1993:TIE

[Kap93] Gadi Kaplan. Technology 1993 — industrial electronics. IEEE Spectrum, 30(1): 58–60, January 1993. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kaplan:1994:ATP

[Kap94]  Gadi Kaplan.  Appropri-

ate technologies [photovoltaic power generation]. IEEE Spectrum, 31(10):32, October 1994. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kaplan:1995:IE

[Kap95] Gadi Kaplan. Industrial electronics. IEEE Spectrum, 32 (1):84–86, January 1995. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kaplan:1996:IE

[Kap96] Gadi Kaplan. Industrial electronics. IEEE Spectrum, 33(1):96–100, January 1996. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kaplan:1997:AMD

[Kap97a] Gadi Kaplan. Auto manufacture digitizes in depth. IEEE Spectrum, 34(11):62– 69, November 1997. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kaplan:1997:IET

[Kap97b] Gadi Kaplan. Industrial electronics [technology analysis and forecast]. IEEE Spectrum, 34(1):79–83, January 1997. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kaplan:1998:IET

[Kap98a] Gadi Kaplan. Industrial electronics [technology 1998 analysis and forecast]. IEEE Spectrum, 35(1):73–78, January 1998. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kaplan:1998:IHT

[Kap98b] Gadi Kaplan. Israel: a hightech haven. IEEE Spectrum, 35(5):22–32, May 1998. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kaplan:1999:CCP

[Kap99a] Gadi Kaplan. Charles Concordia [power engineer biography]. IEEE Spectrum, 36 (6):28–33, June 1999. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kaplan:1999:IET

[Kap99b] Gadi Kaplan. Industrial electronics [technology 1999 analysis and forecast]. IEEE Spectrum, 36(1):68–72, January 1999. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kaplan:1999:USE

[Kap99c] Gadi Kaplan. Update on software engineering. IEEE Spectrum, 36(4):29, April 1999. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Karnofsky:1996:SDA

[Kar96a] K. Karnofsky. Speeding DSP algorithm design. IEEE Spectrum, 33(7):79–82, July 1996. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Karray:1996:AFB

[Kar96b] F. Karray. Applied fuzziness [books]. IEEE Spectrum, 33(10):10–14, October 1996. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kasturirangan:1994:ATT

[Kas94] K. Kasturirangan. Aerospace technologies: a terrestrial focus. IEEE Spectrum, 31(3): 39–43, March 1994. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kelley:1997:FEM

[Kel97] E. W. Kelley. The future of electronic money: a regulator’s perspective. IEEE Spectrum, 34(2):20–22, February 1997. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kelley:1999:HSE

[Kel99] R. E. Kelley. How to be a star engineer. IEEE Spectrum, 36(10):51–58, October 1999. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Khan:1996:WEE

[Kha96] A. R. Khan. Workhorses of the electronic era [microcontrollers]. IEEE Spectrum, 33(10):36–39, October 1996. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Khermouch:1994:LC

[Khe94] G. Khermouch. Large computers. IEEE Spectrum, 31 (1):46–49, January 1994. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Khermouch:1995:LC

[Khe95] G. Khermouch. Large computers. IEEE Spectrum, 32 (1):48–51, January 1995. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

King:1995:HET

[KHSK95] R. D. King, K. B. Haefner, L. Salasoo, and R. A. Koegl. Hybrid electric transit bus pollutes less, conserves fuel. IEEE Spectrum, 32(7):26–31, July 1995. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

King:1993:NPI

[Kin93] J. King. New peripherals increase options. IEEE Spectrum, 30(5):69–76, May 1993. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kirtley:1996:IMF

[Kir96] J. Kirtley. Imaging magnetic fields. IEEE Spectrum, 33(12): 40–43, December 1996. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Koch:1996:NVC

[KM96] C. Koch and B. Mathur. Neuromorphic vision chips. IEEE Spectrum, 33(5):38–46, May 1996. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kobb:1993:PW

[Kob93] B. Z. Kobb. Personal wireless.

IEEE Spectrum, 30(6):20–25, June 1993. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kobetic:1994:ASS

[Kob94] R. Kobetic. Advancing step by step. IEEE Spectrum, 31(10): 27–31, October 1994. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kobb:1995:T

[Kob95] B. Z. Kobb. Telecommunications. IEEE Spectrum, 32 (1):30–34, January 1995. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Koenen:1999:MMO

[Koe99] R. Koenen. MPEG-4 multimedia for our time. IEEE Spectrum, 36(2):26–33, February 1999. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kokubo:1993:CI

[Kok93] A. Kokubo. Competitive intelligence. IEEE Spectrum, 30 (8):44–46, August 1993. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kornbluh:1993:ESS

[Kor93] Ken Kornbluh. Engineering software — seeing data in action. IEEE Spectrum, 30(11): 60–64, November 1993. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kornbluh:1994:ADA

[Kor94] Ken Kornbluh. Active data analysis: advanced software for the ’90s. IEEE Spectrum, 31(11):57–70, November 1994. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kornbluh:1999:MSS

[Kor99] Ken Kornbluh. Math and science software [technology 1999 analysis and forecast]. IEEE Spectrum, 36(1):88–91, January 1999. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kaplan:1998:EE

[KP98] Gadi Kaplan and T. Perry. Engineers as entrepreneurs. IEEE Spectrum, 35(8):14– 23, August 1998. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kaplan:1990:EWR

[KR90] Gadi Kaplan and A. Rosenblatt. The expanding world of R&D. IEEE Spectrum, 27(10): 28–33, October 1990. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kramish:1996:EEB

[Kra96] A. Kramish. Eavesdropping at its electronic best. IEEE Spectrum, 33(5):11–12, May 1996. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kramish:1998:BBB

[Kra98] A. Kramish. Blueprinting the bomb [books]. IEEE Spectrum, 35(4):12–14, April 1998. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Karr:1997:SSM

[KRF97] C. R. Karr, D. Reece, and R. Franceschini. Synthetic soldiers [military training simulators]. IEEE Spectrum, 34 (3):39–45, March 1997. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Koopman:1992:EWI

[KS92] P. Koopman and D. P.

Siewiorek. Engineering workstations — ICs: the brains of a workstation. IEEE Spectrum, 29(4):52–54, April 1992. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kuhn:1991:IPM

[Kuh91] D. R. Kuhn. IEEE’s Posix:

making progress. IEEE Spectrum, 28(12):36–39, December 1991. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kukuljian:1993:GAP

[Kuk93] V. Kukuljian. Growing an automated production line. IEEE Spectrum, 30(9):37– 39, September 1993. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kurland:1992:SHH

[Kur92] M. Kurland. Surviving hell and high water (electronic equipment design). IEEE Spectrum, 29(5):44–47, May 1992. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kurz:1999:RAE

[Kur99]  H. Kurz.  Rolling across

Europe’s vanishing frontiers [electric railway technology]. IEEE Spectrum, 36(2):44– 49, February 1999. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kahaner:1992:SJC

[KW92] D. K. Kahaner and U. Wattenberg. Supercomputing-Japan: a competitive assessment. IEEE Spectrum, 29(9): 42–47, September 1992. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kephart:1993:CE

[KWC93] J. O. Kephart, S. R. White, and D. M. Chess. Computers and epidemiology. IEEE Spectrum, 30(5):20–26, May 1993. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kaplan:1992:CP

[KWCR92] Gadi Kaplan, G. F. Watson, L. Curran, and M. J. Riezenman. Companies in profile. IEEE Spectrum, 29(12): 37–45, December 1992. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kassakian:1996:AES

[KWMH96] John G. Kassakian, Hans-Christoph Wolf, John M. Miller, and Charles J. Hurton. Automotive electrical systems circa 2005. IEEE Spectrum, 33 (8):22–27, August 1996. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lane:1994:AKV

[Lan94] J. Lane. ATM knits voice, data on any net. IEEE Spectrum, 31(2):42–45, February 1994. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lange:1998:ITA

[Lan98] L. Lange. The Internet [technology 1998 analysis and forecast]. IEEE Spectrum, 35(1): 37–42, January 1998. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lange:1999:ITA

[Lan99] L. Lange. The Internet [technology 1999 analysis and forecast]. IEEE Spectrum, 36(1): 35–40, January 1999. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Larsen:1993:RR

[Lar93] R. S. Larsen. Running out of resources. IEEE Spectrum, 30(7):43–45, July 1993. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Larocque:1994:CST

[Lar94] J. Larocque. Client–server trends. IEEE Spectrum, 31 (4):48–50, April 1994. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Large:1995:CNI

[Lar95] D. Large. Creating a network for interactivity. IEEE Spectrum, 32(4):58–63, April 1995. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Latham:1997:CDB

[Lat97] C. Latham. A crucial demarche [books]. IEEE Spectrum, 34(8):10–12, August 1997. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lauterbach:1997:GME

[Lau97] M. Lauterbach. Getting more out of eye diagrams [circuit testing]. IEEE Spectrum, 34 (3):60–63, March 1997. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lazzaro:1999:HWH

[Laz99] J. J. Lazzaro. Helping the Web help the disabled. IEEE Spectrum, 36(3):54–59, March 1999. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lapsley:1996:HED

[LB96] P. Lapsley and G. Blalock.

How to estimate DSP processor performance. IEEE Spectrum, 33(7):74–78, July 1996. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lassen:1999:S

[LC99] C. L. Lassen and M. V. Christensen. The via squeeze. IEEE Spectrum, 36(10):36–41, October 1999. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Liboni:1996:LEP

[LD96] C. Liboni and C. De Lacerda. On the leading edge of the process control world. IEEE Spectrum, 33(7):44–47, July 1996. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lenstra:1996:VNS

[Len96] A. K. Lenstra. Viewpoint: Network security — elusive, essential. IEEE Spectrum, 33 (1):32–33, January 1996. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Leung:1997:SPP

[Leu97] E. Leung. Surge protection for power grids. IEEE Spectrum, 34(7):26–30, July 1997. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Levitt:1992:ATU

[Lev92] M. E. Levitt. ASIC testing upgraded. IEEE Spectrum, 29 (5):26–29, May 1992. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lewis:1995:SGW

[Lew95] M. A. Lewis. Sleuthing out gravitational waves. IEEE Spectrum, 32(5):57–61, May 1995. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

lHunt:1998:GBS

[lH98] G. l. Hunt. The great battery search [electric vehicles]. IEEE Spectrum, 35(11):21– 28, November 1998. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lathrop:1990:FAA

[LHD^{+}90] R. H. Lathrop, R. J. Hall, G. Duffy, K. M. Alexander, and R. S. Kirk. ‘Functional abstraction’ anticipates timing glitches. IEEE Spectrum, 27(4):41–42, April 1990. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Likourezos:1992:JLB

[Lik92] G. Likourezos. Jan. 28, 1958:

A laser is born. IEEE Spectrum, 29(5):43, May 1992. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lindley:1996:NMK

[Lin96] D. Lindley. No more to know?

[books]. IEEE Spectrum, 33 (7):10–12, July 1996. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lazarus-Karaoglan:1999:BOE

[LK99a] T. Lazarus-Karaoglan. Bringing order to the embedded space. IEEE Spectrum, 36 (4):65–72, April 1999. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lazarus-Karaoglan:1999:MET

[LK99b] T. Lazarus-Karaoglan. Medical electronics [technology 1999 analysis and forecast]. IEEE Spectrum, 36(1):79– 83, January 1999. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lucas:1995:FGR

[LLKS95] H. C. Lucas, H. Levecq, R. Kraut, and L. Streeter. France’s grass roots data net. IEEE Spectrum, 32(11):71– 77, November 1995. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Locke:1995:TA

[Loc95] T. Locke. Technology for all. IEEE Spectrum, 32(12): 8–??, December 1995. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Logan:1999:BBE

[Log99] J. Logan. Balancing the books on energy pricing. IEEE Spectrum, 36(12):59–63, December 1999. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lowe:1992:TDC

[Low92] S. J. Lowe. Technology, 1992data communications. IEEE Spectrum, 29(1):39–41, January 1992. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lowe:1995:DC

[Low95] S. J. Lowe. Data communications. IEEE Spectrum, 32 (1):26–29, January 1995. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lowery:1997:CAP

[Low97] A. J. Lowery. Computer-aided photonics design. IEEE Spectrum, 34(4):26–31, April 1997. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lewis:1997:FCB

[LP97a] G. N. Lewis and T. A. Postol.

Future challenges to ballistic missile defense. IEEE Spectrum, 34(9):60–68, September 1997. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lewis:1997:WPR

[LP97b] G. N. Lewis and T. A. Postol. Why the picture is right: a reply to Stein and Little. IEEE Spectrum, 34(11): 73–74, November 1997. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic). See [SLLP97].

Lefkon:1998:PEM

[LP98] D. Lefkon and B. Payne.

The practical engineer. Making embedded systems Year 2000 compliant. IEEE Spectrum, 35(6):74–79, June 1998. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lelieveld:1999:GEA

[LRC99] J. Lelieveld, V. Ramanathan, and P. J. Crutzen. The global effects of Asian haze [air pollution]. IEEE Spectrum, 36(12): 50–54, December 1999. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lodin:1998:FFI

[LS98] S. W. Lodin and C. L. Schuba. Firewalls fend off invasions from the net. IEEE Spectrum, 35(2):26–34, February 1998. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lubell:1995:GSH

[Lub95] P. D. Lubell. The gathering storm in high-density compact disks. IEEE Spectrum, 32 (8):32–37, August 1995. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lubell:1997:HW

[Lub97] P. Lubell. HDTV in Wonderland. IEEE Spectrum, 34(9): 16–18, September 1997. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Luciano:1993:DTY

[Luc93]  V. P. Luciano.  Display

terminals as you like them FMS. IEEE Spectrum, 30(9): 35–37, September 1993. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lucky:1994:BBB

[Luc94a] R. W. Lucky. A bit is a bit is a bit? [reflections]. IEEE Spectrum, 31(7):15, July 1994. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lucky:1994:DSC

[Luc94b] R. W. Lucky. The disappearing space crunch [reflections]. IEEE Spectrum, 31(1):4, January 1994. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lucky:1996:BBR

[Luc96a] R. W. Lucky. Bozos on the bus [reflections]. IEEE Spectrum, 33(7):19, July 1996. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lucky:1996:GFR

[Luc96b] R. W. Lucky. The glitz factor [reflections]. IEEE Spectrum, 33(9):24, September 1996. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lucky:1996:LWT

[Luc96c] R. W. Lucky. Have laptop, will travel [reflections]. IEEE Spectrum, 33(1):21, January 1996. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lucky:1996:HIY

[Luc96d] R. W. Lucky. Honk if you’re for anarchy [reflections]. IEEE Spectrum, 33(11):24, November 1996. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lucky:1996:WHR

[Luc96e] R. W. Lucky. Why am i here? [reflections]. IEEE Spectrum, 33(3):19, March 1996. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lucky:1997:TAR

[Luc97a] R. W. Lucky. Any time, anywhere [reflections]. IEEE Spectrum, 34(7):17, July 1997. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lucky:1997:BOR

[Luc97b] R. W. Lucky. Blinking 12 o’clock [reflections]. IEEE Spectrum, 34(3):17, March 1997. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lucky:1997:FEN

[Luc97c] R. W. Lucky. Fatal error number 27 [reflections]. IEEE Spectrum, 34(5):18, May 1997. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lucky:1997:JPR

[Luc97d] R. W. Lucky. The junk plague [reflections]. IEEE Spectrum, 34(9):25, September 1997. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lucky:1997:R

[Luc97e]  R. W. Lucky.  Reflec-

tions. IEEE Spectrum, 34 (1):21, January 1997. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lucky:1997:WDS

[Luc97f] R. W. Lucky. When is dumb smart? [reflections]. IEEE Spectrum, 34(11):21, November 1997. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lucky:1998:CSR

[Luc98a] R. W. Lucky. Clock speed [reflections]. IEEE Spectrum, 35(7):18, July 1998. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lucky:1998:EED

[Luc98b] R. W. Lucky. Electrical engineering: A diminishing role [reflections]. IEEE Spectrum, 35(5):21, May 1998. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lucky:1998:JCR

[Luc98c] R. W. Lucky. Job churn [reflections]. IEEE Spectrum, 35 (11):17, November 1998. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lucky:1998:LIM

[Luc98d] R. W. Lucky. Life in the interrupt mode [reflections]. IEEE Spectrum, 35(3):25, March 1998. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lucky:1998:MLR

[Luc98e] R. W. Lucky. Moore’s Law redux [reflections]. IEEE Spectrum, 35(9):17, September 1998. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lucky:1998:WAP

[Luc98f] R. W. Lucky. The world according to power point [reflections]. IEEE Spectrum, 35 (1):17, January 1998. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lucky:1999:DPR

[Luc99a] R. W. Lucky. Design for people — NOT [reflections]. IEEE Spectrum, 36(7):20–21, July 1999. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lucky:1999:FSR

[Luc99b] R. W. Lucky. Free software [reflections]. IEEE Spectrum, 36(5):24, May 1999. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lucky:1999:IAR

[Luc99c] R. W. Lucky. Information addiction [reflections]. IEEE Spectrum, 36(9):32, September 1999. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lucky:1999:Ra

[Luc99d]  R. W. Lucky.  Reflec-

tions. IEEE Spectrum, 36 (1):24, January 1999. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lucky:1999:Rb

[Luc99e]  R. W. Lucky.  Reflec-

tions. IEEE Spectrum, 36 (3):19, March 1999. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lucky:1999:SPR

[Luc99f] R. W. Lucky. Special places [reflections]. IEEE Spectrum, 36(11):27, November 1999. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Li:1996:ETB

[LY96] Ping-Chung Li and T. K. Young. Electromigration: the time bomb in deep-submicron ICs. IEEE Spectrum, 33(9): 75–78, September 1996. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lynch:1993:CCE

[Lyn93] D. Lynch. The Calstart consortium (electric vehicle industry). IEEE Spectrum, 30 (7):54–57, July 1993. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lysen:1994:PV

[Lys94] E. H. Lysen. Photovolts for villages. IEEE Spectrum, 31(10):34–39, October 1994. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Mountford:1999:KLP

[MA99] J. D. Mountford and R. R. Austria. Keeping the lights on [power system reliability]. IEEE Spectrum, 36(6):34–39, June 1999. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Maccoby:1991:IMW

[Mac91] M. Maccoby. The innovative mind at work. IEEE Spectrum, 28(12):23–35, December 1991. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Mackail:1998:RSB

[Mac98] L. Mackail. Recipes for success (book review). IEEE Spectrum, 35(1):11, January 1998. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Magee:1997:EEI

[Mag97]  S. Magee.  The environ-

ment [electronics industry aspects]. IEEE Spectrum, 34 (1):94–98, January 1997. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Makukatin:1994:WAS

[Mak94] S. Makukatin. Water from the African sun. IEEE Spectrum, 31(10):40–43, October 1994. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Malozemoff:1993:ATE

[Mal93] A. P. Malozemoff. Advanced technology/electromagnetism — superconducting wire gets hotter. IEEE Spectrum, 30 (12):26–30, December 1993. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Martzloff:1990:PCS

[Mar90] F. Martzloff. Protecting computer systems against power transients. IEEE Spectrum, 27 (4):37–40, April 1990. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Martin:1996:TTD

[Mar96] C. D. Martin. Three takes on the digital age,book reviews. IEEE Spectrum, 33(1): 10–116T2, January 1996. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Martin:1997:MEH

[Mar97a] C. D. Martin. The media equation: How people treat computers, television and new media like real people and places [book review]. IEEE Spectrum, 34(3):9–10, March 1997. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Martin:1997:SV

[Mar97b] C. D. Martin. Single vision. IEEE Spectrum, 34(11): 15–16, November 1997. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Marcopulos:1998:FBC

[Mar98a] T. Marcopulos. Faster, better, cheaper space exploration. IEEE Spectrum, 35 (8):68–74, August 1998. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Marcus:1998:CER

[Mar98b]  S. J. Marcus.  Consort-

ing with the enemy, R&Dwise. IEEE Spectrum, 35(9):

64–70, September 1998. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Marcus:1998:GIR

[Mar98c] S. J. Marcus. Grand illusion:

the risks of Cassini [nuclear space power]. IEEE Spectrum, 35(3):58–65, March 1998. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Marcus:1998:WDA

[Mar98d] S. J. Marcus. What to do about bolts from the blue. IEEE Spectrum, 35(12):34– 41, December 1998. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Marcus:1998:WM

[Mar98e] S. J. Marcus. Woe is mir.

IEEE Spectrum, 35(6):38–44, June 1998. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Martin:1999:EDA

[Mar99a] B. Martin. Electronic design automation [1999 technology analysis and forecast]. IEEE Spectrum, 36(1):57–61, January 1999. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Martin:1999:RYR

[Mar99b] C. D. Martin. Reading your rights. IEEE Spectrum, 36(9):

16–18, September 1999. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Masuoka:1990:YRN

[Mas90] F. Masuoka. Are you ready for next-generation dynamic RAM chips? IEEE Spectrum, 27(11):110–112, November 1990. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Masi:1995:REE

[Mas95] C. G. Masi. Re-engineering engineering education. IEEE Spectrum, 32(9):44–47, September 1995. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Masiello:1996:PSG

[Mas96] R. D. Masiello. It’s put up or shut up for grid controls. IEEE Spectrum, 33(8): 50–51, August 1996. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Masi:1998:TMT

[Mas98] C. G. Masi. Test measurement [technology 1998 analysis and forecast]. IEEE Spectrum, 35 (1):79–83, January 1998. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Masi:1999:SRO

[Mas99a] C. C. Masi. Software reviews — one highly adaptable software package. IEEE Spectrum, 36(6):94–95, June 1999. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Masi:1999:SSD

[Mas99b] C. G. Masi. Software smartens up data manipulation. IEEE Spectrum, 36(4):30–35, April 1999. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Masi:1999:TMT

[Mas99c] C. G. Masi. Test and measurements [1999 technology analysis and forecast]. IEEE Spectrum, 36(1):92–97, January 1999. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Mather:1990:NDS

[Mat90] B. C. Mather. Needed DSP software emerges. IEEE Spectrum, 27(11):52–59, November 1990. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Mather:1991:ED

[Mat91] B. G. Mather. Embedding DSP. IEEE Spectrum, 28(11): 52–55, November 1991. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Maunder:1994:DT

[Mau94] C. Maunder. Documentation on tap. IEEE Spectrum, 31(9):

52–56, September 1994. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Mehta:1997:TTS

[MAW97] S. P. Mehta, N. Aversa, and M. S. Walker. Transforming transformers [superconducting windings]. IEEE Spectrum, 34(7):43–49, July 1997. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

May:1994:MIN

[May94] G. S. May. Manufacturing ICs the neural way. IEEE Spectrum, 31(9):47–51, September 1994. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Morreale:1990:MAN

[MC90] P. A. Morreale and G. M. Campbell. Metropolitan-area networks. IEEE Spectrum, 27 (5):40–42, May 1990. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Mackey:1994:MSS

[MC94] W. A. Mackey and J. C. Carter. Measure the steps to success [concurrent engineering]. IEEE Spectrum, 31 (6):33–38, June 1994. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

McAdoo:1998:CSN

[McA98]  M. McAdoo.  Companies

steer new K–12 courses [education]. IEEE Spectrum, 35 (4):27–33, April 1998. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

McCracken:1992:CDT

[McC92] W. P. McCracken. CCDing in the dark (thermal imaging sensors). IEEE Spectrum, 29 (5):30–34, May 1992. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

McChesney:1997:BCI

[McC97] M. C. McChesney. Banking in cyberspace: an investment in itself. IEEE Spectrum, 34 (2):54–59, February 1997. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

McMillan:1996:VCC

[McM96] K. McMillan. Verifying cache coherence protocols. IEEE Spectrum, 33(6):67, June 1996. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

McTavish:1996:ECC

[McT96] D. McTavish. Equivalence checking at Cray Research. IEEE Spectrum, 33(6):71, June 1996. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Mecham:1995:STS

[Mec95] M. Mecham. To the skies through the sea [Hong Kong airport project]. IEEE Spectrum, 32(10):40–43, October 1995. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Memon:1997:DCE

[Mem97] N. Memon. Digital compression exegesis, straight from the apostles. IEEE Spectrum, 34(9):18–19, September 1997. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Memon:1998:VSC

[Mem98] N. Memon. A video standards compendium [books]. IEEE Spectrum, 35(4):14–15, April 1998. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Meserve:1998:SCB

[Mes98] R. A. Meserve. Science in the courts [books]. IEEE Spectrum, 35(6):10–16, June 1998. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Metcalfe:1997:WWB

[MF97] B. Metcalfe and K. Foster. Winning with Wintel business strategies. IEEE Spectrum, 34(5):7–12, May 1997. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

McMichael:1998:RB

[MH98] F. C. McMichael and C. Henderson. Recycling batteries. IEEE Spectrum, 35(2): 35–42, February 1998. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Miller:1994:HA

[Mil94a] A. Miller. From here to ATM.

IEEE Spectrum, 31(6):20–24, June 1994. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Miller:1994:VSI

[Mil94b] B. Miller. Vital signs of identity [biometrics]. IEEE Spectrum, 31(2):22–30, February 1994. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Miller:1998:SFM

[Mil98a]  B. Miller.  Satellites free

the mobile phone. IEEE Spectrum, 35(3):26–35, March 1998. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Miller:1998:AT

[Mil98b] P. Miller. Art technology.

IEEE Spectrum, 35(7):19, July 1998. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Miller:1998:ECB

[Mil98c] P. Miller. The engineer as catalyst: Billy Kluver on working with artists. IEEE Spectrum, 35(7):20–29, July 1998. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Miller:1998:TAS

[Mil98d] P. Miller. Technology for art’s sake. IEEE Spectrum, 35 (7):30–37, July 1998. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Miller:1998:WSa

[Mil98e] P. Miller. Web sights. IEEE Spectrum, 35(2):81, February 1998. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Miller:1998:WSb

[Mil98f] P. Miller. Web sights. IEEE Spectrum, 35(6):94–95, June 1998. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Miller:1998:WSc

[Mil98g] P. Miller. Web sights. IEEE Spectrum, 35(10):80, October 1998. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Miller:1999:CE

[Mil99a] P. Miller. Consumer electronics. IEEE Spectrum, 36(1): 41–45, January 1999. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Miller:1999:WS

[Mil99b] P. Miller. Web sights. IEEE Spectrum, 36(11):84, November 1999. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Mirotznik:1995:HCG

[Mir95] Mark Mirotznik. How to chart and graph like a pro. IEEE Spectrum, 32(12):58– ??, December 1995. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Mirotznik:1996:AES

[Mir96a] Mark S. Mirotznik. Affordable electromagnetics solver [software reviews]. IEEE Spectrum, 33(10):63–64, October 1996. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Mirotznik:1996:TSM

[Mir96b] Mark S. Mirotznik. Tools for schools [mathematical software packages]. IEEE Spectrum, 33(9):41–45, September 1996. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Mirotznik:1996:TMP

[Mir96c] Mark S. Mirotznik. Translating Matlab programs into C code [software reviews]. IEEE Spectrum, 33(2):63–64, February 1996. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Mirotznik:1997:MCT

[Mir97] Mark S. Mirotznik. Matlabto-C++ time-saver. IEEE Spectrum, 34(4):14–15, April 1997. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Mirotznik:1998:CEA

[Mir98]  Mark Mirotznik.  Confor-

mal EM analysis [software reviews]. IEEE Spectrum, 35 (4):84–85, April 1998. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Mirotznik:1999:BME

[Mir99a] Mark S. Mirotznik. Bringing Maxwell’s equations to heel [software reviews]. IEEE Spectrum, 36(8):82–84, August 1999. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Mirotznik:1999:SR

[Mir99b] Mark S. Mirotznik. Software reviews. IEEE Spectrum, 36 (2):64, February 1999. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Mori:1990:KKW

[MK90] K. Mori and T. Kawada. From kana to kanji: word processing in Japan. IEEE Spectrum, 27 (8):46–48, August 1990. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Morgan:1998:BTB

[MM98] M. G. Morgan and M. G. Morgan. Bigger than ‘the bomb’ [books]. IEEE Spectrum, 35 (8):8–10, August 1998. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Morgan:1990:EFJ

[MN90] M. G. Morgan and I. Nair.

Electromagnetic fields: the jury’s still out. 3. Managing the risks. IEEE Spectrum, 27 (8):32–35, August 1990. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Mogal:1993:WM

[Mog93] J. Mogal. Workstation metrics. IEEE Spectrum, 30 (5):50–52, May 1993. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Mora:1991:SNL

[Mor91] J. Mora. A streetcar named light rail. IEEE Spectrum, 28 (2):54–56, February 1991. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Morgan:1996:MPP

[Mor96]  M. G. Morgan.  Misdi-

rected policy prescription. IEEE Spectrum, 33(4):11– 82T4, April 1996. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Morgan:1997:SBG

[Mor97a] M. G. Morgan. The science behind global change. IEEE Spectrum, 34(10):14, October 1997. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Morgan:1997:S

[Mor97b] M. G. Morgan. Speakout. IEEE Spectrum, 34(12):59– 60, December 1997. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Morris:1997:CPT

[Mor97c] S. E. Morris. Crime and prevention: a Treasury viewpoint. IEEE Spectrum, 34(2): 38–39, February 1997. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Morse:1997:ME

[Mor97d] W. Morse. Medical electronics. IEEE Spectrum, 34(1): 99–102, January 1997. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Morgan:1998:EWS

[Mor98a] G. Morgan. An EE who swayed the world: review of Endless Frontier: Vannevar Bush, Engineer of the American Century. Zachary, G. Pascal, The Free Press, New York, 1997, 518 pp., $32.50. IEEE Spectrum, 35(3):10–12, March 1998. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Morreale:1998:AMM

[Mor98b]  P. Morreale.  Agents on

the move [mobile software agents]. IEEE Spectrum, 35 (4):34–41, April 1998. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Morgan:1999:BRP

[Mor99a] M. G. Morgan. Book review:

Pasteur’s Quadrant: Basic Science And Technological Innovation. IEEE Spectrum, 36 (1):12–13, January 1999. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Morgan:1999:NSP

[Mor99b] M. G. Morgan. A new science policy classic [book reviews]. IEEE Spectrum, 36(6):16–22, June 1999. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Moskun:1996:VIS

[Mos96] J. Moskun. Viewpoint: Improving software quality requires a ‘check early, check often’ approach. IEEE Spectrum, 33(1):63, January 1996. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Mosher:1997:GPB

[Mos97] D. E. Mosher. The grand plans [ballistic missile defense]. IEEE Spectrum, 34(9): 28–39, September 1997. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Mirotznik:1997:HCE

[MP97] Mark S. Mirotznik and D. Prather. How to choose electromagnetic software. IEEE Spectrum, 34(12):53–58, December 1997. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Maunder:1992:TT

[MT92] C. M. Maunder and R. E. Tulloss. Testability on TAP. IEEE Spectrum, 29(2):34– 37, February 1992. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Morgan:1996:NRN

[MT96] M. G. Morgan and S. Talukdar. Nurturing R&D in the new electric power regime. IEEE Spectrum, 33(7):32–33, July 1996. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Munn:1996:MTS

[Mun96] W. Munn. More than significant others-much more [books]. IEEE Spectrum, 33 (12):10–13, December 1996.

CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Murphy:1990:ADT

[Mur90a] E. E. Murphy. Applications ’90: design tools. IEEE Spectrum, 27(2):34–36, February 1990. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Murphy:1990:RCD

[Mur90b] E. E. Murphy. Reconciling conflicting design-automation standards. IEEE Spectrum, 27 (3):44–45, March 1990. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Murphy:1990:SRA

[Mur90c] E. E. Murphy. Software R&D:

from an art to a science. IEEE Spectrum, 27(10):44–46, October 1990. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Murphy:1990:TT

[Mur90d] E. E. Murphy. Technology ’90: transportation. IEEE Spectrum, 27(1):47–48, January 1990. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Murphy:1990:T

[Mur90e] E. E. Murphy. Transportation. IEEE Spectrum, 27 (6):26–28, June 1990. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Murphy:1996:TPP

[Mur96] J. Murphy. Ten points to ponder in picking an oscilloscope. IEEE Spectrum, 33(7):69–73, July 1996. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Murphy:1990:TSS

[MV90] E. E. Murphy and J. Voelcker. Technology ’90: systems software. IEEE Spectrum, 27 (1):38–40, January 1990. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Moore:1996:VV

[MVJ96] Carl E. Moore, Alfred Von Smolinski, and Bruno Jaselskis. Volta as voltmeter. IEEE Spectrum, 33(3):38–41, March 1996. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Morgan:1996:SDU

[MW96] M. G. Morgan and R. M. White. Speakout, dismantling the U.S. technology infrastructure. IEEE Spectrum, 33(1): 22–24, January 1996. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Magin:1997:MMR

[MWP97] R. L. Magin, A. G. Webb, and T. L. Peck. Miniature magnetic resonance machines.

IEEE Spectrum, 34(10):51– 61, October 1997. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Nahin:1991:BLT

[Nah91] Paul J. Nahin. Behind the Laplace transform. IEEE Spectrum, 28(3):60, March 1991. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Nahin:1992:MGU

[Nah92] Paul J. Nahin. Maxwell’s grand unification. IEEE Spectrum, 29(3):45, March 1992. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Natarajan:1995:VST

[Nat95] K. Natarajan. Video servers take root. IEEE Spectrum, 32 (4):66–69, April 1995. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Ng:1992:MFC

[Ng92] R. Ng. Memory — fast computer memories. IEEE Spectrum, 29(10):36–39, October 1992. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Nagle:1998:HWE

[NGOS98] H. T. Nagle, R. Gutierrez-Osuna, and S. S. Schiffman. The how and why of electronic noses. IEEE Spectrum, 35(9):

22–31, September 1998. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Ninomiya:1995:JSD

[Nin95] Y. Ninomiya. The Japanese scene [digital HDTV]. IEEE Spectrum, 32(4):54–57, April 1995. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Nair:1990:EFJ

[NM90] I. Nair and G. Morgan. Electromagnetic fields: the jury’s still out. 1. Biological effects. IEEE Spectrum, 27(8): 23–27, August 1990. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Normann:1996:CIB

[NMGW96] R. A. Normann, E. M. Maynard, K. S. Guillory, and D. J. Warren. Cortical implants for the blind. IEEE Spectrum, 33(5):54–59, May 1996. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Noah:1997:RAT

[Noa97] J. O. Noah. A rational approach to testing MPEG-2. IEEE Spectrum, 34(5):67–72, May 1997. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Norris:1995:BSW

[Nor95] G. Norris. Boeing’s seventh wonder. IEEE Spectrum, 32(10):20–23, October 1995. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Norman:1999:BRT

[Nor99] D. A. Norman. Backing the right technical horse. IEEE Spectrum, 36(2):57–61, February 1999. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Nickerson:1990:TMC

[NP90] R. S. Nickerson and R. W.

Pew. Toward more compatible human-computer interfaces. IEEE Spectrum, 27 (7):40–43, July 1990. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Oberg:1995:RSP

[Obe95] J. Oberg. Russia space program. IEEE Spectrum, 32(12): 18–35, December 1995. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Oberg:1997:FF

[Obe97] J. Oberg. Faults failures.

IEEE Spectrum, 34(8):17, August 1997. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Oberg:1998:SML

[Obe98]  J. Oberg.  Shuttle–Mir’s

lessons for the International Space Station. IEEE Spectrum, 35(6):28–37, June 1998. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Oberc:1999:SD

[Obe99a] J. Oberc. Spying for dummies. IEEE Spectrum, 36(11): 62–69, November 1999. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Oberg:1999:MAN

[Obe99b] J. Oberg. Missiles for all: the new global threat. IEEE Spectrum, 36(3):20–28, March 1999. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Oberg:1999:URS

[Obe99c] J. Oberg. The U.S.–Russian space relationship: symbolism at any cost? IEEE Spectrum, 36(7):74–81, July 1999. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Oberg:1999:WMP

[Obe99d] J. Oberg. Why the Mars probe went off course [accident investigation]. IEEE Spectrum, 36(12):34–39, December 1999. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Oliphant:1999:MPM

[Oli99] M. W. Oliphant. The mobile phone meets the Internet. IEEE Spectrum, 36(8): 20–28, August 1999. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Overly:1997:NWS

[Ove97]  M. R. Overly.  Negotiat-

ing a Web site agreement. IEEE Spectrum, 34(5):60–66, May 1997. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Perry:1991:ATCb

[PA91] Tekla S. Perry and J. A.

Adam. Air traffic control: expert observers define top priorities to ensure safe and accessible skies. IEEE Spectrum, 28(2):33–36, February 1991. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Perry:1992:EMMa

[PA92] Tekla S. Perry and J. A.

Adam. Electronic mail — Email: pervasive and persuasive. IEEE Spectrum, 29(10): 22–23, October 1992. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Proctor:1997:OAC

[PA97] F. M. Proctor and J. S. Albus.

Open architecture controllers. IEEE Spectrum, 34(6):60–64, June 1997. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Packer:1999:SRB

[Pac99] R. Packer. Software review — breezy DSP designs. IEEE Spectrum, 36(6):95–97, June 1999. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Padgett:1998:SR

[Pad98] W. T. Padgett. Software reviews. IEEE Spectrum, 35 (8):75, August 1998. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Panofsky:1997:DDE

[Pan97] W. K. H. Panofsky. Degrees of defense effectiveness. IEEE Spectrum, 34(11):74– 75, November 1997. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Partridge:1996:VSA

[Par96]  C. Partridge.  Viewpoint:

Staying ahead of the elephants. IEEE Spectrum, 33 (1):36–37, January 1996. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Paradiso:1997:EMN

[Par97] J. A. Paradiso. Electronic music: new ways to play. IEEE Spectrum, 34(12):18– 30, December 1997. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Passino:1995:ICA

[Pas95] K. M. Passino. Intelligent control for autonomous systems. IEEE Spectrum, 32(6):55–62, June 1995. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Poggio:1996:LS

[PB96] T. Poggio and D. Beymer.

Learning to see. IEEE Spectrum, 33(5):60–69, May 1996. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Paul:1998:SCS

[PC98] W. Paul and M. Chen. Superconducting control for surge currents. IEEE Spectrum, 35 (5):49–54, May 1998. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Peleeger:1997:LTL

[Pel97] S. L. Peleeger. Love them or leave them? IEEE Spectrum, 34(11):14–15, November 1997. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Pensak:1999:QVH

[Pen99] D. Pensak. Quality visualization for the hierarchical data format [software reviews]. IEEE Spectrum, 36 (8):80, August 1999. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Perry:1990:BEA

[Per90a]  Tekla S. Perry.  Biome-

chanically engineered athletes. IEEE Spectrum, 27 (4):43–44, April 1990. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Perry:1990:CE

[Per90b] Tekla S. Perry. Consumer electronics. IEEE Spectrum, 27 (6):33–34, June 1990. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Perry:1990:TPT

[Per90c] Tekla S. Perry. Teamwork plus technology cuts development time. IEEE Spectrum, 27(10): 61–67, October 1990. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Perry:1990:TCP

[Per90d] Tekla S. Perry. Telephone challenges: a plethora of services. IEEE Spectrum, 27 (7):25–28, July 1990. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Perry:1990:WEM

[Per90e] Tekla S. Perry. When East meets West (management problems). IEEE Spectrum, 27 (8):53–55, August 1990. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Perry:1991:ATCa

[Per91] Tekla S. Perry. Air traffic control: a system-wide upgrade races to replace twodecade-old equipment. IEEE Spectrum, 28(2):22–27, February 1991. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Perry:1992:EMMb

[Per92a] Tekla S. Perry. Electronic mail — E-mail at work. IEEE Spectrum, 29(10):24–28, October 1992. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Perry:1992:EMF

[Per92b] Tekla S. Perry. Electronic mail — forces for social change. IEEE Spectrum, 29(10):30– 32, October 1992. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Perry:1992:EP

[Per92c] Tekla S. Perry. Engineers in profile. IEEE Spectrum, 29(12):29–36, December 1992. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Perry:1993:CEI

[Per93a] Tekla S. Perry. Cleaning (electronics industry safety). IEEE Spectrum, 30(2):20–26, February 1993. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Perry:1993:EMW

[Per93b] Tekla S. Perry. Environment — modeling the world’s climate. IEEE Spectrum, 30 (7):33–37, July 1993. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Perry:1994:CE

[Per94a] Tekla S. Perry. Consumer electronics. IEEE Spectrum, 31 (1):30–34, January 1994. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Perry:1994:EBE

[Per94b]  Tekla S. Perry.  EEs in

the boardroom [electrical engineers]. IEEE Spectrum, 31 (4):20–24, April 1994. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Perry:1994:GR

[Per94c] Tekla S. Perry. ‘Green’ refrigerators. IEEE Spectrum, 31 (8):25–30, August 1994. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Perry:1994:MHJ

[Per94d] Tekla S. Perry. Marcian E.

Hoff, Jr. IEEE Spectrum, 31 (2):46–49, February 1994. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Perry:1994:ME

[Per94e] Tekla S. Perry. The media event. IEEE Spectrum, 31 (1):21, January 1994. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Perry:1994:TVM

[Per94f] Tekla S. Perry. Today’s view of magnetic fields. IEEE Spectrum, 31(12):14–23, December 1994. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Perry:1995:CE

[Per95a] Tekla S. Perry. Consumer electronics. IEEE Spectrum, 32 (1):40–43, January 1995. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Perry:1995:EER

[Per95b] Tekla S. Perry. The environment [energy/resource conservation]. IEEE Spectrum, 32 (1):60–61, January 1995. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Perry:1995:HND

[Per95c] Tekla S. Perry. HDTV and the new digital television. IEEE Spectrum, 32(4):34–35, April 1995. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Perry:1995:LZF

[Per95d] Tekla S. Perry. Lotfi A. Zadeh [fuzzy logic inventor biography]. IEEE Spectrum, 32 (6):32–35, June 1995. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Perry:1995:MGC

[Per95e] Tekla S. Perry. M. George Craford [biography]. IEEE Spectrum, 32(2):52–55, February 1995. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Perry:1996:TTI

[Per96] Tekla S. Perry. The trials and travails of interactive TV. IEEE Spectrum, 33 (4):22–28, April 1996. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Perry:1997:CE

[Per97a] Tekla S. Perry. Consumer electronics. IEEE Spectrum, 34 (1):43–48, January 1997. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Perry:1997:EMT

[Per97b] Tekla S. Perry. Electronic money: Toward a virtual wallet. IEEE Spectrum, 34(2): 18–19, February 1997. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Perry:1997:SFA

[Per97c] Tekla S. Perry. In search of the future of air traffic control. IEEE Spectrum, 34(8): 18–35, August 1997. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Perry:1998:DPE

[Per98a] T. Perry. Donald O. Pederson [electronic engineering biography]. IEEE Spectrum, 35 (6):22–27, June 1998. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Perry:1998:CET

[Per98b] Tekla S. Perry. Consumer electronics [technology 1998 analysis and forecast]. IEEE Spectrum, 35(1):57–61, January 1998. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Perry:1999:CY

[Per99a] Tekla S. Perry. Checking up on y2k. IEEE Spectrum, 36 (7):61–70, July 1999. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Perry:1999:HS

[Per99b] Tekla S. Perry. Henry Samueli. IEEE Spectrum, 36(9):70– 75, September 1999. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Pettersson:1995:ICC

[Pet95] G. Pettersson. ISDN: from custom to commodity service. IEEE Spectrum, 32(6):26–31, June 1995. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Piedra:1996:DSP

[PF96] R. M. Piedra and A. Fritsch. Digital signal processing comes of age. IEEE Spectrum, 33 (5):70–74, May 1996. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Pfleeger:1992:MSR

[Pfl92] S. L. Pfleeger. Measuring software reliability. IEEE Spectrum, 29(8):56–60, August 1992. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Pfleeger:1996:MSD

[Pfl96] S. L. Pfleeger. Managing software development [book review]. IEEE Spectrum, 33 (8):11, August 1996. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Perry:1996:DPE

[PG96] Tekla S. Perry and L. Geppert. Do portable electronics endanger flight? The evidence mounts. IEEE Spectrum, 33(9):26–33, September 1996. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Phillips:1998:TTA

[Phi98] D. Phillips. Transportation [technology 1998 analysis and forecast]. IEEE Spectrum, 35 (1):84–89, January 1998. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Pickover:1995:EIR

[Pic95] C. A. Pickover. Exploring infinite realms. IEEE Spectrum, 32(11):58–60, November 1995. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Pixley:1996:DCF

[Pix96] C. Pixley. Debugging a cache flush unit. IEEE Spectrum, 33(6):63, June 1996. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Picraux:1998:BSI

[PM98] S. T. Picraux and P. J.

McWhorter. The broad sweep of integrated microsystems. IEEE Spectrum, 35(12):24– 33, December 1998. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Powers:1996:PW

[Pow96] C. B. Powers. Preparing for the worst. IEEE Spectrum, 33(12):49–54, December 1996. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Pennino:1993:DM

[PP93] T. P. Pennino and J. Potechin. Design for manufacture. IEEE Spectrum, 30(9):51–53, September 1993. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Persoz:1992:CEP

[PR92] H. Persoz and J. Remondeulaz. Consolidating European power. IEEE Spectrum, 29(10):62–65, October 1992. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Pradhan:1994:TLP

[Pra94] B. D. Pradhan. Telecommunications: at least a phone in every village. IEEE Spectrum, 31(3):43–47, March 1994. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Prince:1994:MFL

[Pri94] B. Prince. Memory in the fast lane. IEEE Spectrum, 31(2): 38–41, February 1994. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Prince:1992:IGV

[PS92] B. Prince and R. H. W.

Salters. ICs going on a 3-V diet. IEEE Spectrum, 29 (5):22–25, May 1992. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Peterson:1998:PER

[PS98] J. Peterson and M. Sharp. Pan European R&D: a perspective. IEEE Spectrum, 35(10): 54–59, October 1998. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Radford:1996:SSD

[Rad96] D. Radford. Spread spectrum data leap through AC power wiring. IEEE Spectrum, 33(11):48–53, November 1996. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Raji:1994:SNC

[Raj94] R. S. Raji. Smart networks for control. IEEE Spectrum, 31(6):49–55, June 1994. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Ramamoorty:1994:EPG

[Ram94] M. Ramamoorty. Electric power: a government priority. IEEE Spectrum, 31(3):

51–52, March 1994. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Rauch:1993:LRE

[Rau93] J. G. Rauch. The law on reverse engineering. IEEE Spectrum, 30(8):47–48, August 1993. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Rush:1990:CHS

[RDK90] K. Rush, S. Draving, and J. Kerley. Characterizing high-speed oscilloscopes. IEEE Spectrum, 27(9):38– 39, September 1990. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Rechtin:1992:ASA

[Rec92] E. Rechtin. The art of systems architecting. IEEE Spectrum, 29(10):66–69, October 1992. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Rechtin:1997:SCS

[Rec97] E. Rechtin. The synthesis of complex systems. IEEE Spectrum, 34(7):50–55, July 1997. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Reinhart:1992:B

[Rei92a] E. E. Reinhart. Broadcasting. IEEE Spectrum, 29(2):

24–26, February 1992. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Reinhart:1992:MC

[Rei92b] F. E. Reinhart. Mobile communications. IEEE Spectrum, 29(2):27–29, February 1992. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Riezenman:1992:CCE

[RGG^{+}92] M. J. Riezenman, J. S.

Gansler, J. J. Guarrera, H. Gracon, G. L. Brown, D. L. Schilling, E. B. Woollen, R. J. Polutchoko, and R. F. Rogers. The courage to convert (engineers). IEEE Spectrum, 29(12):62–68, December 1992. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Rouse:1990:CAF

[RGH90] W. B. Rouse, N. D. Geddes, and J. M. Hammer. Computer-aided fighter pilots. IEEE Spectrum, 27(3): 38–41, March 1990. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Riordan:1997:OPJ

[RH97] Michael Riordan and Lillian Hoddeson. The origins of the pn junction. IEEE Spectrum, 34(6):46–51, June 1997. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Rhodes:1996:DLA

[Rho96] D. L. Rhodes. A design language for analog circuits. IEEE Spectrum, 33(10):43– 48, October 1996. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Richards:1992:DHT

[Ric92] S. Richards. Diversity in the high-tech workplace: making engineers feel at home. IEEE Spectrum, 29(6):29–30, June 1992. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Riddle:1994:MPT

[Rid94]  A. Riddle.  Mathematical

power tools. IEEE Spectrum, 31(11):35–38, November 1994. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Riddle:1998:TIS

[Rid98] A. N. Riddle. A truly interactive spice tool. IEEE Spectrum, 35(4):84, April 1998. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Riezenman:1991:GLR

[Rie91a] M. Riezenman. Gulf legacy: revising the script after Patriot. IEEE Spectrum, 28(9): 49–52, September 1991. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Riezenman:1991:WCC

[Rie91b] M. J. Riezenman. Wanlass’s CMOS circuit. IEEE Spectrum, 28(5):44, May 1991. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Riezenman:1992:EV

[Rie92a] M. J. Riezenman. Electric vehicles. IEEE Spectrum, 29 (11):18–21, November 1992. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Riezenman:1992:EVP

[Rie92b] M. J. Riezenman. Electric vehicles — pursuing efficiency. IEEE Spectrum, 29(11):22– 24, November 1992. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Riezenman:1992:EVA

[Rie92c] M. J. Riezenman. Electric vehicles-architecting the system. IEEE Spectrum, 29(11): 94–96, November 1992. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Riezenman:1992:EVG

[Rie92d] M. J. Riezenman. Electric vehicles-the great battery barrier. IEEE Spectrum, 29(11): 97–101, November 1992. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Riezenman:1992:TTM

[Rie92e] M. J. Riezenman. Technology 1992 — test and measurement. IEEE Spectrum, 29(1): 45–46, January 1992. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Riezenman:1993:ESE

[Rie93]  M. J. Riezenman.  Engi-

neering software — efficient data acquisition. IEEE Spectrum, 30(11):76–84, November 1993. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Riezenman:1994:TM

[Rie94] M. J. Riezenman. Test and measurement. IEEE Spectrum, 31(1):54–57, January 1994. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Riezenman:1995:SBB

[Rie95a] M. J. Riezenman. The search for better batteries. IEEE Spectrum, 32(5):51–56, May 1995. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Riezenman:1995:TM

[Rie95b] M. J. Riezenman. Test and measurement. IEEE Spectrum, 32(1):52–55, January 1995. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Riezenman:1996:ETTa

[Rie96a] M. J. Riezenman. EEs’ tools toys. IEEE Spectrum, 33(1): 119–121, January 1996. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Riezenman:1996:ETTb

[Rie96b] M. J. Riezenman. EEs’ tools toys. IEEE Spectrum, 33 (3):69–70, March 1996. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Riezenman:1996:ETTc

[Rie96c] M. J. Riezenman. Ees’ tools toys. IEEE Spectrum, 33(11): 81–84, November 1996. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Riezenman:1996:TM

[Rie96d] M. J. Riezenman. Tests and measurement. IEEE Spectrum, 33(1):65–69, January 1996. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Riezenman:1997:ETTa

[Rie97a] M. J. Riezenman. Ees’ tools toys. IEEE Spectrum, 34(1): 106–110, January 1997. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Riezenman:1997:ETTb

[Rie97b] M. J. Riezenman. EEs’ tools toys. IEEE Spectrum, 34 (3):75–76, March 1997. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Riezenman:1997:ETTc

[Rie97c] M. J. Riezenman. EEs’ tools toys. IEEE Spectrum, 34 (4):53–56, April 1997. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Riezenman:1997:ETTd

[Rie97d] M. J. Riezenman. EEs’ tools toys. IEEE Spectrum, 34 (5):73–76, May 1997. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Riezenman:1997:ETTe

[Rie97e] M. J. Riezenman. EEs’ tools toys. IEEE Spectrum, 34 (6):73–75, June 1997. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Riezenman:1997:ETTf

[Rie97f]  M. J. Riezenman.  EEs’

tools toys. IEEE Spectrum, 34(7):76, July 1997. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Riezenman:1997:ETTg

[Rie97g] M. J. Riezenman. Ees’ tools toys. IEEE Spectrum, 34(8): 78–80, August 1997. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Riezenman:1997:ETTh

[Rie97h] M. J. Riezenman. EEs’ tools toys. IEEE Spectrum, 34(12): 61–64, December 1997. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Riezenman:1997:EWa

[Rie97i]  M. J. Riezenman.  EV

watch. IEEE Spectrum, 34 (3):77–78, March 1997. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Riezenman:1997:EWb

[Rie97j] M. J. Riezenman. EV watch.

IEEE Spectrum, 34(6):68, June 1997. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Riezenman:1997:EWc

[Rie97k]  M. J. Riezenman.  EV

watch. IEEE Spectrum, 34 (9):24, September 1997. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Riezenman:1997:EWd

[Rie97l] M. J. Riezenman. EV watch.

IEEE Spectrum, 34(12):68– 70, December 1997. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Riezenman:1998:ETT

[Rie98a] M. J. Riezenman. EEs’ tools toys. IEEE Spectrum, 35(1): 106–107, January 1998. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Riezenman:1998:EEF

[Rie98b] M. J. Riezenman. Engineering the EV future. IEEE Spectrum, 35(11):18–20, November 1998. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Riezenman:1998:EWa

[Rie98c] M. J. Riezenman. EV watch. IEEE Spectrum, 35(6):93, June 1998. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Riezenman:1998:EWb

[Rie98d] M. J. Riezenman. EV watch. IEEE Spectrum, 35(9):19– 20, September 1998. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Riezenmann:1998:CTA

[Rie98e] M. J. Riezenmann. Communications [technology 1998 analysis and forecast]. IEEE Spectrum, 35(1):29–36, January 1998. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Riezenman:1999:EWa

[Rie99a] M. J. Riezenman. EV watch.

IEEE Spectrum, 36(6):89–90, June 1999. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Riezenman:1999:EWb

[Rie99b]  M. J. Riezenman.  EV

watch. IEEE Spectrum, 36 (9):100, September 1999. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Riezenman:1999:EWc

[Rie99c] M. J. Riezenman. EV watch.

IEEE Spectrum, 36(12):112– 114, December 1999. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Rivera:1995:TA

[Riv95] C. Rivera. Technology for all. IEEE Spectrum, 32(12): 8–??, December 1995. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Ralston:1992:CS

[RKGB92] R. W. Ralston, M. A. Kastner, W. J. Gallagher, and B. Batlogg. Cooperating on superconductivity. IEEE Spectrum, 29(8):50–55, August 1992. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Rush:1993:TWC

[RLH93] K. Rush and T. Lucero-Hall. Two ways to catch a wave (oscilloscopes). IEEE Spectrum, 30(2):38–41, February 1993. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Rachowitz:1991:GEW

[RMAA91] B. I. Rachowitz, R. K. Maue, N. P. Angrisano, and B. Abramson. A guide to engineering workstations: using workstations efficiently. IEEE Spectrum, 28(4):38–40, April 1991. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Rahman:1997:HCC

[RN97] M. M. Rahman and M. Nassi. High capacity cable’s role in once and future grids. IEEE Spectrum, 34(7):31–35, July 1997. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Robison:1992:EST

[Rob92] R. W. Robison. Engineering software — tools for embedding DSP. IEEE Spectrum, 29(11):81–82, November 1992. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Robison:1995:TTK

[Rob95] R. Robison. The twin towers of Kuala Lumpur. IEEE Spectrum, 32(10):44–47, October 1995. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Robinson:1997:BRF

[Rob97] W. A. Robinson. Book review:

Forecasting the Next Century’s Weather [books]. IEEE Spectrum, 34(1):10–14, January 1997. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Robinson:1999:GYS

[Rob99] W. A. Robinson. Greenhouse: The 200-Year story of global warming, Christianson, Gale E., Walker & Co., New York, 1999, 303 pp., ISBN 0-8027-1346-7, $25 [book review]. IEEE Spectrum, 36(11): 16–17, November 1999. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Rockwell:1997:ISS

[Roc97] R. Rockwell. An infrastructure for social software. IEEE Spectrum, 34(3):26–31, March 1997. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Rodrigo:1998:RLH

[Rod98] G. C. Rodrigo. A revisionist look at hits and flops. IEEE Spectrum, 35(9):12, September 1998. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Rodrigo:1999:SCE

[Rod99] G. C. Rodrigo. Software as Capitol: an economic perspective on software engineering [book review]. IEEE Spectrum, 36(4):14–16, April 1999. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Roehle:1997:CDF

[Roe97] B. Roehle. Channeling the data flood. IEEE Spectrum, 34 (3):32–38, March 1997. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Rosenblatt:1991:TDC

[Ros91a]  A. Rosenblatt.  Technol-

ogy 1991: data communications. IEEE Spectrum, 28(1): 48–51, January 1991. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Rosenblatt:1991:TPW

[Ros91b] A. Rosenblatt. Technology 1991: PCs and workstations. IEEE Spectrum, 28(1): 32–35, January 1991. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Rosenblatt:1991:WAH

[Ros91c] A. Rosenblatt. Who’s ahead in hi-tech? IEEE Spectrum, 28(4):22–27, April 1991. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Rosenblatt:1992:EWW

[Ros92a] A. Rosenblatt. Engineering workstations — what the vendors wrought. IEEE Spectrum, 29(4):38–42, April 1992. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Rosenblatt:1992:TPW

[Ros92b] A. Rosenblatt. Technology 1992 — PCs and workstations. IEEE Spectrum, 29(1): 27–29, January 1992. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Rosenblatt:1996:CMT

[Ros96] A. Rosenblatt. Catching up with modern times [modernisation of Brazil’s infrastructure]. IEEE Spectrum, 33 (6):22–28, June 1996. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Rosengren:1998:MET

[Ros98] J. Rosengren. Medical electronics [technology 1998 analysis and forecast]. IEEE Spectrum, 35(1):62–66, January 1998. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Rosenblatt:1991:FCE

[RPD^{+}91] A. Rosenblatt, Tekla S. Perry, S. M. Dambrot, P. Gwynne, G. Watson, and K. Fitzgerald. Formula for competitiveness (electronics). IEEE Spectrum, 28(6):49–62, June 1991. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Riezenman:1999:CTA

[RS99a] M. J. Riezenman and W. Sweet. Communications [technology 1999 analysis and forecast]. IEEE Spectrum, 36(1):29– 34, January 1999. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Ronald:1999:S

[RS99b] E. Ronald and M. Sipper. Speakout. IEEE Spectrum, 36(9):28–31, September 1999. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Rogowitz:1998:DVE

[RT98] B. E. Rogowitz and L. A. Treinish. Data visualization: the end of the rainbow. IEEE Spectrum, 35(12): 52–59, December 1998. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Redl:1997:PEP

[RTD97] R. Redl, P. Tenti, and J. Daan van Wyk. Power electronics’ polluting effects. IEEE Spectrum, 34(5):32–39, May 1997. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Reinhart:1992:WLA

[RTHM92] E. E. Reinhart, R. M. Taylor, A. O. Heyward, and J. Miller. WARC’s last act? IEEE Spectrum, 29(2):20–23, February 1992. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Rudnick:1996:PER

[Rud96] H. Rudnick. Pioneering electricity reform in South America. IEEE Spectrum, 33(8): 38–44, August 1996. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Ruiu:1994:TAS

[Rui94] B. Ruiu. Testing ATM systems. IEEE Spectrum, 31 (6):25–27, June 1994. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Runyon:1999:TBC

[Run99]  S. Runyon.  Testing big

chips becomes an internal affair. IEEE Spectrum, 36 (4):49–55, April 1999. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Reizenman:1998:ERA

[RW98] M. J. Reizenman and V. Wouk.

EVs: The road ahead. IEEE Spectrum, 35(12):42– 51, December 1998. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Reddy:1991:CED

[RWC91] R. Reddy, R. T. Wood, and K. J. Cleetus. Concurrent engineering: The DARPA initiative: encouraging new industrial practices. IEEE Spectrum, 28(7):26–27, July 1991. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Ryan:1998:SUM

[Rya98] L. Ryan. Software usage metrics for real-world software testing. IEEE Spectrum, 35 (4):64–68, April 1998. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Ryder:1991:MM

[Ryd91] J. D. Ryder. Magnetism without magnets. IEEE Spectrum, 28(10):28, October 1991. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Saito:1990:JTC

[Sai90] T. Saito. Japan telecommunications at the crossroads. IEEE Spectrum, 27(11):126– 128, November 1990. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Salters:1992:MFD

[Sal92] R. H. W. Salters. Memory — fast DRAMs for sharper TV. IEEE Spectrum, 29(10):

40–42, October 1992. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Santori:1990:IIR

[San90]  M. Santori.  An instru-

ment that isn’t really (Laboratory Virtual Instrument Engineering Workbench). IEEE Spectrum, 27(8):36–39, August 1990. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Santos:1996:PCH

[San96] A. Santos. Personal computers are hot. IEEE Spectrum, 33(7):34–39, July 1996. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Sarch:1993:TDC

[Sar93] R. Sarch. Technology 1993 — data communications. IEEE Spectrum, 30(1):42–45, January 1993. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Sawyer:1995:BC

[Saw95] P. B. Sawyer. Basic chemistry. IEEE Spectrum, 32(12): 8–??, December 1995. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Sweet:1998:CRL

[SB98a] W. Sweet and E. A. Bretz.

Corporate R&D leaders reflect on their mutual European future. IEEE Spectrum, 35(10): 60–68, October 1998. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Sweet:1998:OER

[SB98b] W. Sweet and E. A. Bretz. One Europe, raring to go. IEEE Spectrum, 35(10):19– 22, October 1998. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Singh:1999:IPS

[SB99a] V. Singh and E. A. Bretz. India’s power struggles. IEEE Spectrum, 36(11):48–56, November 1999. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Sweet:1999:SVI

[SB99b] W. Sweet and E. A. Bretz. Site visits illustrate advanced grid design. IEEE Spectrum, 36(12):64–68, December 1999. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Sweet:1999:TCF

[SB99c] W. Sweet and E. A. Bretz. Toward carbon free energy. IEEE Spectrum, 36(11):28– 33, November 1999. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Santo:1990:TSS

[SC90] B. Santo and K. T. Chen.

Technology ’90: solid state (semiconductor storage). IEEE Spectrum, 27(1):41–43, January 1990. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Schmalzel:1991:DH

[Sch91] J. L. Schmalzel. Data handling. IEEE Spectrum, 28(11): 38–42, November 1991. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Schaller:1997:MLP

[Sch97]  R. R. Schaller.  Moore’s

Law: past, present and future. IEEE Spectrum, 34(6):52–59, June 1997. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Socarras:1991:AXT

[SCS91] A. E. Socarras, R. S. Cooper, and W. F. Stoneycypher. Anatomy of an X terminal. IEEE Spectrum, 28(3): 52–55, March 1991. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Serapio:1994:FRU

[SD94] M. G. Serapio and D. H. Dalton. Foreign R&D in the United States. IEEE Spectrum, 31(11):26–30, November 1994. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Seidel:1999:LC

[Sei99] R. W. Seidel. Lives of computers. IEEE Spectrum, 36 (7):13–14, July 1999. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Self:1990:DFL

[Sel90] K. Self. Designing with fuzzy logic. IEEE Spectrum, 27(11): 42–44, November 1990. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Self:1993:TSS

[Sel93] K. Self. Technology 1993 — solid state. IEEE Spectrum, 30 (1):46–49, January 1993. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Self:1996:TSa

[Sel96a] K. Self. Technically speaking. IEEE Spectrum, 33 (10):17, October 1996. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Self:1996:TSb

[Sel96b] K. Self. Technically speaking. IEEE Spectrum, 33(12): 23, December 1996. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Self:1996:TPQ

[Sel96c] K. Self. Toward a postmodern, quantum-gravitational ruckus [technically speaking]. IEEE Spectrum, 33(8):20, August 1996. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Self:1997:TSa

[Sel97a] K. Self. Technically speaking. IEEE Spectrum, 34 (4):17, April 1997. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Self:1997:TSb

[Sel97b] K. Self. Technically speaking. IEEE Spectrum, 34(6):17, June 1997. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Self:1997:TSc

[Sel97c] K. Self. Technically speaking.

IEEE Spectrum, 34(8):15, August 1997. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Self:1997:TSd

[Sel97d] K. Self. Technically speaking. IEEE Spectrum, 34(12): 17, December 1997. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Self:1997:TSN

[Sel97e] K. Self. Technically speaking: a node named Nosferatu. IEEE Spectrum, 34 (10):18, October 1997. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Self:1998:TSa

[Sel98a] K. Self. Technically speaking. IEEE Spectrum, 35(2): 17, February 1998. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Self:1998:TSb

[Sel98b] K. Self. Technically speaking. IEEE Spectrum, 35 (4):17, April 1998. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Self:1998:TSc

[Sel98c] K. Self. Technically speaking. IEEE Spectrum, 35(6):21, June 1998. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Self:1998:TSd

[Sel98d] K. Self. Technically speaking. IEEE Spectrum, 35(8):13, August 1998. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Self:1998:TSe

[Sel98e] K. Self. Technically speaking. IEEE Spectrum, 35 (10):18, October 1998. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Self:1998:TSf

[Sel98f] K. Self. Technically speaking. IEEE Spectrum, 35(12): 17, December 1998. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Self:1999:TSa

[Sel99a] K. Self. Technically speaking. IEEE Spectrum, 36(2): 17, February 1999. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Self:1999:TSb

[Sel99b] K. Self. Technically speaking. IEEE Spectrum, 36 (4):26, April 1999. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Self:1999:TSc

[Sel99c] K. Self. Technically speaking. IEEE Spectrum, 36(6):27, June 1999. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Self:1999:TSd

[Sel99d] K. Self. Technically speaking.

IEEE Spectrum, 36(8):18, August 1999. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Self:1999:TSe

[Sel99e] K. Self. Technically speaking. IEEE Spectrum, 36 (10):27, October 1999. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Sweet:1999:CCC

[SH99] W. Sweet and M. Hood. Can China consume less coal? IEEE Spectrum, 36(11):39– 47, November 1999. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Sharma:1991:GDC

[Sha91] R. L. Sharma. A guide to data communications: interconnecting LANs. IEEE Spectrum, 28(8):32–34 37, August 1991. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Shapira:1993:JLA

[Sha93] P. Shapira. Japan looks after the little guys (manufacturing industries). IEEE Spectrum, 30(9):70–73, September 1993. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Shapiro:1994:GHH

[Sha94]  J. Shapiro.  George H.

Heilmeier. IEEE Spectrum, 31(6):56–59, June 1994. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Shapiro:1995:TA

[Sha95] D. Shapiro. Technology for all. IEEE Spectrum, 32(12):

8–??, December 1995. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Shahram:1999:CST

[Sha99] M. Shahram. Calibrating simulation tools for nanometer designs. IEEE Spectrum, 36 (6):77–82, June 1999. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Shina:1991:CEN

[Shi91] S. G. Shina. Concurrent engineering: new rules for worldclass companies. IEEE Spectrum, 28(7):22–26, July 1991. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Soden:1996:IDD

[SHM96] J. M. Soden, C. F. Hawkins, and A. C. Miller. Identifying defects in deep-submicron CMOS ICs. IEEE Spectrum, 33(9):66–71, September 1996. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Singh:1996:AM

[Sin96] M. P. Singh. Agents misunderstood. IEEE Spectrum, 33 (2):10–64, February 1996. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Sirbu:1997:CDI

[Sir97] M. A. Sirbu. Credits and debits on the Internet. IEEE Spectrum, 34(2):23–29, February 1997. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Strobel:1993:PPL

[SJ93] R. A. Strobel and A. Johnson. Pocket pagers in lots of one CIM. IEEE Spectrum, 30(9): 29–32, September 1993. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Sabonnadiere:1992:ESC

[SK92a] J.-C. Sabonnadiere and A. Konrad. Engineering software — computing EM fields. IEEE Spectrum, 29(11):52– 54, November 1992. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Schwartz:1992:FLF

[SK92b] D. G. Schwartz and G. J. Klir. Fuzzy logic flowers in Japan. IEEE Spectrum, 29(7):32–35, July 1992. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Skytte:1994:ESS

[Sky94] K. Skytte. Engineering a small system. IEEE Spectrum, 31 (3):63–65, March 1994. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Stoyenko:1994:DEU

[SLHM94] A. D. Stoyenko, P. A. Laplante, R. Harrison, and T. J. Marlowe. Doubling the engineer’s utility. IEEE Spectrum, 31(12):32–39, December 1994. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Stein:1997:FCB

[SLLP97] R. M. Stein, J. Little, G. N.

Lewis, and T. A. Postol. ‘Future challenges to ballistic missile defense’ paints wrong picture [comments and reply]. IEEE Spectrum, 34(11): 70–72, November 1997. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic). See reply [LP97b].

Slovick:1996:TC

[Slo96] M. Slovick. Table of contents. IEEE Spectrum, 33 (7):68, July 1996. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Smith:1992:ESM

[Smi92a] M. J. S. Smith. Engineering software — more logic synthesis for ASICs. IEEE Spectrum, 29(11):44–48, November 1992. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Smith:1992:DHT

[Smi92b] T. J. Smith. Diversity in the high-tech workplace: preparing for responsibility. IEEE Spectrum, 29(6):31, June 1992. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Smith:1993:SSD

[Smi93] B. Smith. Six-sigma design (quality control). IEEE Spectrum, 30(9):43–47, September 1993. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Stempel:1998:NMH

[SOGC98] Robert C. Stempel, Stanford R. Ovshinsky, Paul R. Gifford, and Dennis A. Corrigan. Nickel metal hydride: ready to serve. IEEE Spectrum, 35(11):29–34, November 1998. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Soota:1994:POS

[Soo94] A. Soota. A partner on the other side of the globe. IEEE Spectrum, 31(3):34–36, March 1994. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Soule:1998:PEG

[Sou98] C. A. Soule. The practical engineer [give priority to heat sinks to reduce design time]. IEEE Spectrum, 35(2): 61–66, February 1998. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Spadaro:1997:TMT

[Spa97] J. Spadaro. Test and measurement [technology analysis and forecast]. IEEE Spectrum, 34 (1):74–78, January 1997. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Spicer:1994:BN

[Spi94] J. Spicer. Backing up the network. IEEE Spectrum, 31 (4):66–70, April 1994. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Spicer:1996:MWM

[Spi96] D. Spicer. Meeting world market requirements. IEEE Spectrum, 33(2):10, February 1996. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Schilling:1990:SSG

[SPM90] D. L. Schilling, R. L. Pickholtz, and L. B. Milstein. Spread spectrum goes commercial. IEEE Spectrum, 27 (8):40–41, August 1990. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Souders:1991:CHC

[SS91] T. M. Souders and G. N. Stenbakken. Cutting the high cost of testing. IEEE Spectrum, 28 (3):48–51, March 1991. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Schoch:1992:CP

[SS92] K. F. Schoch and H. E. Saunders. Conducting polymers. IEEE Spectrum, 29(6):52–55, June 1992. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Sabin:1996:QER

[SS96] D. D. Sabin and A. Sundaram. Quality enhances reliability [power supplies]. IEEE Spectrum, 33(2):34–41, February 1996. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Samant:1999:HRF

[SS99] A. Samant and S. Shearman. High-resolution frequency analysis with a small data record. IEEE Spectrum, 36(9):82–86, September 1999. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Simmons:1992:TBE

[SSD92] J. A. Simmons, P. A. Saillant, and S. P. Dear. Through a bat’s ear. IEEE Spectrum, 29 (3):46–48, March 1992. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Sathbye:1999:DBL

[SSH99] J. Sathbye, J. Sinton, and T. Heller. Doing better with less energy [fuel-efficient power generation]. IEEE Spectrum, 36(12):42–49, December 1999. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Schwab:1993:CE

[SSZ93] A. J. Schwab, B. Schilli, and K. Zinser. Concurrent engineering. IEEE Spectrum, 30(9):56–60, September 1993. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

StJohn:1992:GSR

[St.92] Anthony N. St. John. Grounds for signal referencing (earthing). IEEE Spectrum, 29 (6):42–45, June 1992. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Stamen:1993:SDA

[Sta93] J. P. Stamen. Structuring databases for analysis. IEEE Spectrum, 30(10):55–58, October 1993. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Sternglass:1992:FPC

[Ste92] D. Sternglass. The future is in the PC cards. IEEE Spectrum, 29(6):46–50, June 1992. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Stephenson:1993:TME

[Ste93] J. Stephenson. Technology 1993 — medical electronics. IEEE Spectrum, 30(1): 76–79, January 1993. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Stevens:1996:FEC

[Ste96] C. F. Stevens. Five equations that changed the world: the power and poetry of mathematics [books]. IEEE Spectrum, 33(3):11, March 1996. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Stern:1998:SEI

[Ste98a] L. Stern. Staying employed in interesting times (book review). IEEE Spectrum, 35 (1):8, January 1998. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Stevens:1998:DC

[Ste98b] J. Stevens. DSPs in communications. IEEE Spectrum, 35(9):39–46, September 1998. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Stern:1999:WLB

[Ste99] L. Stern. Winning and losing, big (review of Turbo capitalism: winners and losers in the global economy; Luttwak, E.; 1999). IEEE Spectrum, 36(12):10–12, December 1999. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Stix:1990:APL

[Sti90a] Gary Stix. Americans in Paris … (European Community). IEEE Spectrum, 27(6):55–56, June 1990. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Stix:1990:TDC

[Sti90b] Gary Stix. Technology ’90:

data communications. IEEE Spectrum, 27(1):35–37, January 1990. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Stix:1990:T

[Sti90c] Gary Stix. Telecommunications. IEEE Spectrum, 27 (6):25–26, June 1990. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Stone:1995:BAH

[Sto95] B. Stone. Between the Alps and a hard place [rail network modernisation]. IEEE Spectrum, 32(10):34–39, October 1995. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Stoll:1997:NLA

[Sto97] C. Stoll. Networking locally, area-wise. IEEE Spectrum, 34(6):8, June 1997. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Stone:1998:RBH

[Sto98] A. Stone. (re)butting heads over privacy. IEEE Spectrum, 35(7):10–13, July 1998. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Stoneking:1999:IME

[Sto99]  D. Stoneking.  Improving

the manufacturability of electronic designs. IEEE Spectrum, 36(6):70–76, June 1999. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Strong:1996:PWB

[Str96] S. J. Strong. Power windows. Building-integrated photovoltaics. IEEE Spectrum, 33(10):49–55, October 1996. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Suran:1990:VOS

[Sur90] J. J. Suran. A view from the other side (engineering education). IEEE Spectrum, 27(9): 52–54, September 1990. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Stahlkopf:1997:TCB

[SW97] K. E. Stahlkopf and M. R. Wilhelm. Tighter controls for busier systems [power systems]. IEEE Spectrum, 34 (4):48–52, April 1997. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Swanson:1991:SEF

[Swa91] D. G. Swanson. Simulating EM fields. IEEE Spectrum, 28(11):34–37, November 1991. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Swamidass:1993:TPM

[Swa93] P. M. Swamidass. Technology, people, and management. IEEE Spectrum, 30(9): 68–69, September 1993. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Swann:1996:TCB

[Swa96] H. Swann. A tour of the calculus [books]. IEEE Spectrum, 33(11):13–18, November 1996. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Swann:1998:MM

[Swa98] H. Swann. A mathematical messiah. IEEE Spectrum, 35(12):10–12, December 1998. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic). Review of two books: The Man Who Loved Only Numbers: Hoffman (Hyperion, P.; 1998) and My Brain is Open (Bruce, Schechter; 1998).

Sweet:1994:ESP

[Swe94a] W. Sweet. Europe’s supercollider project. IEEE Spectrum, 31(10):47–54, October 1994. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Sweet:1994:NFA

[Swe94b] W. Sweet. Nuclear fusion advances. IEEE Spectrum, 31(2): 31–36, February 1994. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Swetnam:1994:CEI

[Swe94c] G. F. Swetnam. A Capitol experience [IEEE Congressional Fellow, experience]. IEEE Spectrum, 31(7):40–43, July 1994. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Sweet:1995:GCF

[Swe95] W. Sweet. The glass cockpit [flight deck automation]. IEEE Spectrum, 32(9):30– 38, September 1995. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Sweet:1996:BNT

[Swe96a] W. Sweet. Better networks for test ban monitoring. IEEE Spectrum, 33(2):24–33, February 1996. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Sweet:1996:CSA

[Swe96b] W. Sweet. Chernobyl’s stressful after-effects. IEEE Spectrum, 33(11):26–34, November 1996. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Sweet:1996:PE

[Swe96c] W. Sweet. Power and energy. IEEE Spectrum, 33(1): 70–75, January 1996. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Sweet:1996:RM

[Swe96d] W. Sweet. Robert N. Metcalfe.

IEEE Spectrum, 33(6):48–49, June 1996. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Sweet:1996:SNC

[Swe96e] W. Sweet. A science for the next century [books]. IEEE Spectrum, 33(7):12–15, July 1996. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Sweet:1997:ARD

[Swe97a] W. Sweet. Advanced reactor development rebounding. IEEE Spectrum, 34(11): 41–48, November 1997. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Sweet:1997:CBM

[Swe97b]  W. Sweet.  Chiariglione

and the birth of MPEG. IEEE Spectrum, 34(9):70– 77, September 1997. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Sweet:1997:ID

[Swe97c]  W. Sweet.  Interim dan-

gers. IEEE Spectrum, 34(11): 16–18, November 1997. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Sweet:1997:NR

[Swe97d] W. Sweet. A nuclear reconnaissance. IEEE Spectrum, 34(11):23–28, November 1997. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Sweet:1997:PE

[Swe97e] W. Sweet. Power and energy. IEEE Spectrum, 34(1): 38–42, January 1997. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Sweet:1997:TYA

[Swe97f] W. Sweet. Thirty years ago …. IEEE Spectrum, 34(1): 20T12, January 1997. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Sweet:1998:EEE

[Swe98a] W. Sweet. Educating ethical engineering [roundtable]. IEEE Spectrum, 35(6):51–61, June 1998. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Sweet:1998:FF

[Swe98b] W. Sweet. Faults failures.

IEEE Spectrum, 35(7):87–88, July 1998. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Sweet:1998:MFG

[Swe98c] W. Sweet. A many-faceted gem (book review) [The Scientific Revolution. Shapin, Steven, University of Chicago Press, Chicago, 1996 and 1998, 218 pp., $19.95(hardcover), $12 (paper)]]. IEEE Spectrum, 35(3):12–16, March 1998. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Sweet:1998:PSS

[Swe98d] W. Sweet. A peep at some social pitfalls [books]. IEEE Spectrum, 35(5):10–12, May 1998. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Sweet:1998:PET

[Swe98e] W. Sweet. Power and energy [technology 1998 analysis and forecast]. IEEE Spectrum, 35 (1):67–72, January 1998. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Sweet:1998:TSC

[Swe98f] W. Sweet. Two sides at the (cyber) force. IEEE Spectrum, 35(11):12–14, November 1998. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Sweet:1998:UCM

[Swe98g] W. Sweet. The unnumbered charms of math [books]. IEEE Spectrum, 35(5):10, May 1998. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Sweet:1999:BLP

[Swe99a] W. Sweet. Battling for local phone customers. IEEE Spectrum, 36(5):32–37, May 1999. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Sweet:1999:FF

[Swe99b] W. Sweet. Faults failures. IEEE Spectrum, 36(12):115– 117, December 1999. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Sweet:1999:BB

[Swe99c] W. Sweet. Is bigger better? IEEE Spectrum, 36(5):38–44, May 1999. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Sweet:1999:TTM

[Swe99d] W. Sweet. Of two (or three)

minds on computationalism [books]. IEEE Spectrum, 36 (3):9–12, March 1999. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Sweet:1999:POS

[Swe99e] W. Sweet. Philosopher of the old school [books]. IEEE Spectrum, 36(3):12, March 1999. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Sweet:1999:PET

[Swe99f] W. Sweet. Power and energy [1999 technology analysis and forecast]. IEEE Spectrum, 36 (1):62–67, January 1999. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Symons:1998:TSV

[Sym98] R. S. Symons. Tubes: still vital after all these years. IEEE Spectrum, 35(4):52–63, April 1998. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Santori:1996:FBP

[SZ96] M. Santori and K. Zech. Fieldbus brings protocol to process control. IEEE Spectrum, 33 (3):60–64, March 1996. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Tabors:1996:LUN

[Tab96] R. D. Tabors. Lessons from the UK and Norway. IEEE Spectrum, 33(8):45–49, August 1996. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Tadmor:1998:PHE

[Tad98] Z. Tadmor. Proactive higher education for high technology. IEEE Spectrum, 35(5):39–42, May 1998. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Tandon:1998:OET

[Tan98] V. Tandon. Opening Europe’s telephone business to competition. IEEE Spectrum, 35(10): 43–48, October 1998. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Taur:1999:IST

[Tau99]  Y. Taur.  The incredible

shrinking transistor. IEEE Spectrum, 36(7):25–29, July 1999. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Taylor:1992:SC

[Tay92] R. M. Taylor. Space communications. IEEE Spectrum, 29 (2):30–33, February 1992. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Taylor:1999:IGB

[Tay99] C. W. Taylor. Improving grid behaviour. IEEE Spectrum, 36(6):40–45, June 1999. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

TL:1997:NLI

[TC97] Tsu-Chang Lee and J. Cong.

The new line in IC design. IEEE Spectrum, 34 (3):52–58, March 1997. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

terMaat:1997:EC

[ter97] M. ter Maat. The economics of e-cash. IEEE Spectrum, 34 (2):68–73, February 1997. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Thompson:1996:TTC

[Tho96] M. Thompson. The thick an thin of car cabling. IEEE Spectrum, 33(2):42–45, February 1996. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Thompson:1997:CSP

[Tho97] P. Thompson. Chip-scale packaging. IEEE Spectrum, 34 (8):36–43, August 1997. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Tiwary:1994:BHM

[Tiw94] G. Tiwary. Below the halfmicron mark. IEEE Spectrum, 31(11):84–87, November 1994. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Teegarden:1998:HMS

[TLN98] D. Teegarden, G. Lorenz, and R. Neuf. How to model and simulate microgyroscope systems. IEEE Spectrum, 35 (7):66–75, July 1998. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Toepfer:1998:CEP

[Toe98] C. B. Toepfer. Charge! EVs power up for the long haul. IEEE Spectrum, 35(11):41– 47, November 1998. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Topper:1991:ASD

[Top91] Andrew Topper. Automating software development. IEEE Spectrum, 28(11):56, November 1991. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Trovato:1995:SAN

[TPB95] S. A. Trovato, J. O. Parry, and J. M. Burger. Slowing the aging of nuclear power plants. IEEE Spectrum, 32 (3):32–36, March 1995. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Tracht:1993:ALD

[Tra93] A. E. Tracht. Adapting laptops to data acquisition. IEEE Spectrum, 30(10):45–47, October 1993. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Trlica:1996:SA

[Trl96] C. Trlica. Software applications. IEEE Spectrum, 33(1): 56–59, January 1996. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Trlica:1997:STT

[Trl97] C. Trlica. Software tools [technology analysis and forecast]. IEEE Spectrum, 34(1): 60–64, January 1997. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Trope:1999:HFY

[Tro99] R. L. Trope. How NOT to forfeit Y2K Act protections [law]. IEEE Spectrum, 36(10): 65–69, October 1999. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Trumble:1998:GLL

[Tru98] B. Trumble. Get the lead out! [lead free solder]. IEEE Spectrum, 35(5):55–60, May 1998. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Tremblay:1995:PPD

[TT95] M. Tremblay and P. Tirumalai. Partners in platform design. IEEE Spectrum, 32 (4):20–26, April 1995. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Turino:1991:CEM

[Tur91] J. Turino. Concurrent engineering: making it work calls for input from everyone. IEEE Spectrum, 28(7):30–32, July 1991. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Unger:1996:DRT

[Ung96] S. H. Unger. Doing the right thing. IEEE Spectrum, 33(12): 25–32, December 1996. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Ungson:1999:EKB

[UT99] G. R. Ungson and J. D. Trudel. The emerging knowledge-based economy. IEEE Spectrum, 36(5):60–65, May 1999. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

VanWyk:1996:VMM

[Van96] J. D. VanWyk. Viewpoint: The muscle of multikilowatts. IEEE Spectrum, 33(1):98– 99, January 1996. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Vendelin:1990:ENM

[Ven90] G. D. Vendelin. Evaluating nonlinear models for microwave GaAs FETs. IEEE Spectrum, 27(9):48–50, September 1990. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Vaglica:1990:HSM

[VG90] J. J. Vaglica and P. S.

Gilmour. How to select a microcontroller. IEEE Spectrum, 27(11):106–109, November 1990. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Vieira:1996:CAT

[VG96] P. Vieira and F. Gomide.

Computer-aided train dispatch. IEEE Spectrum, 33 (7):51–53, July 1996. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Voelcker:1990:APW

[Voe90a] J. Voelcker. Applications ’90:

PCs and workstations. IEEE Spectrum, 27(2):24–27, February 1990. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Voelcker:1990:AP

[Voe90b] J. Voelcker. Applications ’90:

peripherals. IEEE Spectrum, 27(2):28–30, February 1990.

CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

vonAulock:1994:SCM

[von94] W. H. von Aulock. Smart cars may not be smart. IEEE Spectrum, 31(3):16, March 1994. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Vallett:1997:FFD

[VS97] D. P. Vallett and J. M. Soden. Finding fault with deepsubmicron ICs. IEEE Spectrum, 34(10):39–50, October 1997. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Waff:1993:RDS

[Waf93] C. B. Waff. The road to the deep space network. IEEE Spectrum, 30(4):50–57, April 1993. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Wallach:1996:VWE

[Wal96] S. J. Wallach. Viewpoint: Windows everywhere? IEEE Spectrum, 33(1):44–45, January 1996. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Wallman:1997:TTS

[Wal97] S. M. H. Wallman. Technology takes to securities trading. IEEE Spectrum, 34(2): 60–65, February 1997. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Warner:1996:EMF

[War96] R. M. Warner. Earl Masterson: a fresh slant on videorecording. IEEE Spectrum, 33 (2):51–57, February 1996. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Watson:1991:TME

[Wat91a]  G. Watson.  Technology

1991: the main event 64 Mb RAM. IEEE Spectrum, 28 (1):30, January 1991. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Watson:1991:TSS

[Wat91b] G. F. Watson. Technology 1991: solid state. IEEE Spectrum, 28(1):52–55, January 1991. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Watson:1992:MRN

[Wat92a] G. F. Watson. MIL reliability: a new approach. IEEE Spectrum, 29(8):46–49, August 1992. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Watson:1992:RC

[Wat92b] G. F. Watson. Refreshing curricula. IEEE Spectrum, 29 (3):31–35, March 1992. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Watson:1992:SIP

[Wat92c] G. F. Watson. Supercomputers — interconnections and packaging. IEEE Spectrum, 29(9):69–71, September 1992. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Watson:1992:TSS

[Wat92d] G. F. Watson. Technology 1992 — solid state. IEEE Spectrum, 29(1):42–44, January 1992. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Waters:1997:RSV

[WB97] R. C. Waters and J. W. Barrus. The rise of shared virtual environments. IEEE Spectrum, 34(3):20–25, March 1997. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Wheeler:1991:CES

[WBR91] R. Wheeler, R. W. Burnett, and A. Rosenblatt. Concurrent engineering: success stories in instrumentation, communications. IEEE Spectrum, 28(7):32–37, July 1991. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Werner:1993:FPF

[Wer93]  K. I. Werner.  The flat

panel’s future. IEEE Spectrum, 30(11):18–22, November 1993. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Werner:1994:HVL

[Wer94] K. Werner. Higher visibility for LEDs. IEEE Spectrum, 31(7):30–34, July 1994. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Werner:1995:UDI

[Wer95] K. Werner. US display industry on the edge. IEEE Spectrum, 32(5):62–69, May 1995. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Werner:1997:FFD

[Wer97] K. Werner. The flowering of flat displays. IEEE Spectrum, 34(5):40–49, May 1997. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Wang:1999:APP

[WH99] Yanjia Wang and Kebin He.

The air pollution picture in China. IEEE Spectrum, 36 (12):55–58, December 1999. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

White:1992:CUC

[Whi92] R. M. White. A catalyst for US competitiveness. IEEE Spectrum, 29(3):49–50, March 1992. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

White:1993:CM

[Whi93a] R. M. White. Competitive measures. IEEE Spectrum, 30 (4):29–33, April 1993. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Whitney:1993:WPT

[Whi93b] D. Whitney. When people are too large and dirty (manufacturing computer control). IEEE Spectrum, 30(9): 39–42, September 1993. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Wickelgren:1996:SSO

[Wic96a] I. Wickelgren. The strange senses of other species. IEEE Spectrum, 33(3):32–37, March 1996. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Wickelgren:1996:LAN

[Wic96b]  I. J. Wickelgren.  Local-

area networks go wireless. IEEE Spectrum, 33(9):34– 40, September 1996. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Wickelgren:1997:FAF

[Wic97] I. J. Wickelgren. The facts about FireWire [serial communication bus]. IEEE Spectrum, 34(4):19–25, April 1997.

CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Williams:1993:MVV

[Wil93] J. Williams. Mixing 3-V and 5-V ICs. IEEE Spectrum, 30 (3):40–42, March 1993. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Willner:1997:MOB

[Wil97] A. E. Willner. Mining the optical bandwidth for a terabit per second. IEEE Spectrum, 34(4):32–41, April 1997. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Winters:1996:FMA

[Win96] J. M. Winters. Flatland movement analyzed interactively. IEEE Spectrum, 33 (4):17–18, April 1996. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Wise:1996:FAE

[Wis96] G. Wise. First among equals [book review]. IEEE Spectrum, 33(5):10–11, May 1996. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Wise:1999:ELI

[Wis99] G. Wise. Edison: A Life of Invention, Israel, Paul, John Wiley, New York, 1998, 552 pp., $30 [book review]. IEEE Spectrum, 36(1):10–12, January 1999. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Werblin:1996:CE

[WJT96] F. Werblin, A. Jacobs, and J. Teeters. The computational eye. IEEE Spectrum, 33(5):30–37, May 1996. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Whittaker:1990:JRA

[WK90] W. L. Whittaker and T. Kanade. Japan robotics aim for unmanned space exploration. IEEE Spectrum, 27(12):64– 67, December 1990. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Werb:1998:DPS

[WL98] J. Werb and C. Lanzl. Designing a positioning system for finding things and people indoors. IEEE Spectrum, 35(9): 71–78, September 1998. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Wolff:1990:VGG

[Wol90] M. F. Wolff. Van de Graaff’s generator. IEEE Spectrum, 27(7):46, July 1990. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Wolff:1992:LNS

[Wol92] H. Wolff. Learning new skills.

IEEE Spectrum, 29(12):58– 61, December 1992. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Wolff:1993:HEV

[Wol93] H. Wolff. How engineers view themselves. IEEE Spectrum, 30(4):24–28, April 1993. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Wolk:1999:FCH

[Wol99] R. H. Wolk. Fuel cells for homes and hospitals. IEEE Spectrum, 36(5):45–52, May 1999. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Wood:1990:MM

[Woo90] R. Wood. Magnetic megabits.

IEEE Spectrum, 27(5):32–33, May 1990. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Wouk:1995:HTN

[Wou95] V. Wouk. Hybrids: then and now. IEEE Spectrum, 32 (7):16–21, July 1995. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Wyatt:1996:OIB

[WR96] J. Wyatt and J. Rizzo. Ocular implants for the blind.

IEEE Spectrum, 33(5):47–53, May 1996. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Weiss:1993:PDD

[WS93] J. Weiss and D. Schremp. Putting data on a diet. IEEE Spectrum, 30(8):36–39, August 1993. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Yale:1997:PRD

[Yal97] K. Yale. Preparing the right data diet for training neural networks. IEEE Spectrum, 34 (3):64–66, March 1997. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Yeager:1996:VCV

[Yea96] K. E. Yeager. Viewpoint: Count the value of R&D, not just its cost. IEEE Spectrum, 33(1):74–75, January 1996. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Youst:1990:TLG

[YL90] D. B. Youst and L. Lipsett. The technical ladder gets harder to climb (engineering career advancement). IEEE Spectrum, 27(9):46– 47, September 1990. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Youst:1994:SFT

[YL94] D. B. Youst and L. Lipsett.

Survival of the fittest [technical skills development]. IEEE Spectrum, 31(10):67–69, October 1994. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Yoffa:1996:VMT

[Yof96]  E. J. Yoffa.  Viewpoint:

Methodology, not tools, is the key to winning the design race. IEEE Spectrum, 33 (1):58, January 1996. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Younse:1993:MC

[You93] J. M. Younse. Mirrors on a chip. IEEE Spectrum, 30(11): 27–31, November 1993. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Yuan:1991:PLS

[YS91] J. Yuan and C. Svensson.

Pushing the limits of standard CMOS. IEEE Spectrum, 28 (2):52–53, February 1991. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Zachary:1995:VBB

[Zac95] G. P. Zachary. Vannevar Bush [biography]. IEEE Spectrum, 32(7):65–69, July 1995. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Zakheim:1999:ORN

[Zak99] D. S. Zakheim. Old rivalries, new arsenals: should the US worry? IEEE Spectrum, 36 (3):29–31, March 1999. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Zinser:1994:MPP

[ZF94] K. Zinser and F. Frischenschlager. Multimedia’s push into power. IEEE Spectrum, 31(7):44–48, July 1994. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Zimmerman:1997:EEW

[Zim97]  D. Zimmerman.  Exploit-

ing enemy weaknesses [books]. IEEE Spectrum, 34(8):10, August 1997. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Zorpette:1990:DSC

[Zor90a] Glenn Zorpette. Defense: suppliers are challenged to downsize gracefully as they seek to diversify. IEEE Spectrum, 27(11):33–36, November 1990. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Zorpette:1990:ECI

[Zor90b] Glenn Zorpette. Electronics consortia to impact products for generations. IEEE Spectrum, 27(10):50–52, October 1990. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Zorpette:1990:FHE

[Zor90c] Glenn Zorpette. Fish and hydroelectricity: engineering a better coexistence. IEEE Spectrum, 27(12):68– 70, December 1990. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Zorpette:1990:MM

[Zor90d] Glenn Zorpette. Multinationals on the move. IEEE Spectrum, 27(6):50–52, June 1990. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Zorpette:1990:NPP

[Zor90e] Glenn Zorpette. New pay phones hit the street. IEEE Spectrum, 27(5):28–31, May 1990. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Zorpette:1990:PE

[Zor90f] Glenn Zorpette. Power and energy. IEEE Spectrum, 27 (6):28–30, June 1990. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Zorpette:1990:TMM

[Zor90g] Glenn Zorpette. Technology ’90: minis and mainframes. IEEE Spectrum, 27 (1):30–34, January 1990. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Zorpette:1990:TPE

[Zor90h] Glenn Zorpette. Technology ’90: power and energy. IEEE Spectrum, 27(1):44–46, January 1990. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Zorpette:1991:EML

[Zor91a] Glenn Zorpette. Energy management — loosening the bonds of oil. IEEE Spectrum, 28(3):34–38, March 1991. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Zorpette:1991:EMD

[Zor91b] Glenn Zorpette. Energy management — the DOE: agency under fire. IEEE Spectrum, 28 (3):32–33, March 1991. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Zorpette:1991:GLE

[Zor91c] Glenn Zorpette. Gulf legacy: emulating the battlefield. IEEE Spectrum, 28(9):36– 39, September 1991. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Zorpette:1991:GLF

[Zor91d] Glenn Zorpette. Gulf legacy: from factory floor to desert war. IEEE Spectrum, 28(9):

40–42, September 1991. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Zorpette:1991:TMM

[Zor91e] Glenn Zorpette. Technology 1991: minis and mainframes. IEEE Spectrum, 28 (1):40–43, January 1991. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Zorpette:1991:TPE

[Zor91f] Glenn Zorpette. Technology 1991: power and energy. IEEE Spectrum, 28(1):61–64, January 1991. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Zorpette:1991:UGS

[Zor91g] Glenn Zorpette. Utilities get serious about efficiency. IEEE Spectrum, 28(5):42–43, May 1991. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Zorpette:1992:EL

[Zor92a] Glenn Zorpette. Elizabeth Laverick. IEEE Spectrum, 29 (5):40–42, May 1992. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Zorpette:1992:HGP

[Zor92b] Glenn Zorpette. Hawaii’s geothermal program. IEEE Spectrum, 29(2):49–50, February 1992. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Zorpette:1992:RXR

[Zor92c] Glenn Zorpette. Rethinking X-ray lithography. IEEE Spectrum, 29(6):33–36, June 1992. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Zorpette:1992:SNS

[Zor92d]  Glenn Zorpette.  Seeking

nuclear safeguards. I. How Iraq reverse-engineered the bomb. IEEE Spectrum, 29 (4):20–24, April 1992. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Zorpette:1992:SPP

[Zor92e] Glenn Zorpette. Supercomputers — the power of parallelism. IEEE Spectrum, 29(9): 28–33, September 1992. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Zorpette:1992:TLC

[Zor92f] Glenn Zorpette. Technology 1992 — large computers. IEEE Spectrum, 29(1): 33–35, January 1992. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Zorpette:1992:TPE

[Zor92g] Glenn Zorpette. Technology 1992 — power and energy. IEEE Spectrum, 29(1):

49–51, January 1992. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Zorpette:1993:ESC

[Zor93a] Glenn Zorpette. Environment — sensing climate change. IEEE Spectrum, 30(7):20–27, July 1993. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Zorpette:1993:TLC

[Zor93b] Glenn Zorpette. Technology 1993 — large computers. IEEE Spectrum, 30(1): 34–37, January 1993. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Zorpette:1993:TPE

[Zor93c] Glenn Zorpette. Technology 1993 — power and energy. IEEE Spectrum, 30(1): 61–64, January 1993. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Zorpette:1994:AD

[Zor94a] Glenn Zorpette. Autopilots of the deep. IEEE Spectrum, 31 (8):38–44, August 1994. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Zorpette:1994:PPG

[Zor94b] Glenn Zorpette. A peripatetic prodigy goes home. IEEE Spectrum, 31(7):56–58, July 1994. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Zorpette:1994:PE

[Zor94c] Glenn Zorpette. Power and energy. IEEE Spectrum, 31 (1):58–61, January 1994. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Zorpette:1994:SSE

[Zor94d] Glenn Zorpette. A slow start for emissions trading. IEEE Spectrum, 31(7):49–52, July 1994. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Zorpette:1994:TI

[Zor94e] Glenn Zorpette. Technology in India. IEEE Spectrum, 31 (3):24–32, March 1994. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Zorpette:1995:PE

[Zor95a] Glenn Zorpette. Power and energy. IEEE Spectrum, 32 (1):56–59, January 1995. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Zorpette:1995:RAN

[Zor95b] Glenn Zorpette. Radio astronomy: new windows on the Universe. IEEE Spectrum, 32 (2):18–19, February 1995. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Zorpette:1997:ETM

[Zor97] Glenn Zorpette. Entertainment for the technically mature [book reviews]. IEEE Spectrum, 34(10):12–14, October 1997. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Zorian:1999:TMC

[Zor99] Y. Zorian. Testing the monster chip. IEEE Spectrum, 36(7):54–60, July 1999. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Zorpette:1990:NWC

[ZS90] Glenn Zorpette and Gary Stix.

Nuclear waste: the challenge is global. IEEE Spectrum, 27 (7):18–24, July 1990. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).